### v1

# Egorov-Type Semiclassical Limits for Open Quantum Systems with a Bi-Lindblad Structure

Abstract
This paper develops a bridge between bi-Hamiltonian structures of Poisson–Lie type, contact Hamiltonian dynamics, and the Gorini–Kossakowski–Sudarshan–Lindblad (GKSL) formalism for quantum open systems. On the classical side, we consider bi-Hamiltonian systems defined by a Poisson pencil with non-trivial invariants. Using an exact symplectic realization, these invariants are lifted and projected onto a contact manifold, yielding a completely integrable contact Hamiltonian system in terms of dissipated quantities and a Jacobi-commutative algebra of observables. On the quantum side, we introduce a class of contact-compatible Lindblad generators: GKSL evolutions whose dissipative part preserves a commutative $C^{\ast}$-subalgebra generated by the quantizations of the classical dissipated quantities, and whose Hamiltonian part admits an Egorov-type semiclassical limit to the contact dynamics. This construction provides a mathematical mechanism compatible with the semiclassical limit for pure dephasing, compatible with integrability and contact dissipation. An explicit Euler-top-type Poisson–Lie pencil, inspired by deformed Euler top models, is developed as a fully worked-out example illustrating the resulting bi-Lindblad structure and its semiclassical behavior.
Keywords: open quantum systems, semiclassical limits, integrable systems, Lindblad equation, GKSL formalism, bi-Hamiltonian, contact geometry.
MSC 2020 codes: 81S22 (primary); 37J35, 37J55, 46L60, 53Z05, 81Q70, 81R12, 81S10.
1 Introduction
Bi-Hamiltonian geometry plays a central role in the theory of integrable systems. A bi-Hamiltonian system consists of a smooth manifold endowed with two compatible Poisson tensors $\Pi_{0},\Pi_{1}$, meaning that every linear combination
$\Pi_{\lambda}\mathrel{\vbox{\hbox{\scriptsize.}\hbox{\scriptsize.}}}=\Pi_{1}-\lambda\Pi_{0},\qquad\lambda\in\mathbb{R},$
is again a Poisson tensor. A vector field $X$ is called bi-Hamiltonian if there exist Hamiltonian functions $H_{0},H_{1}$ such that
$X=\Pi_{0}^{\sharp}(\mathrm{d}H_{0})=\Pi_{1}^{\sharp}(\mathrm{d}H_{1})\,.$
Compatibility ensures the existence of a recursion operator (on a symplectic leaf where one of the Poisson structures is non-degenerate) and hierarchies of commuting Hamiltonians, which in turn give rise to rich algebraic and geometric structures and, under suitable regularity assumptions, to complete integrability in the sense of Liouville; see, for example, [M.M1984, C.M.P1993, M.C.F+1997, F.M.P2000, Fernandes1994, G.M.S2003].
A natural and geometrically significant source of bi-Hamiltonian systems is provided by Poisson–Lie groups. A Poisson tensor $\pi$ on a Lie group $G$ is called multiplicative if the group multiplication $m\colon G\times G\longrightarrow G$ is a Poisson map from $(G\times G,\pi\oplus\pi)$ to $(G,\pi)$. Multiplicative Poisson structures and their infinitesimal counterparts, Lie bialgebras, were introduced by Drinfel’d in the context of quantum groups [Drinfeld1988, Drinfeld1990] and further developed in the foundational work of Semenov–Tian–Shansky on the classical $r$-matrix formalism, Poisson pencils and factorization dynamics [STS1983, STS1985]. The global geometry of Poisson–Lie groups, including dressing transformations and symplectic groupoids, was elucidated in [LuWeinstein1990].
When two multiplicative Poisson tensors $\pi_{0},\pi_{1}$ on $G$ are compatible, the resulting Poisson pencil $\pi_{\lambda}=\pi_{1}-\lambda\pi_{0},\,\lambda\in\mathbb{R}$ encodes a rich class of bi-Hamiltonian vector fields on $G$. In particular, if $H_{0},H_{1}\in\mathscr{C}^{\infty}(G)$ satisfy
$X=\pi_{0}^{\sharp}(\mathrm{d}H_{0})=\pi_{1}^{\sharp}(\mathrm{d}H_{1}),$
then $X$ is a bi-Hamiltonian vector field whose integrability properties reflect the algebraic structure of the underlying Lie bialgebra and classical $r$-matrix. This mechanism underlies many classical integrable models on Lie groups and Poisson–Lie groups [STS1983, STS1985, KS2004, KS2008, LuWeinstein1990].
On the other hand, contact geometry provides a natural framework for describing non-Hamiltonian classical dynamics in a plethora of contexts, including certain dissipative systems and thermodynamically relevant flows; see, e.g., [d.L2019a, d.L2019b, G.G2022a, G.G2024a, Grabowski2013a, Lopez-Gordon2024]. There is a one-to-one correspondence between contact manifolds and homogeneous symplectic manifolds, which has been known for a while [Arnold1978, L.M1987], and whose applications to contact Hamiltonian dynamics have been recently analyzed in detail by Grabowska and Grabowski [Grabowski2013a, G.G2022a, G.G2023, G.G2024a]. We have exploited this correspondence in order to establish a novel characterisation of completely integrable contact Hamiltonian systems as completely integrable Hamiltonian systems (in the standard Liouville sense) that satisfy certain homogeneity conditions [C.d.L+2023, Lopez-Gordon2024, Colombo2025].
A key point, which we emphasise here, is that:
  • the integrability of the underlying symplectic dynamics is still controlled by Poisson (or Poisson–Nijenhuis) structures and their compatibility, exactly as in the standard bi-Hamiltonian theory;
  • the homogeneity is an independent geometric ingredient that guarantees that the dynamics and its commuting integrals can be projected onto a contact manifold.
In particular, in [Colombo2025], it is shown, in concrete examples, that one may take a Poisson structure on the symplectization compatible with the canonical one, use the corresponding recursion operator to construct a maximal family of homogeneous functions in involution, and then “project” these functions to a $(2n+1)$-dimensional contact manifold as functions in Jacobi-involution. What turns out to be too rigid for contact integrability is not Poisson compatibility on the symplectization, but the use of compatible Jacobi structures (see [NunesdaCosta1998a, M.M.P1999, C.D.N2010, P.N2003, Petalidou2002]) directly on the original contact manifold: compatible Jacobi pairs do not yield a recursion operator with enough independent eigenvalues to produce a full set of contact integrals.
This paper aims to unify bi-Hamiltonian Poisson geometry with the homogeneous/contact correspondence and to connect the resulting classical framework with the quantum realm. In doing so, we bring together three layers of structure:
  1. 1.
    bi-Hamiltonian dynamics on Poisson–Lie groups and their symplectic (or exact symplectic) realizations;
  2. 2.
    the projection of homogeneous Hamiltonian systems to contact dynamics via Liouville-transverse hypersurfaces;
  3. 3.
    quantum open-system evolution of GKSL type.
To avoid conceptual ambiguity, we stress a fundamental distinction:
  • Systems defined on Poisson–Lie groups, and on any of their symplectic or exact symplectic realizations, are entirely classical Hamiltonian systems. Their integrability is governed by Poisson, Poisson–Nijenhuis and bi-Hamiltonian structures in the usual sense.
  • The Lindblad equation that we construct is a genuinely quantum evolution equation (of GKSL type), acting on density matrices over a suitable Hilbert space obtained by quantizing $G$ or one of its symplectic realizations. Its dissipative part has no classical analog at the level of closed Hamiltonian dynamics, although it admits a meaningful semiclassical description.
A contact Hamiltonian system is a triple $(C,\alpha,h)$ consisting of a manifold $C$ endowed with a contact form $\alpha$ and a fixed function $h\in\mathscr{C}^{\infty}(M)$. In this context, by a dissipated quantity we mean a classical observable $I\in\mathscr{C}^{\infty}(C)$ such that
$\{I,h\}_{\alpha}=0,$
where $\{\cdot,\cdot\}_{\alpha}$ is the Jacobi bracket defined by $\alpha$. The trivial symplectization of $(C,\alpha,h)$ is given by $(M=C\times(0,+\infty),\omega=\mathrm{d}(r\alpha),H=rh)$, where $r$ denotes the global coordinate of $(0,+\infty)$. There is a Lie algebra isomorphism between the Jacobi bracket $\{\cdot,\cdot\}_{\alpha}$ and the Poisson bracket $\{\cdot,\cdot\}_{\omega}$, which implies that conserved quantities for $(C,\alpha,h)$ are conserved quantities for $(M,\omega,H)$.
A symplectic form $\omega$ on a manifold $M$ is exact (i.e, there exists a one-form $\theta$ such that $\mathrm{d}\theta=-\omega$) if and only if there exists a Liouville vector field $\Delta$ on $M$ such that $\mathcal{L}_{\Delta}\omega=\omega$. Indeed, one can define $\theta=-\iota_{\Delta}\omega$. A tensor field $A$ on $M$ is called homogeneous of degree $w\in\mathbb{R}$ if $\mathcal{L}_{\Delta}A=w\cdot A$. On each hypersurface $C$ of $M$ which is transverse to $\Delta$, the restriction $\alpha={\left.\kern-1.2pt\theta\right|_{C}}$ is a contact form, and homogeneous functions of degree $1$ on $M$ project onto functions on $C$. Conversely, given a function $g$ on $C$, one can extend it, via the flow $\phi_{t}$ of $\Delta$, to a homogeneous function $\hat{g}$ on $M$ given by $\hat{g}(\phi_{t}(x))=e^{t}g(x)$ for each $x\in C$.
By quantum dissipation we refer to the irreversible, non-unitary component of the reduced dynamics of an open quantum system, described at the Markovian level by a GKSL (Lindblad) generator. This dissipative term arises from tracing out environmental degrees of freedom and accounts for phenomena such as decoherence, relaxation and thermalisation [BreuerPetruccione]. While such dissipation has no direct analog in closed Hamiltonian mechanics, it admits effective classical limits. In the present work we exhibit a class of GKSL generators whose semiclassical limit is naturally described by contact Hamiltonian dynamics, with dissipated quantities playing the role of robust invariants of the resulting reduced evolution.
Our main goal is to show that the classical dissipated quantities arising (via realization and projection) from a bi-Hamiltonian Poisson–Lie system admit quantum counterparts that are invariant under a suitably constructed GKSL generator, thereby establishing a precise bridge between Poisson–Lie integrability, homogeneous/contact Hamiltonian geometry, and physically admissible quantum dissipation.
Before introducing the specific quantization scheme, it is useful to explain why such a quantum layer is natural in the present setting. A bi-Hamiltonian Poisson–Lie system produces, through an exact and homogeneous symplectic realization, a completely integrable contact system whose dynamics is intrinsically non-Hamiltonian, as contact vector fields does not preserve either the contact form or the associated “Hamiltonian” function. From the viewpoint of physics, contact systems appear precisely as effective or reduced descriptions of systems interacting with an environment, where certain degrees of freedom are not explicitly modelled. Quantum open systems [BreuerPetruccione, Lindblad1976], governed by Lindblad master (GKSL) equations, arise from a conceptually analogous mechanism: a unitary evolution on a larger Hilbert space is reduced to a non-unitary dynamics on the subsystem of interest.
This parallel suggests that the contact dynamics obtained from the homogeneous projection of a Poisson–Lie system should admit a natural quantum analog. However, making this correspondence precise requires a quantization of the classical observables on the symplectic realization $(M,\omega)$ (or directly on the Poisson–Lie group $G$), together with a quantum evolution whose semiclassical limit reproduces the contact flow. To interpret the classical dissipated quantities $I_{0},\dots,I_{n}$ as quantum constants of motion, and to construct a GKSL generator whose dissipative part reflects the geometric content of the contact structure, one must assign to each classical observable $f\in\mathscr{C}^{\infty}(M)$ a quantum operator acting on a Hilbert space.
In more physical terms, by “quantizing” the above classical data we mean choosing a Hilbert space $\mathcal{H}$ and a family of maps
$Q_{\hbar}\colon\mathscr{C}^{\infty}(M)\;\longrightarrow\;\mathcal{B}(\mathcal{H}),\qquad f\longmapsto\widehat{f}=Q_{\hbar}(f),$
defined for sufficiently small $\hbar>0$, such that $\widehat{f}$ is interpreted as the quantum observable corresponding to the classical observable $f$. Here $\mathcal{H}$ denotes a (finite-dimensional) Hilbert space, and $\mathcal{B}(\mathcal{H})$ the algebra of linear operators on $\mathcal{H}$. Since $\mathcal{H}$ is finite-dimensional, every linear operator on $\mathcal{H}$ is automatically bounded, so that $\mathcal{B}(\mathcal{H})=\mathrm{End}(\mathcal{H})$ and no domain or topological issues arise. In particular, all operator topologies coincide and the algebraic formulation of the quantum dynamics is completely sufficient for our purposes.
The assumption of finite-dimensionality is made for conceptual clarity and to avoid technicalities related to unbounded operators; it does not impose any restriction on the classical phase space $M$ itself. Rather, it should be understood as a semiclassical or effective quantization of the Poisson algebra of observables, capturing the contact dynamics of interest in a finite quantum state space. All constructions below extend, under standard hypotheses, to separable infinite-dimensional Hilbert spaces.
Depending on the context, $Q_{\hbar}$ may be realized by geometric quantization of a symplectic realization of $G$, by (unitary) representation theory of $G$, or by deformation quantization of the Poisson manifold $(G,\pi_{0})$ or $(M,\omega)$. Quantum states are positive, trace-one operators $\rho$ on $\mathcal{H}$, and their time evolution is described by a GKSL generator of the form
$\mathcal{L}_{\hbar}(\rho)=-\frac{i}{\hbar}\big[\,\widehat{H}_{\hbar},\rho\,\big]+\sum_{k=1}^{N(\hbar)}\Big(L_{k,\hbar}\,\rho\,L_{k,\hbar}^{\dagger}-\tfrac{1}{2}\,\big\{\,L_{k,\hbar}^{\dagger}L_{k,\hbar},\,\rho\,\big\}\Big),$ (1.1)
where $\widehat{H}_{\hbar}$ is the quantization of a classical Hamiltonian, $\cdot$$\cdot$] is the operator commutator, and $\{\cdot,\cdot\}$ denotes the operator anti-commutator, {A,B} . . $\{A,B\}\mathrel{\vbox{\hbox{\scriptsize.}\hbox{\scriptsize.}}}=AB$+BA. The family $\{L_{k,\hbar}\}_{k=1}^{N(\hbar)}$ encodes the coupling to the environment; here $N(\hbar)\in\mathbb{N}$ is finite for each fixed $\hbar$ and may depend on $\hbar$ in the chosen semiclassical regime. In the constructions considered in this paper, $\widehat{H}_{\hbar}$ is chosen as the quantization of the homogeneous Hamiltonian $H$ on the exact symplectic realization $(M,\omega=-\mathrm{d}\theta)$ of a Poisson–Lie group, whose restriction to a Liouville-transverse hypersurface $C\subset M$ yields the contact Hamiltonian $h={\left.\kern-1.2ptH\right|_{C}}$.
The semiclassical limit $\hbar\to 0$ is understood either in the sense of Wigner transforms or within deformation quantization. In this limit, commutators of quantized observables reproduce the underlying classical brackets: for observables $a,b$ on the symplectic realization $(M,\omega)$ one has
$\frac{1}{i\hbar}\big[\,Q_{\hbar}(a),Q_{\hbar}(b)\,\big]\;\longrightarrow\;\{a,b\}_{\omega}\qquad\text{as }\hbar\to 0,$
and, after restriction to a Liouville-transverse hypersurface $C\subset M$, this induces the corresponding Jacobi bracket $\{\cdot,\cdot\}_{\alpha}$ on the contact manifold $(C,\alpha)$.
Accordingly, the GKSL evolution (1.1) reduces in the semiclassical limit to the classical Hamiltonian dynamics on $(M,\omega)$ or, after projection, to the contact Hamiltonian dynamics on $(C,\alpha,h)$. In particular, the quantum observables associated with the classical dissipated quantities $I_{0},\dots,I_{n}$ are required to become constants of motion in the Heisenberg picture, and their semiclassical limit recovers the dissipated quantities of the underlying contact integrable system.
At the technical level, the semiclassical passage from the Heisenberg-picture GKSL evolution to the underlying classical contact dynamics will be formulated through an Egorov-type requirement: quantized observables evolved by $\mathcal{L}_{\hbar}^{\ast}$ must be approximated, to leading order in $\hbar$, by classical transport along the corresponding (contact) flow, where $\mathcal{L}_{\hbar}^{\ast}$ denotes the adjoint GKSL generator in the Heisenberg picture, defined with respect to the trace duality. This viewpoint naturally leads to invariant commutative $C^{\ast}$-subalgebras generated by quantizations of the classical dissipated quantities, which will play a key role in our analysis of dephasing-type mechanisms.
The classical part of our construction proceeds as follows. Starting from a bi-Hamiltonian Poisson–Lie system, we consider an exact symplectic realization where the Hamiltonian is homogeneous of degree one with respect to the Liouville vector field. On this exact symplectic manifold we exploit bi-Hamiltonian techniques to build a family of commuting integrals. Homogeneity then allows us to project the dynamics and its integrals to a contact manifold, obtaining a completely integrable contact system in the sense of [C.d.L+2023, Lopez-Gordon2024].
The main question we address is whether one can construct an open quantum dynamics naturally associated with this classical framework, whose semiclassical limit reproduces the contact Hamiltonian flow.
Problem 1.
Let $G$ be a Poisson–Lie group endowed with two compatible multiplicative Poisson bivectors $\pi_{0},\pi_{1}$, and let $H_{0},H_{1}\in\mathscr{C}^{\infty}(G)$ be Hamiltonian functions such that the corresponding Hamiltonian vector field
$X=\pi_{0}^{\sharp}(\mathrm{d}H_{0})=\pi_{1}^{\sharp}(\mathrm{d}H_{1})$
is bi-Hamiltonian on $(G,\pi_{0},\pi_{1})$.
Suppose that $(G,\pi_{0})$ admits an exact symplectic realization $(M,\theta)$ and that $X$ lifts to a homogeneous Hamiltonian system $(M,\theta,H)$, with $H$ of degree $1$ with respect to the Liouville vector field. Let $C\subset M$ be a hypersurface transverse to the Liouville flow, and set
$\alpha\mathrel{\vbox{\hbox{\scriptsize.}\hbox{\scriptsize.}}}=\theta|_{C},\qquad h\mathrel{\vbox{\hbox{\scriptsize.}\hbox{\scriptsize.}}}=H|_{C},$
so that $(C,\alpha,h)$ is a $(2n+1)$-dimensional completely integrable contact Hamiltonian system, with classical dissipated quantities
$I_{0}\mathrel{\vbox{\hbox{\scriptsize.}\hbox{\scriptsize.}}}=h,\quad I_{1},\dots,I_{n}\in\mathscr{C}^{\infty}(C)$
coming from the bi-Hamiltonian hierarchy on $G$.
Can one construct a Lindblad generator—that is, a physically admissible GKSL generator $\mathcal{L}$ of quantum open-system dynamics on a suitable quantum state space (for instance, on density matrices over a Hilbert space $\mathcal{H}$ obtained by quantizing $G$ or $M$) such that:
  1. 1.
    its semiclassical limit (in the sense of $\hbar\to 0$ or deformation quantization) reproduces the contact Hamiltonian flow of $(C,\alpha,h)$;
  2. 2.
    its conserved observables $\widehat{I}_{0},\dots,\widehat{I}_{n}$ are quantizations of the classical dissipated quantities $I_{0},\dots,I_{n}$?
Our answer is affirmative. We introduce the notion of a contact-compatible Lindblad generator: a GKSL dynamics whose Hamiltonian and dissipative parts are organized so that
  • the Heisenberg-picture evolution admits a well-defined semiclassical limit on the contact phase space $(C,\alpha)$, reproducing the contact flow of $h$;
  • the quantum constants of motion of the Lindblad evolution correspond, in the semiclassical limit, to the classical dissipated quantities of the contact integrable system.
We also introduce a notion of bi-Lindblad structure, inspired by bi-Hamiltonian geometry: a pair of Lindblad generators whose convex combinations remain Lindbladian and whose common invariant observables encode the classical integrals of a bi-Hamiltonian system on a Poisson–Lie group. This provides a geometric bridge between Poisson–Lie geometry, contact integrable dynamics and physically admissible quantum dissipation.
An Euler-top-type Poisson–Lie system, inspired by the deformed Euler top introduced in [ballesteros2017poisson, G.I.M+2023], is developed here as a fully explicit example illustrating the classical–quantum correspondence between bi-Hamiltonian contact dynamics and Lindblad open-system evolution.
Structure of the paper. Section 2 reviews bi-Hamiltonian systems, Poisson pencils and homogeneous Hamiltonian dynamics, with particular emphasis on how Poisson compatibility and homogeneity give rise to Jacobi-commutative families of integrals. Section 3 develops the passage from Poisson–Lie bi-Hamiltonian data to completely integrable contact systems via homogeneous symplectization and Liouville reduction, introducing the notion of dissipated quantities and their associated Jacobi algebras. Section 4 introduces contact-compatible and bi-Lindblad generators in the GKSL formalism, establishes an Egorov-type semiclassical criterion, and analyzes the emergence of invariant commutative $C^{\ast}$-algebras and pure dephasing mechanisms. Section 5 illustrates the construction on an explicit low-dimensional Poisson–Lie pencil of Euler-top type, inspired by deformed Euler-top models, showing how integrability, dissipation and decoherence are combined in a bi-Lindblad framework. Finally, we conclude with a discussion of the physical interpretation and connections with integrable open quantum systems.
Notation and conventions. All manifolds and maps are $\mathscr{C}^{\infty}$-smooth. Given a vector field $X$ and a $p$-form $\alpha$ on a manifold $M$, the exterior derivative of $\alpha$ is denoted by $\mathrm{d}\alpha$, the contraction of $\alpha$ with $X$ is denoted by $\iota_{X}\alpha$. The Lie derivative of a tensor field $A$ on $M$ with respect to $X$ is denoted by $\mathcal{L}_{X}A$. The positive real numbers will be denoted by $\mathbb{R}_{>0}$. Given a vector bundle $E\to M$, the space of sections of $E$ is denoted by $\Gamma(E)$. The Jacobi (resp. Poisson) bracket defined by a contact form $\alpha$ (resp. symplectic form $\omega$) is denoted by $\{\cdot,\cdot\}_{\alpha}$ (resp. $\{\cdot,\cdot\}_{\omega}$). If $M\simeq\mathbb{R}^{2d}$, the standard class of semiclassical symbols of order zero on $M$ will be denoted by $S^{0}(M)$, and $\mathcal{S}(\mathbb{R}^{d})$ will denote the Schwartz space of rapidly decreasing smooth functions on $\mathbb{R}^{d}$. For simplicity’s sake, we will assume Hilbert spaces to be finite-dimensional. Given a Hilbert space $\mathcal{H}$, we will denote the $C^{\ast}$-algebra of (bounded) linear operators on $\mathcal{H}$ by $\mathcal{B}(\mathcal{H})$. The operator anticommutator is {A,B} . . $\{A,B\}\mathrel{\vbox{\hbox{\scriptsize.}\hbox{\scriptsize.}}}=AB$+BA for each $A,B\in\mathcal{B}(\mathcal{H})$.
2 Bi-Hamiltonian geometry and contact integrability
This section reviews classical bi-Hamiltonian geometry and homogeneous Hamiltonian systems on exact symplectic manifolds, and recalls the notion of contact integrability introduced in [C.d.L+2023, Lopez-Gordon2024].
In this section we work entirely in the classical realm: Poisson, symplectic and contact structures live on smooth manifolds, and no quantization procedure is involved. The contact systems and homogeneous Hamiltonian systems constructed here will later serve as the classical (or semiclassical) phase-space models for the GKSL generators introduced in Section 4.
On the one hand, bi-Hamiltonian geometry is entirely formulated in the Poisson (or symplectic) category: compatibility of Poisson structures leads to recursion operators and hierarchies of commuting Hamiltonians [M.M1984, M.C.F+1997, F.M.P2000, Fernandes1994, C.M.P1993, G.M.S2003]. On the other hand, homogeneity is a property with respect to the Liouville vector field on an exact symplectic manifold. In our context these two ingredients are logically independent: bi-Hamiltonian (or Poisson–Nijenhuis) structures will be used to generate integrals in involution in the usual symplectic sense, while homogeneity will be the geometric mechanism that allows us to project these integrals to a $(2n+1)$-dimensional contact manifold via symplectization.
2.1 Bi-Hamiltonian structures
Let $M$ be a smooth manifold endowed with two Poisson tensors $\Pi_{0},\Pi_{1}\in\Gamma(\wedge^{2}{\mathsf{T}}M)$.
Definition 2.1.
The pair $(\Pi_{0},\Pi_{1})$ is compatible if every linear combination
$\Pi_{\lambda}\mathrel{\vbox{\hbox{\scriptsize.}\hbox{\scriptsize.}}}=\Pi_{1}-\lambda\Pi_{0},\qquad\lambda\in\mathbb{R},$
is a Poisson tensor.
The family $\Pi_{\lambda}$ of Poisson tensors is called a Poisson pencil, and it defines a family of Poisson brackets
$\{f,g\}_{\lambda}\mathrel{\vbox{\hbox{\scriptsize.}\hbox{\scriptsize.}}}=\Pi_{\lambda}(\mathrm{d}f,\mathrm{d}g)\,,\quad\forall\,f,g\in\mathscr{C}^{\infty}(M).$
Definition 2.2.
A vector field $X\in\mathfrak{X}(M)$ is bi-Hamiltonian with respect to $(\Pi_{0},\Pi_{1})$ if there exist smooth functions $H_{0},H_{1}\in\mathscr{C}^{\infty}(M)$ such that
$X=\Pi_{0}^{\sharp}(\mathrm{d}H_{0})=\Pi_{1}^{\sharp}(\mathrm{d}H_{1}).$
On any symplectic leaf where $\Pi_{0}$ is nondegenerate, one can introduce the recursion operator (see, for instance, [M.M1984, M.C.F+1997, F.M.P2000, Fernandes1994])
$N\mathrel{\vbox{\hbox{\scriptsize.}\hbox{\scriptsize.}}}=\Pi_{1}^{\sharp}\circ(\Pi_{0}^{\sharp})^{-1}.$
When the Nijenhuis torsion of $N$ vanishes, spectral invariants of $N$ yield hierarchies of commuting Hamiltonians, providing the standard bi-Hamiltonian framework for Liouville integrability [C.M.P1993, G.M.S2003]. This picture can be recast in the more general language of Poisson–Nijenhuis manifolds, Jacobi–Nijenhuis manifolds and their reductions [K.M1990, Wade2001, NunesdaCosta1998a, M.M.P1999, P.N2003, C.D.N2010, D.M.P2011], but in this section we restrict ourselves to the basic Poisson setting.
2.2 Exact symplectic manifolds and homogeneity
Throughout the paper we adopt the convention that an exact symplectic manifold $(M,\theta)$ carries the symplectic form $\omega\mathrel{\vbox{\hbox{\scriptsize.}\hbox{\scriptsize.}}}=-\mathrm{d}\theta$. Hamiltonian vector fields are defined by $\iota_{X_{H}}\omega=\mathrm{d}H$, and the Liouville vector field $\Delta$ is characterized by $\iota_{\Delta}\omega=-\theta$.
Let $(M,\theta)$ be an exact symplectic manifold with symplectic form $\omega=-\mathrm{d}\theta$. The one-form $\theta$ is called a symplectic potential.
Definition 2.3.
The Liouville vector field $\Delta$ is defined by $\iota_{\Delta}\omega=-\theta$.
A tensor field $A$ on $M$ is said to be homogeneous of degree $w\in\mathbb{R}$ (or simply $w$-homogeneous) if
$\mathcal{L}_{\Delta}A=w\,A.$
In particular, a function $H\in\mathscr{C}^{\infty}(M)$ is homogeneous of degree $1$ if $\mathcal{L}_{\Delta}H=H$. A triple $(M,\theta,H)$ with $H$ of degree $1$ is called a homogeneous Hamiltonian system.
Remark 2.4.
The symplectic potential $\theta$ and the symplectic form $\omega$ are $1$-homogeneous with respect to the Liouville vector field $\Delta$. Moreover, if $f\in\mathscr{C}^{\infty}(M)$ is a $1$-homogeneous function, then its Hamiltonian vector field $X_{f}$, defined by $\iota_{X_{f}}\omega=\mathrm{d}f$, is 0-homogeneous, i.e. $\mathcal{L}_{\Delta}X_{f}=0$. In other words, the Hamiltonian flow of a $1$-homogeneous function is invariant under the Liouville flow.
Following [C.d.L+2023, Lopez-Gordon2024], we say that a homogeneous Hamiltonian system is integrable if it admits a full set of independent, homogeneous first integrals in involution.
Definition 2.5.
A homogeneous integrable system $(M,\theta,H)$ consists of a $2n$-dimensional exact symplectic manifold $(M,\theta)$, a $1$-homogeneous Hamiltonian $H$, and functions $f_{1},\dots,f_{n}\in\mathscr{C}^{\infty}(M)$ such that:
  1. 1.
    Each $f_{i}$ is a first integral of $X_{H}$, and homogeneous of degree $1$;
  2. 2.
    The $f_{i}$ are in involution for the Poisson bracket induced by $\omega$;
  3. 3.
    The differentials $\mathrm{d}f_{1},\dots,\mathrm{d}f_{n}$ are independent on a dense open subset.
In our later construction, we will start from a Poisson (or Poisson–Lie) bi-Hamiltonian system and, when an exact symplectic realization is available, we will require the Hamiltonian and its integrals to be $1$-homogeneous with respect to the corresponding Liouville vector field. Compatibility of Poisson tensors is responsible for integrability in the Liouville sense, whereas homogeneity is what will allow us to relate this symplectic picture with a contact Hamiltonian system on a $(2n+1)$-dimensional manifold via symplectization. These are therefore two independent geometric requirements that must be arranged simultaneously in our applications.
Homogeneous integrable systems arise naturally as symplectizations of integrable contact systems, as recalled below.
2.3 Contact Hamiltonian systems and integrability
Let $(C,\alpha)$ be a $(2n+1)$-dimensional co-oriented contact manifold, i.e. $\alpha\wedge(\mathrm{d}\alpha)^{n}$ is a volume form on $C$. The Reeb vector field $R$ is uniquely determined by
$\iota_{R}\mathrm{d}\alpha=0,\qquad\alpha(R)=1.$
Given a function $h\in\mathscr{C}^{\infty}(C)$, the associated contact Hamiltonian vector field $X_{h}$ is defined by (see, e.g., [d.L2019b, G.G2022a, Lopez-Gordon2024])
$\begin{cases}\alpha(X_{h})=h,\\[5.69054pt] \iota_{X_{h}}\mathrm{d}\alpha=(\mathrm{d}h)(R)\,\alpha-\mathrm{d}h.\end{cases}$
The triple $(C,\alpha,h)$ is called a contact Hamiltonian system.
The contact form $\alpha$ induces a Jacobi structure $(\Lambda,R)$ on $C$ [Lichnerowicz1977a, Lichnerowicz1978, d.L.M+2003, G.U2003], and hence a Jacobi bracket on $\mathscr{C}^{\infty}(C)$ given by
$\{f,g\}_{\alpha}\mathrel{\vbox{\hbox{\scriptsize.}\hbox{\scriptsize.}}}=X_{f}(g)-g\,R(f)=\Lambda(\mathrm{d}f,\mathrm{d}g)+f\,R(g)-g\,R(f),\qquad f,g\in\mathscr{C}^{\infty}(C).$
Definition 2.6.
Let $(C,\alpha,h)$ be a contact Hamiltonian system. A function $f\in\mathscr{C}^{\infty}(C)$ is called a dissipated quantity (also called a dissipated integral) if it is in Jacobi-involution with $h$, i.e.
$\{f,h\}_{\alpha}=0.$
This notion was introduced in [G.G.M+2020a] and further developed in [d.L2019a, C.d.L+2023, Lopez-Gordon2024] as the natural replacement, in contact geometry, of conserved quantities in Hamiltonian dynamics.
Definition 2.7.
A $(2n+1)$-dimensional contact Hamiltonian system $(C,\alpha,h)$ is said to be completely integrable if there exist $n+1$ dissipated quantities $f_{1},\dots,f_{n+1}$ such that:
  1. 1.
    $\{f_{i},f_{j}\}_{\alpha}=0$ for all $i,j$;
  2. 2.
    the differentials $\mathrm{d}f_{1},\dots,\mathrm{d}f_{n+1}$ have rank at least $n$ on a dense open subset of $C$.
Reordering if necessary, we may and shall assume that $h$ itself is included in the family $\{f_{i}\}$.
In this language, the contact Liouville–Arnol’d theorem proved in [C.d.L+2023] extends the classical symplectic theory [Arnold1978, Audin2004] to the contact category. In particular, completely integrable contact systems admit local action–angle type coordinates adapted to the Reeb dynamics and to the dissipated quantities.
2.4 Symplectization and the contact–homogeneous correspondence
Let $(C,\alpha)$ be a co-oriented contact manifold. Its (trivial) symplectization is the exact symplectic manifold $(C^{\Sigma},\theta)$ with
$C^{\Sigma}\mathrel{\vbox{\hbox{\scriptsize.}\hbox{\scriptsize.}}}=C\times\mathbb{R}_{>0},\qquad\theta\mathrel{\vbox{\hbox{\scriptsize.}\hbox{\scriptsize.}}}=r\,\alpha,$
where $r$ is the coordinate on $\mathbb{R}_{>0}$. The projection
$\Sigma\colon C^{\Sigma}\to C,\qquad\Sigma(x,r)=x,$
is a symplectization in the sense of [D.L.M1991, I.L.M+1997, G.I.M+2004, Petalidou2002, Lopez-Gordon2024]: there exists a nowhere-vanishing function $\sigma$ (here $\sigma=r$) such that $\theta=\sigma\,\Sigma^{\ast}\alpha$.
Given a contact Hamiltonian $h\in\mathscr{C}^{\infty}(C)$, one associates a $1$-homogeneous Hamiltonian $H^{\Sigma}$ on $C^{\Sigma}$ by
$H^{\Sigma}(x,r)\mathrel{\vbox{\hbox{\scriptsize.}\hbox{\scriptsize.}}}=-\,\sigma(x,r)\,h(x)=-r\,h(x),$
following the conventions in [C.d.L+2023, Lopez-Gordon2024]. This defines a homogeneous Hamiltonian system $(C^{\Sigma},\theta,H^{\Sigma})$.
The following result summarizes the correspondence between integrability in the contact and homogeneous symplectic settings (see [C.d.L+2023, Lopez-Gordon2024, G.I.M+2004, Petalidou2002] for proofs and variants, and [NunesdaCosta1998a, M.M.P1999, P.N2003, C.D.N2010] for related Jacobi and Jacobi–Nijenhuis structures).
Proposition 2.8 (Contact–homogeneous correspondence).
Let $(C,\alpha,h)$ be a contact Hamiltonian system and $(C^{\Sigma},\theta,H^{\Sigma})$ its symplectization as above. Then:
  1. 1.
    The map $f\mapsto f^{\Sigma}\mathrel{\vbox{\hbox{\scriptsize.}\hbox{\scriptsize.}}}=-\sigma\,\Sigma^{\ast}f$ defines a bijection between $\mathscr{C}^{\infty}(C)$ and the space of $1$-homogeneous functions on $C^{\Sigma}$.
  2. 2.
    For all $f,g\in\mathscr{C}^{\infty}(C)$, the Poisson bracket on $(C^{\Sigma},\theta)$ and the Jacobi bracket on $(C,\alpha)$ satisfy
    $\{f^{\Sigma},g^{\Sigma}\}_{\theta}=\{f,g\}_{\alpha}^{\Sigma}.$
  3. 3.
    A function $f$ is a dissipated quantity for $(C,\alpha,h)$ if and only if $f^{\Sigma}$ is a first integral of $(C^{\Sigma},\theta,H^{\Sigma})$.
  4. 4.
    The contact system $(C,\alpha,h)$ is completely integrable if and only if the homogeneous Hamiltonian system $(C^{\Sigma},\theta,H^{\Sigma})$ is a homogeneous integrable system.
Remark 2.9.
The above result is classical. Items (1)–(3) follow from the standard identification between Jacobi manifolds and homogeneous Poisson manifolds under symplectization; see [G.I.M+2004, Petalidou2002, D.L.M1991]. The equivalence of contact integrability and homogeneous integrability in item (4) is established in [C.d.L+2023, Lopez-Gordon2024]. We record the proposition here for completeness and to fix notation for the subsequent sections, where this correspondence becomes the bridge to the quantum (Lindbladian) setting.
Remark 2.10.
The previous proposition shows that the study of completely integrable contact Hamiltonian systems can be reduced, via symplectization, to the study of homogeneous integrable systems on exact symplectic manifolds. In particular, one may try to import bi-Hamiltonian techniques to the contact setting by first constructing compatible Poisson structures on the symplectic realization and then projecting the resulting commuting integrals to the contact manifold.
In contrast, starting directly from compatible Jacobi structures on $(C,\alpha)$ [NunesdaCosta1998a, M.M.P1999, P.N2003, C.D.N2010] and using the associated Jacobi–Nijenhuis recursion operators does not, in general, lead to a maximal family of dissipated quantities in involution for a contact system, as we shall discuss later in connection with a no-go result of “Jacobi/bi-Hamiltonian” type. This dichotomy will be crucial when we pass from classical homogeneous and contact dynamics to quantum GKSL generators in the subsequent sections.
Example 2.11.
Let $C=\mathbb{R}^{3}(q,p,z)$ endowed with the standard contact form
[…]
Its Reeb vector field is $R=\partial_{z}$.
We adopt the convention for the contact Hamiltonian vector field $X_{h}$:
$\alpha(X_{h})=h,\qquad\iota_{X_{h}}\mathrm{d}\alpha=(\mathrm{d}h)(R)\,\alpha-\mathrm{d}h.$
For $\alpha=\mathrm{d}z-p\,\mathrm{d}q$ this yields the coordinate expressions
$\dot{q}=-\frac{\partial h}{\partial p},\qquad\dot{p}=\frac{\partial h}{\partial q}+p\frac{\partial h}{\partial z},\qquad\dot{z}=h-p\frac{\partial h}{\partial p}.$
Consider the linear contact Hamiltonian
$h(q,p,z)\mathrel{\vbox{\hbox{\scriptsize.}\hbox{\scriptsize.}}}=z-p.$
Then
$\frac{\partial h}{\partial p}=-1,\qquad\frac{\partial h}{\partial q}=0,\qquad\frac{\partial h}{\partial z}=1,$
and the resulting contact dynamics is
$\dot{q}=1,\qquad\dot{p}=p,\qquad\dot{z}=z.$
Hence the flow is explicitly solvable:
$q(t)=q_{0}+t,\qquad p(t)=p_{0}e^{t},\qquad z(t)=z_{0}e^{t}.$
Define the observable
$I_{1}(q,p,z)\mathrel{\vbox{\hbox{\scriptsize.}\hbox{\scriptsize.}}}=e^{-z}.$
Since $R(h)=1$, the condition for $I_{1}$ to be a dissipated quantity is
$X_{h}(I_{1})+(Rh)\,I_{1}=0.$
A direct computation gives
$X_{h}(I_{1})=-\,\frac{\partial I_{1}}{\partial z}=-(-e^{-z})=e^{-z},$
and therefore
$X_{h}(I_{1})+(Rh)\,I_{1}=e^{-z}-e^{-z}=0.$
Equivalently, in terms of the Jacobi bracket,
$\{I_{1},h\}_{\alpha}=\frac{\partial I_{1}}{\partial z}-\frac{\partial I_{1}}{\partial p}+I_{1}=(-e^{-z})-0+e^{-z}=0.$
Hence $I_{1}$ is a dissipated quantity in the sense of Definition 2.7.
Note that $I_{1}$ is not constant along the contact flow:
$X_{h}(I_{1})=e^{-z}\neq 0,$
which illustrates that dissipated quantities (Jacobi integrals) need not be conserved along the flow, unless additional conditions such as $R(I_{1})=0$ are satisfied.
Moreover, $I_{0}\mathrel{\vbox{\hbox{\scriptsize.}\hbox{\scriptsize.}}}=h$ is trivially dissipated since $\{h,h\}_{\alpha}=0$, and we have
$\{I_{0},I_{1}\}_{\alpha}=0.$
The differentials
$\mathrm{d}h=\mathrm{d}z-\mathrm{d}p,\qquad\mathrm{d}I_{1}=-e^{-z}\,\mathrm{d}z$
are linearly independent on a dense open subset of $C$, so
$\operatorname{rank}\,\langle\mathrm{d}h,\mathrm{d}I_{1}\rangle=2\geq n=1,$
as required for complete contact integrability in dimension $3$.
Therefore $(C,\alpha,h)$ is a completely integrable contact Hamiltonian system with $n=1$, admitting the two commuting dissipated quantities $I_{0}=h$ and $I_{1}=e^{-z}$.
Finally, this contact system arises from symplectization with our sign conventions. Let
[…]
and define the $1$-homogeneous Hamiltonian
$H^{\Sigma}\mathrel{\vbox{\hbox{\scriptsize.}\hbox{\scriptsize.}}}=-\,r\,h=-r(z-p).$
Then $(C^{\Sigma},\theta,H^{\Sigma})$ is an exact symplectic homogeneous Hamiltonian system, and the contact dynamics of $(C,\alpha,h)$ is recovered by restriction to the hypersurface $\{r=1\}$, in accordance with Proposition 2.8.
3 Bi-Hamiltonian Poisson–Lie systems and contact realizations
In this section we explain how bi-Hamiltonian structures on Poisson–Lie groups naturally give rise to contact integrable systems via exact symplectic realizations and homogeneous Hamiltonians. This provides the classical geometric backbone for the Lindblad constructions that will be developed later on.
3.1 Bi-Hamiltonian structures on Poisson–Lie groups
Let $G$ be a Lie group with multiplication $m\colon G\times G\to G$. Recall that a smooth map $\Phi\colon(M_{1},\pi_{1})\to(M_{2},\pi_{2})$ between Poisson manifolds is a Poisson map if it preserves Poisson brackets, i.e.
$\{f\circ\Phi,\,g\circ\Phi\}_{\pi_{1}}=\{f,g\}_{\pi_{2}}\circ\Phi,\qquad\forall\,f,g\in\mathscr{C}^{\infty}(M_{2}).$
A Poisson tensor $\pi_{0}\in\Gamma(\wedge^{2}{\mathsf{T}}G)$ is called multiplicative if the multiplication map is a Poisson map from $(G\times G,\pi_{0}\oplus\pi_{0})$ to $(G,\pi_{0})$, i.e.
$m\colon(G\times G,\pi_{0}\oplus\pi_{0})\longrightarrow(G,\pi_{0})$
is Poisson. In this case $(G,\pi_{0})$ is a Poisson–Lie group; see, e.g., [Lichnerowicz1977, Lichnerowicz1977a, Lichnerowicz1978, D.L.M1991, G.U2003].
Given a Poisson tensor $\pi$ on $G$, we denote by
$\pi^{\sharp}\colon{\mathsf{T}}^{\ast}G\longrightarrow{\mathsf{T}}G,\qquad\pi^{\sharp}(\alpha)\mathrel{\vbox{\hbox{\scriptsize.}\hbox{\scriptsize.}}}=\pi(\cdot,\alpha),$
the associated bundle map. For $H\in\mathscr{C}^{\infty}(G)$, the Hamiltonian vector field $X_{H}$ is defined by $X_{H}=\pi^{\sharp}(\mathrm{d}H)$.
On a symplectic manifold $(M,\omega)$ we define the Hamiltonian vector field $X_{H}$ by $\iota_{X_{H}}\omega=\mathrm{d}H$.
Definition 3.1.
A bi-Hamiltonian Poisson–Lie system on a Lie group $G$ consists of:
  • two Poisson–Lie structures $\pi_{0},\pi_{1}\in\Gamma(\wedge^{2}{\mathsf{T}}G)$ on $G$;
  • two Hamiltonians $H_{0},H_{1}\in\mathscr{C}^{\infty}(G)$;
such that:
  1. 1.
    $\pi_{0}$ and $\pi_{1}$ are compatible Poisson tensors in the sense that every linear combination
    $\pi_{\lambda}\mathrel{\vbox{\hbox{\scriptsize.}\hbox{\scriptsize.}}}=\pi_{1}-\lambda\,\pi_{0},\qquad\lambda\in\mathbb{R},$
    is a Poisson tensor on $G$;
  2. 2.
    the corresponding Hamiltonian vector fields coincide,
    $X\mathrel{\vbox{\hbox{\scriptsize.}\hbox{\scriptsize.}}}=\pi_{0}^{\sharp}(\mathrm{d}H_{0})=\pi_{1}^{\sharp}(\mathrm{d}H_{1}).$
Thus $(\pi_{0},\pi_{1})$ defines a Poisson pencil on $G$, and on each symplectic leaf of $\pi_{0}$ (or $\pi_{1}$) the restriction of $(\pi_{0},\pi_{1})$ and $X$ yields an ordinary bi-Hamiltonian system in the sense of Section 2. The general bi-Hamiltonian theory on Poisson manifolds (and in particular on symplectic leaves) is developed in [C.M.P1993, M.C.F+1997, F.M.P2000, Fernandes1994, G.M.S2003].
Many classical integrable systems, such as the Euler top and its deformations, admit realizations of this type on Poisson–Lie groups [M.P1996, NunesdaCosta1998, Petalidou2000]. In what follows, we shall assume that the bi-Hamiltonian dynamics is (Liouville) integrable on a suitable open dense subset of $G$, and concentrate on how this structure lifts to an exact symplectic manifold and then induces a contact integrable system via homogeneity.
3.2 Exact symplectic realizations and homogeneous lifts
Let $(G,\pi_{0})$ be a Poisson manifold. A symplectic realization of $(G,\pi_{0})$ is a symplectic manifold $(M,\omega)$ together with a Poisson map $J\colon(M,\omega)\to(G,\pi_{0})$ whose image is an open subset of $G$; see, for instance, [L.M1987, D.L.M1991, G.I.M+2004, Lopez-Gordon2024]. We are interested in the case in which $(M,\omega)$ is exact.
Definition 3.2.
An exact symplectic realization of a Poisson manifold $(G,\pi_{0})$ is a triple $(M,\theta,J)$ such that:
  1. 1.
    $\theta\in\Omega^{1}(M)$ is a one-form with $\omega\mathrel{\vbox{\hbox{\scriptsize.}\hbox{\scriptsize.}}}=-\mathrm{d}\theta$ symplectic;
  2. 2.
    $J\colon(M,\omega)\to(G,\pi_{0})$ is a Poisson map with open image, where $(M,\omega)$ is regarded as a Poisson manifold via the Poisson bivector $\pi$ induced by $\omega$, i.e. $\pi^{\sharp}=\omega^{-1}$.
The pair $(M,\theta)$ is then an exact symplectic manifold in the sense of Section 2.
Remark 3.3.
Here and throughout, when we say that $J\colon(M,\omega)\to(G,\pi_{0})$ is a Poisson map, we mean that $M$ is equipped with the Poisson structure $\pi=\omega^{-1}$ induced by the symplectic form $\omega$, so that $J$ satisfies $J_{*}\pi=\pi_{0}$.
Remark 3.4.
Requiring a symplectic realization $(M,\omega)$ to be exact, $\omega=-\mathrm{d}\theta$, imposes topological restrictions on the realizing manifold $M$ (in particular, $M$ cannot be compact and $\omega]=0$ in de Rham cohomology). This requirement, however, does not constrain the underlying Lie–Poisson or Poisson–Lie structure itself. For a large class of Lie–Poisson systems, including linear and solvable cases, canonical exact realizations are provided by cotangent bundles ${\mathsf{T}}^{\ast}Q$, which are always non-compact and carry a natural exact symplectic structure. The example in Section 3.4 illustrates this situation explicitly. More generally, since our construction only requires working on open dense subsets where the homogeneous integrals are independent, the assumption of exactness should be regarded as a natural geometric convenience rather than a substantive restriction of the Poisson–Lie framework.
Let $\Delta$ be the Liouville vector field of $(M,\theta)$, defined by
$\iota_{\Delta}\omega=-\theta.$
As recalled in Section 2, a function $H\in\mathscr{C}^{\infty}(M)$ is said to be $1$-homogeneous if
$\mathcal{L}_{\Delta}H=H,$
and a triple $(M,\theta,H)$ with $H$ 11-homogeneous is a homogeneous Hamiltonian system. We shall also use the notion of homogeneous integrability from Definition 2.3.
To connect the Poisson–Lie data on $G$ with contact geometry, we make the following structural assumption.
Assumption 3.5.
Let $(G,\pi_{0},\pi_{1};H_{0},H_{1})$ be a bi-Hamiltonian Poisson–Lie system whose Hamiltonian flow $X$ is completely integrable (in the Liouville sense) on a dense open subset of $G$. We assume that there exist:
  • an exact symplectic realization $(M,\theta,J)$ of $(G,\pi_{0})$, with $\dim M=2(n+1)$;
  • a $1$-homogeneous Hamiltonian $H\in\mathscr{C}^{\infty}(M)$ with respect to the Liouville vector field $\Delta$ of $(M,\theta)$,
such that:
  1. 1.
    $H_{0}\circ J=H$;
  2. 2.
    the Hamiltonian vector field $X_{H}$ on $(M,\omega=-\mathrm{d}\theta)$ projects to the bi-Hamiltonian vector field $X$ on $G$, i.e.
    $TJ\circ X_{H}=X\circ J;$
  3. 3.
    there exist $n+1$ independent, $1$-homogeneous first integrals in involution for $(M,\theta,H)$, making $(M,\theta,H)$ a homogeneous integrable system in the sense of Definition 2.3.
In many examples of interest, the Poisson pencil $(\pi_{0},\pi_{1})$ admits a symplectic realization $(M,\omega)$ on which the pullbacks of $\pi_{0},\pi_{1}$ give rise to a Poisson–Nijenhuis structure $(\omega,N)$ on $M$ [K.M1990, Wade2001, D.M.P2011, B.K.M2022, B.K.M2022a, B.K.M2024b, Konyaev2021]. The eigenvalues of $N$ then provide a natural candidate for homogeneous integrals in involution, and the homogeneity of $H$ can often be arranged by a suitable choice of exact symplectic structure (or by considering an appropriate homogeneous extension).
In any case, Assumption 3.5 encodes precisely the two independent geometric ingredients emphasised in the introduction:
  • the bi-Hamiltonian Poisson data on $G$ and its integrability;
  • the existence of a homogeneous Hamiltonian lift on an exact symplectic realization $(M,\theta)$, which will later allow us to pass to contact geometry via a transverse hypersurface.
3.3 Contact realizations from homogeneous lifts
Under Assumption 3.5, we can apply the general homogeneous–contact correspondence of Section 2 to obtain a contact integrable system. We spell this out carefully.
We first recall a standard fact about exact symplectic manifolds (see, e.g., [L.M1987, D.L.M1991, Lopez-Gordon2024]).
Lemma 3.6.
Let $(M,\theta)$ be an exact symplectic manifold of dimension $2(n+1)$, with $\omega=-\mathrm{d}\theta$ and Liouville vector field $\Delta$ defined by $\iota_{\Delta}\omega=-\theta$. Let $C\subset M$ be a hypersurface transverse to $\Delta$. Then the restriction $\alpha\mathrel{\vbox{\hbox{\scriptsize.}\hbox{\scriptsize.}}}=\theta|_{C}$ is a contact form on $C$, i.e.
$\alpha\wedge(\mathrm{d}\alpha)^{n}\neq 0$
everywhere on $C$.
Proof.
Since $\omega$ is symplectic, the top form $\omega^{n+1}$ is a volume form on $M$. Using $\iota_{\Delta}\omega=-\theta$ and the identity $\iota_{\Delta}(\omega^{n+1})=(n+1)(\iota_{\Delta}\omega)\wedge\omega^{n}$, we obtain
$\iota_{\Delta}(\omega^{n+1})=-(n+1)\,\theta\wedge\omega^{n}=-(n+1)(-1)^{n}\,\theta\wedge(\mathrm{d}\theta)^{n}.$
Hence $\Theta\mathrel{\vbox{\hbox{\scriptsize.}\hbox{\scriptsize.}}}=\theta\wedge(\mathrm{d}\theta)^{n}$ is nonvanishing at every point where $\Delta\neq 0$.
Now let $C\subset M$ be a hypersurface transverse to $\Delta$. Transversality implies $\Delta_{x}\neq 0$ and ${\mathsf{T}}_{x}M=\mathbb{R}\Delta_{x}\oplus{\mathsf{T}}_{x}C$ for every $x\in C$. Therefore the contraction $\iota_{\Delta}(\omega^{n+1})$ restricts to a nowhere-vanishing $(2n+1)$-form on $C$, and so does $\Theta$.
Finally, letting $\iota:C\hookrightarrow M$ be the inclusion and $\alpha=\iota^{*}\theta$, we have
$\iota^{*}\Theta=\iota^{*}\bigl(\theta\wedge(\mathrm{d}\theta)^{n}\bigr)=\alpha\wedge(\mathrm{d}\alpha)^{n},$
which is therefore nowhere vanishing on $C$. This proves that $\alpha$ is a contact form. ∎
We now combine this lemma with the homogeneous integrability of $(M,\theta,H)$.
Proposition 3.7.
Let $(G,\pi_{0},\pi_{1};H_{0},H_{1})$ and $(M,\theta,H)$ be as in Assumption 3.5. Let $\Delta$ be the Liouville vector field of $(M,\theta)$ and let $C\subset M$ be a hypersurface transverse to $\Delta$, contained in the open set where $H\neq 0$. Define
$\alpha\mathrel{\vbox{\hbox{\scriptsize.}\hbox{\scriptsize.}}}=\theta|_{C},\qquad h\mathrel{\vbox{\hbox{\scriptsize.}\hbox{\scriptsize.}}}=H|_{C}.$
Then $(C,\alpha,h)$ is a $(2n+1)$-dimensional completely integrable contact Hamiltonian system in the sense of Definition 2.4.
Proof.
First, by Lemma 3.6, the restriction $\alpha=\theta|_{C}$ is a contact form on $C$, so $(C,\alpha)$ is a co-oriented contact manifold.
Since $H$ is $1$-homogeneous with respect to $\Delta$, we have $\mathcal{L}_{\Delta}H=H$. Writing $\varphi_{s}$ for the flow of $\Delta$, this implies
$\frac{\mathrm{d}}{\mathrm{d}s}H(\varphi_{s}(x))=(\mathcal{L}_{\Delta}H)(\varphi_{s}(x))=H(\varphi_{s}(x)),$
so, on any integral curve of $\Delta$,
$H(\varphi_{s}(x))=e^{s}H(x).$
In particular, on the open region where $H\neq 0$, the level sets of $H$ are transversal to the Liouville flow, and every Liouville orbit intersects each nonzero level set exactly once. By construction $C$ is a hypersurface transverse to $\Delta$ and contained in $\{H\neq 0\}$, so every Liouville orbit in the region of interest intersects $C$ exactly once.
By Assumption 3.5, the homogeneous Hamiltonian system $(M,\theta,H)$ is a homogeneous integrable system in the sense of Definition 2.3: there exist $n+1$ first integrals
$F_{0}\mathrel{\vbox{\hbox{\scriptsize.}\hbox{\scriptsize.}}}=H,\;F_{1},\dots,F_{n}\in\mathscr{C}^{\infty}(M)$
such that:
  1. 1.
    each $F_{j}$ is $1$-homogeneous, $\mathcal{L}_{\Delta}F_{j}=F_{j}$;
  2. 2.
    the $F_{j}$ are in involution with respect to the Poisson bracket induced by $\omega=-\mathrm{d}\theta$;
  3. 3.
    the differentials $\mathrm{d}F_{0},\dots,\mathrm{d}F_{n}$ are independent on a dense open subset $U\subset M$.
Consider the restrictions
$I_{j}\mathrel{\vbox{\hbox{\scriptsize.}\hbox{\scriptsize.}}}=F_{j}|_{C}\in\mathscr{C}^{\infty}(C),\qquad j=0,\dots,n.$
We claim that the $(n+1)$-tuple $(I_{0},\dots,I_{n})$ defines a completely integrable contact system on $(C,\alpha)$ with contact Hamiltonian $h=I_{0}$.
(i) Dissipated quantities and Jacobi-involution. Let $X_{H}$ be the Hamiltonian vector field of $H$ on $(M,\omega)$, and let $X_{h}$ be the contact Hamiltonian vector field of $h$ on $(C,\alpha)$, defined by
$\alpha(X_{h})=h,\qquad\iota_{X_{h}}\mathrm{d}\alpha=(\mathrm{d}h)(R)\,\alpha-\mathrm{d}h,$
where $R$ is the Reeb vector field of $(C,\alpha)$.
Standard results on the contactization of homogeneous Hamiltonian systems (see, for instance, [D.L.M1991, I.L.M+1997, G.I.M+2004, Petalidou2002, Lopez-Gordon2024]) ensure that:
  • Since $H$ is $1$-homogeneous, one has $\mathcal{L}_{\Delta}H=H$, and hence $X_{H}$ is 0-homogeneous (equivalently $\Delta,X_{H}]=0$). In particular, the flow of $X_{H}$ commutes with the Liouville flow.
  • Because $C$ is transverse to $\Delta$, a neighborhood of $C$ is identified with an open subset of $\mathbb{R}\times C$ via the Liouville flow. The $\Delta$-invariance of $X_{H}$ implies that $X_{H}$ descends to a well-defined vector field on the local quotient by $\Delta$, and choosing the section $i:C\hookrightarrow M$ yields an induced vector field on $C$.
  • Under this identification, the induced vector field on $C$ coincides with the contact Hamiltonian vector field $X_{h}$ associated with $\alpha=\theta|_{C}$ and $h=H|_{C}$. Moreover, if $F$ is a $1$-homogeneous first integral of $X_{H}$, then $I\mathrel{\vbox{\hbox{\scriptsize.}\hbox{\scriptsize.}}}=F|_{C}$ satisfies $\{I,h\}_{\alpha}=0$.
This is the same mechanism as in Proposition 2.8, but now the contact manifold $C$ is realized as a transversal hypersurface to the Liouville flow in an arbitrary exact symplectic manifold $(M,\theta)$, rather than as the slice $r=1$ of a trivial symplectization.
Applying this to $F_{0},\dots,F_{n}$ we see that each $I_{j}$ is a dissipated quantity for the contact Hamiltonian system $(C,\alpha,h)$, i.e. $\{I_{j},h\}_{\alpha}=0$ for all $j$.
Moreover, since the $F_{j}$ are in Poisson-involution on $M$ and the Jacobi bracket on $C$ is obtained by reduction along the Liouville flow (again, see [D.L.M1991, G.I.M+2004, Petalidou2002, Lopez-Gordon2024]), the restrictions $I_{j}$ are in Jacobi-involution:
$\{I_{j},I_{k}\}_{\alpha}=0,\qquad\forall\,j,k\in\{0,\dots,n\}.$
(ii) Functional independence. Let $U\subset M$ be the dense open subset where $\mathrm{d}F_{0},\dots,\mathrm{d}F_{n}$ are independent. Since $C$ is transverse to the Liouville flow, the intersection $U\cap C$ is a dense open subset of $C$. At any $x\in U\cap C$ the differentials $\mathrm{d}F_{0},\dots,\mathrm{d}F_{n}$ are linearly independent in ${\mathsf{T}}^{\ast}_{x}M$. Restricting to ${\mathsf{T}}_{x}C\subset{\mathsf{T}}_{x}M$ we obtain the differentials $\mathrm{d}I_{0},\dots,\mathrm{d}I_{n}$, which remain linearly independent on ${\mathsf{T}}_{x}C$: otherwise there would be a nontrivial linear relation among the $\mathrm{d}F_{j}$ on ${\mathsf{T}}_{x}M$. Thus,
$\operatorname{rank}\langle\mathrm{d}I_{0},\dots,\mathrm{d}I_{n}\rangle\geq n+1$
on a dense open subset of $C$. In particular, the rank is at least $n$, as required in Definition 2.4 of contact integrability.
(iii) Dimension and completeness. By construction $\dim M=2(n+1)$ and $C$ is a hypersurface, so $\dim C=2n+1$. Together with (i) and (ii), this shows that the contact Hamiltonian system $(C,\alpha,h)$ is a $(2n+1)$-dimensional completely integrable contact system in the sense of Definition 2.4. ∎
In particular, any bi-Hamiltonian Poisson–Lie system satisfying Assumption 3.5 yields, in a canonical way, a contact integrable system $(C,\alpha,h)$ equipped with $n+1$ independent dissipated quantities in Jacobi-involution, obtained from the bi-Hamiltonian hierarchy on $G$ via an exact, homogeneous lift to $(M,\theta,H)$ and restriction to a Liouville-transverse hypersurface.
Remark 3.8.
The construction above crucially uses Poisson compatibility and homogeneity on the exact symplectic realization $(M,\theta)$, rather than compatibility of Jacobi structures directly on the contact manifold $(C,\alpha)$. As shown in our previous work on the Jacobi/bi-Hamiltonian no-go phenomenon [C.d.L+2023, Lopez-Gordon2024] (building on [NunesdaCosta1998a, M.M.P1999, N.P2002, C.D.N2010, P.N2003, Petalidou2002]), compatible Jacobi pairs on a contact manifold do not provide enough independent eigenvalues of the associated recursion operator to recover a full set of contact integrals. By passing through the homogeneous Poisson picture on $(M,\theta)$ we circumvent this obstruction: the bi-Hamiltonian structure lives in the Poisson category, while the contact dynamics is obtained by projection along the Liouville flow.
[…]
Figure 1: From a bi-Hamiltonian Poisson–Lie system on $(G,\pi_{0},\pi_{1})$ to a contact integrable system $(C,\alpha,h)$ via an exact, homogeneous symplectic realization $(M,\theta,H)$ and a Liouville-transverse hypersurface $C$.
This is precisely the classical geometric framework in which, in the next section, we shall construct Lindblad generators whose semiclassical limit reproduces the contact flow of $h$ and whose conserved quantum observables quantize the classical dissipated quantities $I_{0},\dots,I_{n}$.
Corollary 3.9.
Let $(G,\pi_{0},\pi_{1};H_{0},H_{1})$ be a bi-Hamiltonian Poisson–Lie system satisfying Assumption 3.5. Then there exist a Liouville-transverse hypersurface $C\subset M$ and a contact form $\alpha=\theta|_{C}$ such that the induced contact Hamiltonian system $(C,\alpha,h)$, with $h=H|_{C}$, is completely integrable and carries $n+1$ dissipated quantities $I_{0}\mathrel{\vbox{\hbox{\scriptsize.}\hbox{\scriptsize.}}}=h,I_{1},\dots,I_{n}$ in Jacobi-involution.
Proof.
This is an immediate consequence of Proposition 3.7, together with the fact that Liouville-transverse hypersurfaces can be chosen inside any given nonzero level set of $H$ in the region where the homogeneous integrals $(F_{0},\dots,F_{n})$ are independent. ∎
3.4 Example: a Poisson–Nijenhuis bi-Hamiltonian system with two nontrivial contact integrals
We illustrate Proposition 3.7 with the Poisson–Nijenhuis example already discussed in Section 2, now interpreted as a homogeneous bi-Hamiltonian system producing a contact integrable system with the correct number of dissipated quantities for the $3$-dimensional ($n=1$) contact setting.
Let $\mathrm{Aff}^{+}(1)$ denote the connected component of the identity of the group of orientation–preserving affine transformations of the real line, with group law $(a,b)\cdot(a^{\prime},b^{\prime})=(aa^{\prime},\,b+a\,b^{\prime})$. Its Lie algebra $\mathfrak{aff}(1)$ has basis $(e_{1},e_{2})$ with $e_{1},e_{2}]=e_{2}$. For our purposes it suffices to view the construction below as a concrete Poisson–Nijenhuis model that may be realized as coming from a Poisson–Lie pencil on a $2$-dimensional solvable group; the explicit realization map will not be needed.
Consider the cotangent bundle
$M={\mathsf{T}}^{\ast}\mathbb{R}^{2}\cong\mathbb{R}^{4}$
with coordinates $(x^{1},x^{2},p_{1},p_{2})$ and canonical one-form $\theta=p_{i}\,\mathrm{d}x^{i}$. Throughout we use the convention $\omega=-\mathrm{d}\theta$. The associated Poisson tensor is
$\Lambda=\frac{\partial}{\partial x^{1}}\wedge\frac{\partial}{\partial p_{1}}+\frac{\partial}{\partial x^{2}}\wedge\frac{\partial}{\partial p_{2}}.$
Define a second Poisson tensor
$\Lambda_{1}=p_{1}\,\frac{\partial}{\partial x^{1}}\wedge\frac{\partial}{\partial p_{1}}+p_{2}x^{2}\,\frac{\partial}{\partial x^{2}}\wedge\frac{\partial}{\partial p_{2}}.$
One checks that $\Lambda,\Lambda_{1}]_{\mathrm{SN}}=0$, hence $(\Lambda,\Lambda_{1})$ is a Poisson pencil.
On the open dense subset
$U\mathrel{\vbox{\hbox{\scriptsize.}\hbox{\scriptsize.}}}=M\setminus\big(\{p_{2}=0\}\cap\{x^{2}=0\}\big),$
the recursion operator
$N\mathrel{\vbox{\hbox{\scriptsize.}\hbox{\scriptsize.}}}=\sharp_{\Lambda_{1}}\circ\sharp_{\Lambda}^{-1}$
has two eigenvalues
$\lambda_{1}\mathrel{\vbox{\hbox{\scriptsize.}\hbox{\scriptsize.}}}=p_{1},\qquad\lambda_{2}\mathrel{\vbox{\hbox{\scriptsize.}\hbox{\scriptsize.}}}=p_{2}x^{2},$
which are functionally independent on $U$ and pairwise commuting for the two Poisson structures.
The Liouville vector field
$\Delta=p_{i}\frac{\partial}{\partial p_{i}}$
satisfies $\iota_{\Delta}\omega=-\theta$, and $\lambda_{1},\lambda_{2}$ are both $1$-homogeneous:
$\mathcal{L}_{\Delta}\lambda_{i}=\lambda_{i},\qquad i=1,2.$
Define the $1$-homogeneous Hamiltonian
$H\mathrel{\vbox{\hbox{\scriptsize.}\hbox{\scriptsize.}}}=\lambda_{1}+\lambda_{2}=p_{1}+p_{2}x^{2}.$
Then $\mathcal{L}_{\Delta}H=H$ and, since $\{\lambda_{1},\lambda_{2}\}_{\Lambda}=0$, each $\lambda_{i}$ is a first integral of $X_{H}$. Thus $(M,\theta,H)$ is a $1$-homogeneous integrable system in the sense of Definition 2.3.
Moreover, $X_{H}$ is bi-Hamiltonian:
$X_{H}=\Lambda^{\sharp}(\mathrm{d}H)=\Lambda_{1}^{\sharp}(\mathrm{d}H_{1}),$
for instance with
$H_{1}=\log(\lambda_{1}\lambda_{2})\quad\text{on the region where }\lambda_{1}\lambda_{2}>0\quad(\text{or }H_{1}=\log|\lambda_{1}\lambda_{2}|\text{ locally}).$
Introduce coordinates
$x^{1}=q,\qquad x^{2}=z,\qquad p_{1}=-rp,\qquad p_{2}=r,\qquad r>0.$
Then
$\theta=p_{i}\,\mathrm{d}x^{i}=r\,\mathrm{d}z-rp\,\mathrm{d}q=r(\mathrm{d}z-p\,\mathrm{d}q),\qquad H=r(z-p),\qquad\Delta=r\,\partial_{r}.$
In particular, $(q^{i},p_{i})$ define coordinates on hypersurfaces which are transversal to the Liouville vector field. Let
$C\mathrel{\vbox{\hbox{\scriptsize.}\hbox{\scriptsize.}}}=\{r=1\}\subset M.$
Since $\Delta=r\partial_{r}$ is transverse to the level sets of $r$, the hypersurface $C$ is transverse to the Liouville flow. The restriction
$\alpha\mathrel{\vbox{\hbox{\scriptsize.}\hbox{\scriptsize.}}}=\theta|_{C}=\mathrm{d}z-p\,\mathrm{d}q$
is a contact form, since
$\alpha\wedge\mathrm{d}\alpha=(\mathrm{d}z-p\,\mathrm{d}q)\wedge(-\mathrm{d}p\wedge\mathrm{d}q)\neq 0.$
Thus $(C,\alpha)$ is a $3$-dimensional co-oriented contact manifold ($n=1$), and
$h\mathrel{\vbox{\hbox{\scriptsize.}\hbox{\scriptsize.}}}=H|_{C}=z-p$
is the induced contact Hamiltonian.
The eigenvalues restrict to
$\bar{\lambda}_{1}=\lambda_{1}|_{C}=-p,\qquad\bar{\lambda}_{2}=\lambda_{2}|_{C}=z.$
Up to a harmless sign convention we set
$I_{1}\mathrel{\vbox{\hbox{\scriptsize.}\hbox{\scriptsize.}}}=p,\qquad I_{2}\mathrel{\vbox{\hbox{\scriptsize.}\hbox{\scriptsize.}}}=z.$
By Proposition 2.8,
$\{I_{j},h\}_{\alpha}=0,\qquad\{I_{1},I_{2}\}_{\alpha}=0,$
so $I_{1}$ and $I_{2}$ are dissipated quantities in Jacobi-involution. Since
$I_{0}\mathrel{\vbox{\hbox{\scriptsize.}\hbox{\scriptsize.}}}=h=z-p=I_{2}-I_{1},$
the family $\{I_{0},I_{1},I_{2}\}$ contains exactly $n+1=2$ functionally independent integrals for $n=1$; indeed,
$\operatorname{rank}\langle\mathrm{d}I_{0},\mathrm{d}I_{1},\mathrm{d}I_{2}\rangle=2\geq n=1$
on a dense open subset (in fact everywhere). Therefore $(C,\alpha,h)$ is a completely integrable contact system in the sense of Definition 2.4.
Remark 3.10 (The $3$-dimensional case $n=1$).
For a contact manifold of dimension $2n+1=3$, complete contact integrability requires $n+1=2$ independent dissipated quantities in Jacobi-involution. One may take the contact Hamiltonian itself as one of them; any additional Jacobi integrals must then satisfy functional relations, as happens here with $I_{0}=I_{2}-I_{1}$. What matters is that the rank condition in Definition 2.4 holds (here with rank $2$ on a dense set). This example fits the general scheme of Proposition 3.7.
4 Lindblad dynamics compatible with contact geometry
We now move to the quantum setting and introduce a class of Lindblad generators that are compatible with the contact integrable systems constructed in Section 3. For definiteness, we work with finite-dimensional Hilbert spaces; in this context, $\mathcal{B}(\mathcal{H})$ denotes the full $C^{\ast}$-algebra of bounded operators on $\mathcal{H}$ (which coincides with the space of all linear operators), all operator topologies coincide, and domain issues do not arise. Most of the algebraic definitions and arguments below extend to separable infinite-dimensional Hilbert spaces under standard technical assumptions on domains and continuity of the corresponding quantum dynamical semigroups.
Before proceeding, we briefly recall the algebraic notions that will be used throughout this subsection:
  • An algebra $\mathcal{A}$ over $\mathbb{C}$ is a complex vector space equipped with an associative bilinear product. A -algebra is an algebra endowed with an involution $A\mapsto A^{\dagger}$ satisfying $(AB)^{\dagger}=B^{\dagger}A^{\dagger}$.
  • A subalgebra $\mathcal{A}\subset\mathcal{B}(\mathcal{H})$ is a linear subspace that is closed under products and adjoints. It is called unital if it contains the identity operator $\mathbb{I}_{\mathcal{H}}$, and commutative (or abelian) if $AB=BA$ for all $A,B\in\mathcal{A}$.
  • A $C^{\ast}$-algebra is a -algebra $\mathcal{A}$ that is complete with respect to a norm $\|\cdot\|$ satisfying the $C^{\ast}$-identity $\|A^{\dagger}A\|=\|A\|^{2}$. In finite-dimensional Hilbert spaces, every -subalgebra of $\mathcal{B}(\mathcal{H})$ is automatically a $C^{\ast}$-algebra with the operator norm.
  • In particular, a unital commutative $C^{\ast}$-subalgebra of $\mathcal{B}(\mathcal{H})$ can always be identified, via the spectral theorem, with an algebra of functions on a finite classical configuration space. This interpretation will be crucial for understanding the emergence of effectively classical observables and decoherence in open quantum systems.
In the finite-dimensional setting, a state is described by a density matrix $\rho$. Fixing an orthonormal basis $\{|j\rangle\}_{j=1}^{d}$ (often selected by the measurement scheme or by the spectral decomposition of a reference Hamiltonian), we write $\rho_{jk}=\langle j|\rho|k\rangle$. The coherences are the off-diagonal entries $\rho_{jk}$ with $j\neq k$, which encode relative phases and interference effects, while the diagonal entries $\rho_{jj}$ represent populations, i.e., classical probabilities of outcomes in that basis (in the standard terminology of open quantum systems, see e.g., [BreuerPetruccione]). A dynamics exhibits decoherence (in that basis) when the off-diagonal entries are damped in time, typically driving $\rho(t)$ towards a state that is approximately diagonal in the chosen basis. In open quantum systems, decoherence is not an ad hoc ingredient but a robust consequence of the coupling to an environment: while the total evolution on system-plus-environment may be unitary, the reduced dynamics on the system alone is generally non-unitary and may suppress coherences.
There are two equivalent ways to represent the same quantum evolution. It is convenient to introduce a unified notation for expectation values. For an observable $A$ at time $t$, we define
$\mathrm{E}(A)(t)\mathrel{\vbox{\hbox{\scriptsize.}\hbox{\scriptsize.}}}=\operatorname{Tr}\big(\rho(t)A\big)$
in the Schrödinger picture, or equivalently
$\mathrm{E}(A)(t)\mathrel{\vbox{\hbox{\scriptsize.}\hbox{\scriptsize.}}}=\operatorname{Tr}\big(\rho_{0}A(t)\big)$
in the Heisenberg picture.
In the Schrödinger picture, the state evolves according to
$\dot{\rho}(t)=\mathcal{L}(\rho(t)),\qquad\rho(0)=\rho_{0},$
while observables $A_{0}$ are fixed. Expectation values are then given by
$\mathrm{E}(A_{0})(t)=\operatorname{Tr}\big(\rho(t)A_{0}\big),$
where $\mathcal{L}$ is the Gorini–Kossakowski–Sudarshan–Lindblad generator (or simply Lindblad generator).
In the Heisenberg picture, the state remains fixed and observables evolve as
$\dot{A}(t)=\mathcal{L}^{\dagger}(A(t)),\qquad A(0)=A_{0},$
where $\mathcal{L}^{\dagger}$ denotes the adjoint of $\mathcal{L}$ with respect to the trace pairing, i.e.
$\operatorname{Tr}\big(\mathcal{L}(\rho)\,A\big)=\operatorname{Tr}\big(\rho\,\mathcal{L}^{\dagger}(A)\big).$
Expectation values are given by
$\mathrm{E}(A)(t)=\operatorname{Tr}\big(\rho_{0}A(t)\big).$
The two descriptions are equivalent by construction, since they yield identical expectation values $\mathrm{E}(A)(t)$ for all observables and all times.
We will systematically use the Heisenberg picture in what follows. The reason is that our compatibility conditions with contact geometry are most naturally expressed as invariance statements for observables, namely
$\mathcal{L}^{\dagger}(\widehat{I})=0$
for the quantized dissipated quantities $\widehat{I}$, and as a semiclassical generator limit
$\mathcal{L}_{\hbar}^{\dagger}(\widehat{f}_{\hbar})\;\longrightarrow\;Q_{\hbar}(\{f,h\}_{\alpha})$
on the commutative subalgebra generated by the integrals. This viewpoint allows one to identify conserved quantities and invariant subalgebras directly at the level of the observable algebra $\mathcal{B}(\mathcal{H})$, which is particularly well suited to the geometric framework developed here.
By contrast, the Schrödinger picture is often preferable when addressing state-preparation problems, entropy production, contractivity properties of the quantum dynamical semigroup, or the detailed long-time behavior of the density operator $\rho(t)$; see, for instance, [Lindblad1976, BreuerPetruccione] and the references therein. Both pictures are, of course, mathematically equivalent, but emphasize different structural aspects of open quantum dynamics.
4.1 GKSL generators and Heisenberg picture
Let $\mathcal{H}$ be a finite-dimensional Hilbert space and $\mathcal{B}(\mathcal{H})$ the $C^{\ast}$-algebra of linear operators on $\mathcal{H}$. We denote by $\mathcal{S}(\mathcal{H})$ the convex set of density matrices on $\mathcal{H}$, i.e. positive trace-one operators.
Definition 4.1.
[…]
$\mathcal{L}(\rho)=-\frac{i}{\hbar}[H,\rho]+\sum_{k=1}^{N}\Big(L_{k}\rho L_{k}^{\dagger}-\tfrac{1}{2}\{L_{k}^{\dagger}L_{k},\rho\}\Big),$ (4.1)
where $H=H^{\dagger}$ is self-adjoint (the Hamiltonian), $\{L_{k}\}_{k=1}^{N}\subset\mathcal{B}(\mathcal{H})$ is a finite family of bounded operators (the Lindblad operators), with $N\in\mathbb{N}$, and $\{\cdot,\cdot\}$ denotes the anticommutator {A,B} . . $\{A,B\}\mathrel{\vbox{\hbox{\scriptsize.}\hbox{\scriptsize.}}}=AB$+BA.
It is well known that such $\mathcal{L}$ are precisely the generators of one-parameter semigroups $\Phi_{t}=e^{t\mathcal{L}},\,t\geq 0$, of completely positive and trace-preserving (CPTP) maps on $\mathcal{S}(\mathcal{H})$. Conversely, any norm-continuous CPTP semigroup on $\mathcal{S}(\mathcal{H})$ has a generator of the form (4.1) (see [BreuerPetruccione], [Lindblad1976] for instance).
We now recall the Heisenberg picture associated with a Lindblad generator.
Definition 4.2.
Given a Lindblad generator $\mathcal{L}$ acting on density matrices, the Heisenberg adjoint $\mathcal{L}^{\dagger}$ is the linear map on $\mathcal{B}(\mathcal{H})$ defined by
$\mathrm{Tr}\big(\mathcal{L}(\rho)\,A\big)=\mathrm{Tr}\big(\rho\,\mathcal{L}^{\dagger}(A)\big)$
for all $\rho\in\mathcal{B}(\mathcal{H})$ and all $A\in\mathcal{B}(\mathcal{H})$.
Since the trace pairing $\langle\rho,A\rangle=\mathrm{Tr}(\rho A)$ is nondegenerate, $\mathcal{L}^{\dagger}$ is uniquely determined by this relation.
Proposition 4.3.
Let $\mathcal{L}$ be of GKSL form (4.1). Then its Heisenberg adjoint $\mathcal{L}^{\dagger}$ is given, for all $A\in\mathcal{B}(\mathcal{H})$, by
$\mathcal{L}^{\dagger}(A)=\frac{i}{\hbar}[H,A]+\sum_{k=1}^{N}\Big(L_{k}^{\dagger}AL_{k}-\tfrac{1}{2}\{L_{k}^{\dagger}L_{k},A\}\Big).$ (4.2)
Proof.
Let $A\in\mathcal{B}(\mathcal{H})$ and $\rho\in\mathcal{B}(\mathcal{H})$. Using the cyclicity of the trace and the definition (4.1), we compute:
$\mathrm{Tr}\big(\mathcal{L}(\rho)\,A\big)$ $=-\frac{i}{\hbar}\mathrm{Tr}\big([H,\rho]\,A\big)+\sum_{k=1}^{N}\mathrm{Tr}\Big(L_{k}\rho L_{k}^{\dagger}A-\tfrac{1}{2}\{L_{k}^{\dagger}L_{k},\rho\}A\Big)$
$=-\frac{i}{\hbar}\mathrm{Tr}\big(H\rho A-\rho HA\big)+\sum_{k=1}^{N}\Big(\mathrm{Tr}(L_{k}\rho L_{k}^{\dagger}A)-\tfrac{1}{2}\mathrm{Tr}(L_{k}^{\dagger}L_{k}\rho A)-\tfrac{1}{2}\mathrm{Tr}(\rho L_{k}^{\dagger}L_{k}A)\Big).$
By cyclicity of the trace,
$\mathrm{Tr}(H\rho A)=\mathrm{Tr}(\rho AH),\qquad\mathrm{Tr}(L_{k}\rho L_{k}^{\dagger}A)=\mathrm{Tr}(\rho L_{k}^{\dagger}AL_{k}),\qquad\mathrm{Tr}(L_{k}^{\dagger}L_{k}\rho A)=\mathrm{Tr}(\rho AL_{k}^{\dagger}L_{k}),$
so we obtain
$\mathrm{Tr}\big(\mathcal{L}(\rho)\,A\big)=\frac{i}{\hbar}\mathrm{Tr}\big(\rho[H,A]\big)+\sum_{k=1}^{N}\Big(\mathrm{Tr}(\rho L_{k}^{\dagger}AL_{k})-\tfrac{1}{2}\mathrm{Tr}(\rho AL_{k}^{\dagger}L_{k})-\tfrac{1}{2}\mathrm{Tr}(\rho L_{k}^{\dagger}L_{k}A)\Big).$
Factoring out $\rho$ under the trace gives
$\mathrm{Tr}\big(\mathcal{L}(\rho)\,A\big)=\mathrm{Tr}\Big(\rho\Big[\frac{i}{\hbar}[H,A]+\sum_{k=1}^{N}\big(L_{k}^{\dagger}AL_{k}-\tfrac{1}{2}\{L_{k}^{\dagger}L_{k},A\}\big)\Big]\Big),$
and, by the defining relation of $\mathcal{L}^{\dagger}$ and the nondegeneracy of the trace pairing, this shows (4.2). ∎
We now formalise the notion of conserved observable for a given Lindblad dynamics.
Proposition 4.4.
Let $\mathcal{L}$ be a Lindblad generator with adjoint $\mathcal{L}^{\dagger}$, and let $O\in\mathcal{B}(\mathcal{H})$. The following statements are equivalent:
  1. 1.
    $O$ has a constant expectation value, i.e.,
    $\mathrm{Tr}\big(\rho(t)\,O\big)=\mathrm{Tr}\big(\rho(0)\,O\big)\,$,
    for each solution $\rho(t)$ of $\dot{\rho}=\mathcal{L}(\rho)$,
  2. 2.
    for any solution $\rho(t)$ of $\dot{\rho}=\mathcal{L}(\rho)$ one has
    $\frac{\mathrm{d}}{\mathrm{d}t}\,\mathrm{Tr}\big(\rho(t)\,O\big)=0\,$,
  3. 3.
    $O\in\ker\mathcal{L}^{\dagger}$.
Proof.
The equivalence between the first two statements is obvious. For proving the equivalence between the second and third statements, note that, in the Schrödinger picture,
$\frac{\mathrm{d}}{\mathrm{d}t}\,\mathrm{Tr}(\rho(t)\,O)=\mathrm{Tr}(\dot{\rho}(t)\,O)=\mathrm{Tr}(\mathcal{L}(\rho(t))\,O)=\mathrm{Tr}(\rho(t)\,\mathcal{L}^{\dagger}(O))\,.$
Since the trace product is non-degenerate, the right-hand side vanishes for every $\rho$ if and only if $\mathcal{L}^{\dagger}(O)$. ∎
This motivates the following definition:
Definition 4.5.
Let $\mathcal{L}$ be a Lindblad generator and $\mathcal{L}^{\dagger}$ its Heisenberg adjoint. An operator $O\in\mathcal{B}(\mathcal{H})$ is called a (quantum) constant of motion [^1] This terminology is standard in the theory of quantum dynamical semigroups, where such observables are precisely those whose expectation values are preserved under the evolution for all initial states. for $\mathcal{L}$ if it satisfies one (and hence all) of the equivalence statements in Proposition 4.4.
Remark 4.6.
For any Lindblad operator, every matrix proportional to the identity is a constant of the motion, since $\mathbb{I}\in\ker\mathcal{L}^{\dagger}$. Hence, constants of the motion always form an unital subalgebra of $\mathcal{B}(\mathcal{H})$, of dimension at least $1$.
We now give a sufficient condition ensuring that an observable is conserved by the Lindblad dynamics.
Lemma 4.7.
Let $\mathcal{L}$ be a Lindblad generator in GKSL form (4.1), and let $O\in\mathcal{B}(\mathcal{H})$ be self-adjoint. Assume that
$H,O]=0,\qquad[L_{k},O]=0,\qquad[L_{k}^{\dagger},O]=0\quad\text{for all }k.$
Then $\mathcal{L}^{\dagger}(O)=0$, i.e. $O$ is a quantum constant of motion.
Proof.
Using (4.2),
$\mathcal{L}^{\dagger}(O)=\frac{i}{\hbar}[H,O]+\sum_{k=1}^{N}\Big(L_{k}^{\dagger}OL_{k}-\tfrac{1}{2}\{L_{k}^{\dagger}L_{k},O\}\Big).$
By hypothesis, $[H,O]=0$, so the Hamiltonian term vanishes. For each $k$,
$L_{k}^{\dagger}OL_{k}=OL_{k}^{\dagger}L_{k},$
since $L_{k},O]=[L_{k}^{\dagger},O]=0$ implies $O$ commutes with both $L_{k}$ and $L_{k}^{\dagger}$, and hence with $L_{k}^{\dagger}L_{k}$. Similarly,
$\{L_{k}^{\dagger}L_{k},O\}=L_{k}^{\dagger}L_{k}O+OL_{k}^{\dagger}L_{k}=2OL_{k}^{\dagger}L_{k}.$
Therefore
$L_{k}^{\dagger}OL_{k}-\tfrac{1}{2}\{L_{k}^{\dagger}L_{k},O\}=OL_{k}^{\dagger}L_{k}-OL_{k}^{\dagger}L_{k}=0.$
Summing over $k$ we obtain $\mathcal{L}^{\dagger}(O)=0$. ∎
Remark 4.8.
Lemma 4.7 shows that any observable $O$ commuting with the Hamiltonian and with all Lindblad operators (and their adjoints) is invariant under the Heisenberg evolution.
In particular, let $\{O_{0},\dots,O_{n}\}$ be a family of commuting observables generating a commutative $C^{\ast}$-subalgebra
$\mathcal{A}\mathrel{\vbox{\hbox{\scriptsize.}\hbox{\scriptsize.}}}=C^{\ast}(O_{0},\dots,O_{n})\subset\mathcal{B}(\mathcal{H}),$
and denote by
[…]
its commutant. If the Hamiltonian $H$ and all Lindblad operators $L_{k}$ belong to $\mathcal{A}^{\prime}$, then every observable $O\in\mathcal{A}$ satisfies
$\mathcal{L}^{\dagger}(O)=0,$
and is therefore a constant of motion for the Lindblad dynamics.
This will be our basic mechanism to quantize the classical dissipated quantities of a contact integrable system: we will construct Lindblad generators whose Hamiltonian and dissipative parts lie in the commutant of the algebra generated by the quantum integrals. $\diamond$
Example 4.9 (dephasing Lindblad dynamics with conserved observables).
We conclude this subsection with a simple but paradigmatic example in dimension $2$, which illustrates how genuine quantum dissipation may coexist with nontrivial constants of motion. This example will serve as a quantum toy model for the notion of “dissipated” yet conserved quantities that later appear as quantizations of contact integrals.
Let $\mathcal{H}=\mathbb{C}^{2}$ and denote by $\sigma_{x},\sigma_{y},\sigma_{z}$ the Pauli matrices, which form a basis of the space of traceless self-adjoint operators on $\mathcal{H}$. Moreover, $\{\sigma_{x},\sigma_{y},\sigma_{z},\mathbb{I}\}$ is a basis of $\mathcal{B}(\mathcal{H})$. Physically, $\sigma_{z}$ is diagonal in the energy basis and measures the population imbalance between the two energy levels. We consider the Lindblad generator $\mathcal{L}$ defined by (4.1) with
$H\mathrel{\vbox{\hbox{\scriptsize.}\hbox{\scriptsize.}}}=\frac{\hbar\omega}{2}\,\sigma_{z},\qquad L\mathrel{\vbox{\hbox{\scriptsize.}\hbox{\scriptsize.}}}=\sqrt{\gamma}\,\sigma_{z},$
where $\omega\in\mathbb{R}$ is the energy splitting and $\gamma>0$ is a dissipation rate. There is a single Lindblad operator $L_{1}=L$.
For any density matrix $\rho$, the Schrödinger-picture evolution reads
$\mathcal{L}(\rho)=-\frac{i}{\hbar}[H,\rho]+L\rho L^{\dagger}-\tfrac{1}{2}\{L^{\dagger}L,\rho\}.$
Since $\sigma_{z}^{2}=\mathbb{I}$, where $\mathbb{I}$ denotes the identity operator on $\mathcal{H}$, we have
$L^{\dagger}L=\gamma\sigma_{z}^{2}=\gamma\mathbb{I}.$
As a consequence, the dissipative part simplifies to
$L\rho L^{\dagger}-\tfrac{1}{2}\{L^{\dagger}L,\rho\}=\gamma(\sigma_{z}\rho\sigma_{z}-\rho),$
and therefore
$\mathcal{L}(\rho)=-\frac{i\omega}{2}[\sigma_{z},\rho]+\gamma(\sigma_{z}\rho\sigma_{z}-\rho).$
This dynamics is known as pure dephasing: it suppresses the off-diagonal matrix elements of $\rho$ in the eigenbasis of $\sigma_{z}$ (i.e. quantum coherences) while leaving the diagonal entries (the populations) unchanged. Although dissipation is present, no energy exchange with the environment occurs.
While dissipation is most naturally described in the Schrödinger picture at the level of states, the analysis of conserved quantities is more transparent in the Heisenberg picture. The adjoint generator acts as
$\mathcal{L}^{\dagger}(A)=\frac{i}{\hbar}[H,A]+L^{\dagger}AL-\tfrac{1}{2}\{L^{\dagger}L,A\}=\frac{i\omega}{2}[\sigma_{z},A]+\gamma(\sigma_{z}A\sigma_{z}-A).$
We now determine the constants of motion. Take $O=\sigma_{z}$. Then
$H,\sigma_{z}]=0,\qquad[L,\sigma_{z}]=\sqrt{\gamma}[\sigma_{z},\sigma_{z}]=0,$
so the hypotheses of Lemma 4.7 are satisfied and $\mathcal{L}^{\dagger}(\sigma_{z})=0$. Hence, $\sigma_{z}$ is a quantum constant of the motion. Consequently, the population difference between the two energy levels is conserved, even though the system undergoes irreversible dephasing.
In contrast, for $O=\sigma_{x}$ or $O=\sigma_{y}$ one has
$H,O]\neq 0,\qquad[L,O]\neq 0,$
and a direct computation yields
$\mathcal{L}^{\dagger}(\sigma_{x})=-\gamma\sigma_{x}-\omega\sigma_{y},\qquad\mathcal{L}^{\dagger}(\sigma_{y})=-\gamma\sigma_{y}+\omega\sigma_{x}.$
Hence neither $\sigma_{x}$ nor $\sigma_{y}$ is a constant of motion: their expectation values decay (and rotate) in time, reflecting the loss of quantum coherence.
Finally, recall that the identity operator $\mathbb{I}$ always satisfies $\mathcal{L}^{\dagger}(\mathbb{I})=0$, expressing trace preservation of the Lindblad dynamics. Therefore, this system has a nontrivial subalgebra $\operatorname{span}\{\sigma_{z},\mathbb{I}\}\subset\mathcal{B}(\mathcal{H})$ of conserved observables, despite the presence of irreversible dissipation.
4.2 Invariant commutative subalgebras and constants of motion
We now highlight a basic structural feature of GKSL dynamics that will play a central role in our construction. In the following sections, we will require that invariant commutative subalgebras arise as quantizations of Jacobi-commutative algebras of dissipated quantities associated with contact integrable systems.
In the Heisenberg picture, the generator $\mathcal{L}^{\dagger}$ defines a generally non-Hamiltonian evolution on the algebra of observables $\mathcal{B}(\mathcal{H})$, in the sense that it is not a derivation generated by a commutator with a self-adjoint operator, but includes dissipative terms of Lindblad type. As a consequence, the resulting dynamics is not an automorphism of the observable algebra and may exhibit contraction and loss of information.
Such invariant commutative subalgebras encode families of observables whose expectation values are preserved by the open-system dynamics and which therefore behave as robust, effectively classical quantities in the presence of dissipation. As we show below, the GKSL structure provides a simple and natural mechanism ensuring the existence of such subalgebras under elementary commutation assumptions.
Theorem 4.10 (Invariant commutative subalgebras for GKSL dynamics).
Let $\mathcal{H}$ be a finite-dimensional Hilbert space and let $\mathcal{L}$ be a Lindblad generator on $\mathcal{B}(\mathcal{H})$ of the form (4.1), with Heisenberg adjoint $\mathcal{L}^{\dagger}$ given by (4.2). Let $\mathcal{A}\subset\mathcal{B}(\mathcal{H})$ be a unital, commutative $C^{\ast}$-subalgebra. Suppose that
$H,A]=0,\qquad[L_{k},A]=[L_{k}^{\dagger},A]=0\quad\text{for all }A\in\mathcal{A}\text{ and all }k.$
Then, $\mathcal{A}\subseteq\ker\mathcal{L}^{\dagger}$ and every observable $A\in\mathcal{A}$ is a quantum constant of motion. In particular, $\mathcal{A}$ is invariant under the Heisenberg evolution generated by $\mathcal{L}^{\dagger}$, i.e. $\mathcal{L}^{\dagger}(\mathcal{A})\subset\mathcal{A}$.
Proof.
Let $A\in\mathcal{A}$. By assumption $[H,A]=0$, so the Hamiltonian contribution to $\mathcal{L}^{\dagger}(A)$ vanishes. For each Lindblad operator $L_{k}$, the commutation relations imply
$L_{k}^{\dagger}AL_{k}=AL_{k}^{\dagger}L_{k},\qquad\{L_{k}^{\dagger}L_{k},A\}=2AL_{k}^{\dagger}L_{k}.$
Hence each dissipative term satisfies
$L_{k}^{\dagger}AL_{k}-\tfrac{1}{2}\{L_{k}^{\dagger}L_{k},A\}=0.$
Summing over $k$ yields $\mathcal{L}^{\dagger}(A)=0$. Thus, $A$ is a constant of the motion. Since $\mathcal{A}$ is a linear subspace, this shows both invariance and pointwise vanishing of $\mathcal{L}^{\dagger}$ on $\mathcal{A}$. ∎
Remark 4.11.
The existence of an invariant commutative $C^{\ast}$-subalgebra for a Lindblad generator admits a natural physical interpretation in terms of coherence and decoherence.
In the Schrödinger picture, decoherence refers to the irreversible suppression of off-diagonal matrix elements of the density operator in certain preferred bases, typically induced by the dissipative part of the GKSL generator. Equivalently, in the Heisenberg picture, decoherence manifests itself as the contraction of the algebra of observables toward a smaller, dynamically stable subalgebra.
Invariant commutative subalgebras play a distinguished role in this process. Observables belonging to such a subalgebra have constant expectation values for all initial states and are insensitive to the dissipative dynamics. In this sense, they behave as robust, effectively classical quantities selected by the open-system evolution.
From a geometric viewpoint, this selection mechanism parallels the emergence of dissipated quantities in contact Hamiltonian systems: although the full dynamics is non-Hamiltonian and generically dissipative, certain observables retain a privileged status encoded by algebraic invariance conditions. This analogy will be made precise in the following sections by requiring that the invariant quantum observables coincide, in the semiclassical limit, with the dissipated quantities of a completely integrable contact system. $\diamond$
For the purposes of this work, the Heisenberg picture is particularly natural, since conserved observables are precisely characterized as elements of $\ker\mathcal{L}^{\dagger}$, and invariant commutative subalgebras can be directly identified. $\diamond$
Example 4.12 (Invariant commutative subalgebras in dephasing Lindblad dynamics).
The dephasing example above (Example 4.9) can be reinterpreted in purely algebraic terms as the invariance of a commutative subalgebra under the Heisenberg evolution generated by a GKSL operator. This perspective is particularly useful for understanding the role of conserved observables in open quantum systems and anticipates the contact-compatible constructions developed later.
Consider again the dephasing Lindblad generator on $\mathcal{H}=\mathbb{C}^{2}$ defined by
$\mathcal{L}^{\dagger}(A)=\frac{i\omega}{2}[\sigma_{z},A]+\gamma(\sigma_{z}A\sigma_{z}-A),\qquad\omega\in\mathbb{R},\ \gamma>0.$
Let
$\mathcal{A}\mathrel{\vbox{\hbox{\scriptsize.}\hbox{\scriptsize.}}}=\mathrm{span}\{\mathbb{I},\sigma_{z}\}\;\subset\;\mathcal{B}(\mathcal{H})$
be the unital -subalgebra generated by $\sigma_{z}$. This is a commutative $C^{\ast}$-subalgebra, canonically isomorphic to $\mathbb{C}^{2}$, corresponding to the algebra of functions on a two-point classical configuration space.
A direct computation shows that $\mathcal{A}$ is invariant under $\mathcal{L}^{\dagger}$. Indeed,
$\mathcal{L}^{\dagger}(\mathbb{I})=0,\qquad\mathcal{L}^{\dagger}(\sigma_{z})=0,$
and therefore
$\mathcal{L}^{\dagger}(A)=0\quad\text{for all }A\in\mathcal{A}.$
Equivalently,
$\mathcal{L}^{\dagger}(\mathcal{A})\subset\mathcal{A},\qquad\mathcal{L}^{\dagger}|_{\mathcal{A}}=0.$
In contrast, the complementary subspace
$\mathcal{A}^{\perp}\mathrel{\vbox{\hbox{\scriptsize.}\hbox{\scriptsize.}}}=\mathrm{span}\{\sigma_{x},\sigma_{y}\}$
is not invariant. Using the Pauli-matrix identities
$\sigma_{z},\sigma_{x}]=2i\sigma_{y},\qquad[\sigma_{z},\sigma_{y}]=-2i\sigma_{x},\qquad\sigma_{z}\sigma_{x}\sigma_{z}=-\sigma_{x},\qquad\sigma_{z}\sigma_{y}\sigma_{z}=-\sigma_{y},$
one finds
$\mathcal{L}^{\dagger}(\sigma_{x})=-\omega\sigma_{y}-2\gamma\sigma_{x},\qquad\mathcal{L}^{\dagger}(\sigma_{y})=\omega\sigma_{x}-2\gamma\sigma_{y}.$
Thus, writing a generic observable in $\mathcal{A}^{\perp}$ as $A(t)=a_{x}(t)\sigma_{x}+a_{y}(t)\sigma_{y},$ the Heisenberg equation $\dot{A}(t)=\mathcal{L}^{\dagger}(A(t))$ reduces to the linear system
$\binom{\dot{a}_{x}}{\dot{a}_{y}}=\begin{pmatrix}-2\gamma&\ \omega\\ -\omega&-2\gamma\end{pmatrix}\binom{a_{x}}{a_{y}}$.
The eigenvalues of this matrix are $\lambda_{\pm}=-2\gamma\pm i\omega$, and the solution is explicitly
$\binom{a_{x}(t)}{a_{y}(t)}=e^{-2\gamma t}\begin{pmatrix}\cos(\omega t)&\ \sin(\omega t)\\ -\sin(\omega t)&\ \cos(\omega t)\end{pmatrix}\binom{a_{x}(0)}{a_{y}(0)}$.
Equivalently, the evolved observables are
$\sigma_{x}(t)=e^{-2\gamma t}\big(\cos(\omega t)\,\sigma_{x}-\sin(\omega t)\,\sigma_{y}\big),\qquad\sigma_{y}(t)=e^{-2\gamma t}\big(\sin(\omega t)\,\sigma_{x}+\cos(\omega t)\,\sigma_{y}\big).$
For any fixed initial state $\rho_{0}$, the corresponding expectation values satisfy
$\mathrm{E}(\sigma_{x})(t)$ . . $\mathrel{\vbox{\hbox{\scriptsize.}\hbox{\scriptsize.}}}=\mathrm{Tr}(\rho_{0}\,\sigma_{x}(t))=e^{-2\gamma t}\Big(\cos(\omega t)\mathrm{E}(\sigma_{x})(0)-\sin(\omega t)\mathrm{E}(\sigma_{y})(0)\Big),$
$\mathrm{E}(\sigma_{y})(t)$ $=e^{-2\gamma t}\Big(\sin(\omega t)\mathrm{E}(\sigma_{x})(0)+\cos(\omega t)\mathrm{E}(\sigma_{y})(0)\Big).$
In particular,
$\mathrm{E}(\sigma_{x})(t)^{2}+\mathrm{E}(\sigma_{y})(t)^{2}=e^{-4\gamma t}\Big(\mathrm{E}(\sigma_{x})(0)^{2}+\mathrm{E}(\sigma_{y})(0)^{2}\Big),$
showing an exponential decay of the coherences with rate $2\gamma$.
The statement that $\mathcal{A}^{\perp}$ is contracted by the dissipative part” should be understood at the level of the induced semigroup $e^{t\mathcal{L}^{\dagger}}$, not in terms of a negative operator norm. Indeed, the restriction of $\mathcal{L}^{\dagger}$ to $\mathcal{A}^{\perp}$ has spectrum $\{-2\gamma\pm i\omega\}$, and hence
$\|e^{t\mathcal{L}^{\dagger}}|_{\mathcal{A}^{\perp}}\|\;\leq\;C\,e^{-2\gamma t}$
for some $C>0$. This exponential bound is the precise sense in which the dissipative dynamics suppresses the noncommutative (coherence) sector while leaving $\mathcal{A}$ pointwise invariant.
Here, the statement that $\mathcal{A}$ is pointwise invariant” means that the Heisenberg generator vanishes identically on $\mathcal{A}$, namely
$\mathcal{L}^{\dagger}(A)=0\qquad\text{for all }A\in\mathcal{A}.$
Equivalently, every observable in $\mathcal{A}$ is a fixed point of the quantum dynamical semigroup generated by $\mathcal{L}^{\dagger}$, that is,
$e^{t\mathcal{L}^{\dagger}}(A)=A\qquad\text{for all }t\geq 0,\ A\in\mathcal{A}.$
In particular,
$\mathcal{A}\subset\ker\mathcal{L}^{\dagger},$
which is a stronger property than mere invariance of $\mathcal{A}$ as a subspace.
From an algebraic viewpoint, the dephasing dynamics induces a dynamical splitting
$\mathcal{B}(\mathcal{H})=\mathcal{A}\;\oplus\;\mathcal{A}^{\perp},$
where $\mathcal{A}$ is a maximal commutative $C^{\ast}$-subalgebra that is invariant under the Heisenberg evolution generated by $\mathcal{L}^{\dagger}$, and $\mathcal{A}^{\perp}$ is its orthogonal complement with respect to the Hilbert–Schmidt inner product.
The key point is that $\mathcal{L}^{\dagger}$ acts trivially on $\mathcal{A}$,
$\mathcal{L}^{\dagger}|_{\mathcal{A}}=0,$
while it generates a strict contraction semigroup on $\mathcal{A}^{\perp}$. Consequently, for any observable $A\in\mathcal{B}(\mathcal{H})$ one has
$A(t)=e^{t\mathcal{L}^{\dagger}}(A)\;\longrightarrow\;\Pi_{\mathcal{A}}(A)\qquad(t\to+\infty),$
where $\Pi_{\mathcal{A}}$ denotes the conditional expectation (orthogonal projection) onto $\mathcal{A}$.
Thus, in the long-time limit, the Heisenberg evolution effectively reduces the observable algebra to $\mathcal{A}$: all components orthogonal to $\mathcal{A}$ are dynamically suppressed, while observables in $\mathcal{A}$ remain unchanged.
In the present example, $\mathcal{A}=C^{\ast}(\sigma_{z})$, so the asymptotic observable algebra consists precisely of functions of $\sigma_{z}$. This means that the Lindblad dynamics dynamically selects $\sigma_{z}$ as the only nontrivial observable whose expectation values remain stable under the open-system evolution.
More generally, this example illustrates a robust structural phenomenon: for suitable GKSL generators, there exist nontrivial commutative $C^{\ast}$-subalgebras $\mathcal{A}\subset\mathcal{B}(\mathcal{H})$, with $\mathcal{A}\neq\operatorname{span}(\mathbb{I})$, that are invariant under the Heisenberg evolution and lie entirely in $\ker\mathcal{L}^{\dagger}$.
In the geometric framework developed in this paper, such invariant commutative subalgebras provide the natural quantum counterparts of the Jacobi-commutative algebras generated by dissipated quantities in contact integrable systems.
Remark 4.13.
The terms “projection onto a classical observable algebra” and “selection of a preferred observable” are to be understood in the precise algebraic sense described above: the dynamics induces a contraction of the full observable algebra toward an invariant commutative subalgebra, without invoking additional postulates beyond the GKSL evolution itself. $\diamond$
4.3 Contact-compatible Lindblad generators
The goal of this subsection is to introduce a notion of Lindblad dynamics that is compatible, in a precise semiclassical sense, with a given completely integrable contact Hamiltonian system. More precisely, we want to formalize when a quantum open-system evolution preserves, at the level of expectation values and in the semiclassical limit, the distinguished family of dissipated quantities associated with contact integrability.
Let $(C,\alpha,h)$ be a completely integrable contact Hamiltonian system with dissipated quantities $I_{0}\mathrel{\vbox{\hbox{\scriptsize.}\hbox{\scriptsize.}}}=h,I_{1},\dots,I_{n}$ as in Definition 2.7 and Proposition 3.7. Throughout this subsection we assume that $(C,\alpha,h;I_{0},\dots,I_{n})$ arises from a bi-Hamiltonian Poisson–Lie system as described in Section 3.
Consider the algebra $(\mathscr{C}^{\infty}(C),\{\cdot,\cdot\}_{\alpha})$, where $\{\cdot,\cdot\}_{\alpha}$ denotes de Jacobi bracket defined by $\alpha$. Let $\mathcal{A}_{\mathrm{cl}}\subset\mathscr{C}^{\infty}(C)$ denote the commutative subalgebra generated by $\{I_{0},\dots,I_{n}\}$. We refer to $\mathcal{A}_{\mathrm{cl}}$ as the Jacobi commutative algebra of the integrable system. This terminology reflects the fact that $\mathcal{A}_{\mathrm{cl}}$ plays, for contact dynamics, the same role as the Poisson-commutative algebra of integrals in Liouville-integrable Hamiltonian systems: it encodes the distinguished observables whose mutual relations are governed by the underlying Jacobi structure.
We shall work with an abstract semiclassical quantization of the contact data. The reason for doing so is twofold. First, the geometric constructions leading to contact integrable systems are largely independent of any particular choice of quantization scheme. Second, our aim is to identify structural conditions on Lindblad generators that are robust under different realizations (e.g. Weyl quantization, Toeplitz quantization, or finite-dimensional representations), rather than tied to a specific model.
Definition 4.14 (Semiclassical quantization of the Jacobi commutative algebra).
A semiclassical quantization $(\mathcal{H}_{\hbar},Q_{\hbar})$ of $(C,\alpha,h;I_{0},\dots,I_{n})$ consists of a family of Hilbert spaces $\{\mathcal{H}_{\hbar}\}_{\hbar\in(0,\hbar_{0}]}$ and a family of linear maps $Q_{\hbar}\colon\mathcal{A}_{\mathrm{cl}}\longrightarrow\mathcal{B}(\mathcal{H}_{\hbar})$, such that for all $f,g\in\mathcal{A}_{\mathrm{cl}}$ one has:
  1. 1.
    $Q_{\hbar}(\overline{f})=Q_{\hbar}(f)^{\dagger}.$
    In particular, real-valued classical observables are mapped to self-adjoint operators.
  2. 2.
    $Q_{\hbar}(1)=\mathbf{1}_{\mathcal{H}_{\hbar}}.$
    Here $1$ denotes the constant function on $C$, while $\mathbf{1}_{\mathcal{H}_{\hbar}}$ is the identity operator on the Hilbert space $\mathcal{H}_{\hbar}$. This condition ensures that the quantization preserves units and expectation values of constant observables.
  3. 3.
    (Dirac condition for the Jacobi bracket)
    $\lim_{\hbar\to 0}\left\|\frac{1}{i\hbar}\big[Q_{\hbar}(f),Q_{\hbar}(g)\big]-Q_{\hbar}\big(\{f,g\}_{\alpha}\big)\right\|=0.$
    The factor $1/(i\hbar)$ reflects the standard correspondence principle: commutators of quantum observables rescaled by $1/(i\hbar)$ converge, in the semiclassical limit, to the classical Jacobi bracket. This condition is the natural contact analog of the usual Dirac condition in symplectic quantization.
For brevity we set
[…]
$\mathcal{L}_{\hbar}(\rho)=-\frac{i}{\hbar}\big[\widehat{H}_{\hbar},\rho\big]+\sum_{k=1}^{N(\hbar)}\big(L_{k,\hbar}\,\rho\,L_{k,\hbar}^{\dagger}-\tfrac{1}{2}\{L_{k,\hbar}^{\dagger}L_{k,\hbar},\rho\}\big).$
Here $N(\hbar)\in\mathbb{N}$ denotes the (possibly $\hbar$-dependent) number of Lindblad operators. Allowing $N(\hbar)$ to vary reflects the fact that the effective dissipative description of an open quantum system may depend on the semiclassical scale. We denote by $\mathcal{L}_{\hbar}^{\dagger}$ the Heisenberg adjoint of $\mathcal{L}_{\hbar}$ (see Proposition 4.3).
Remark 4.15.
A contact system is naturally encoded by its symplectization. Given a contact manifold $(C,\alpha)$, its symplectization is
$(M,\omega)\mathrel{\vbox{\hbox{\scriptsize.}\hbox{\scriptsize.}}}=(C\times\mathbb{R}_{>0},\,\omega\mathrel{\vbox{\hbox{\scriptsize.}\hbox{\scriptsize.}}}=d(r\alpha)),$
with Liouville (Euler) vector field $\Delta=r\partial_{r}$. Any function $f\in\mathscr{C}^{\infty}(C)$ admits a canonical 11–homogeneous lift on $M$, for instance
$f^{\Sigma}(x,r)\mathrel{\vbox{\hbox{\scriptsize.}\hbox{\scriptsize.}}}=r\,f(x),$
and the Jacobi bracket on $(C,\alpha)$ is recovered from the Poisson bracket on $(M,\omega)$ by restriction to the hypersurface $\{r=1\}$:
$\{f,g\}_{\alpha}\;=\;\{f^{\Sigma},g^{\Sigma}\}_{\omega}\big|_{r=1}.$
In particular, the contact Hamiltonian dynamics on $C$ is the reduction (or, equivalently, the restriction to a transverse section $r=\mathrm{const}$) of the homogeneous Hamiltonian dynamics on $M$.
In the present work, the semiclassical quantization map $Q_{\hbar}$ on the Jacobi commutative algebra $\mathcal{A}_{\mathrm{cl}}$ is precisely obtained by symplectizing first and then applying a semiclassical quantization on $M$ to the homogeneous lifts:
$Q_{\hbar}(f)\;\mathrel{\vbox{\hbox{\scriptsize.}\hbox{\scriptsize.}}}=\;\operatorname{Op}_{\hbar}(f^{\Sigma}),\qquad f\in\mathcal{A}_{\mathrm{cl}}.$
With this choice, “symplectization” and “quantization” commute at the level relevant for our compatibility conditions: the Dirac/Jacobi correspondence and the Egorov-type limit are inherited from the standard semiclassical calculus on $(M,\omega)$ (up to the usual $o(1)$ or $\mathcal{O}(\hbar)$ remainders).
It is also worth mentioning that the underlying contact dynamics depends only on the conformal class of the contact form. If $f$ is nowhere vanishing and $\alpha^{\prime}\mathrel{\vbox{\hbox{\scriptsize.}\hbox{\scriptsize.}}}=\alpha/f$, then the rescaled data $(C,\alpha^{\prime},\,fI_{0},\dots,fI_{n})$ describe the same integrable geometry (same invariant foliations and contact vector fields, $X_{I_{i}}^{\alpha}=X_{fI_{i}}^{\alpha^{\prime}}$). In the symplectization picture this corresponds to a change of the radial normalization/section, and therefore one expects the resulting quantizations to be equivalent (up to unitary equivalence and/or time reparametrization within the same semiclassical class). This “gauge” freedom is the classical counterpart of the inevitable unital sector in GKSL dynamics, since $\mathcal{L}_{\hbar}^{\dagger}(\mathbf{1})=0$ holds for every Lindblad generator. $\diamond$
We can now formulate the central compatibility notion.
Definition 4.16 (Contact-compatible Lindblad generator).
Let $(C,\alpha,h;I_{0},\dots,I_{n})$ and $(\mathcal{H}_{\hbar},Q_{\hbar})$ be as above. A family of Lindblad generators $\{\mathcal{L}_{\hbar}\}_{\hbar\in(0,\hbar_{0}]}$ is said to be contact-compatible if:
  1. (CC1)
    (Quantum constants of motion) For each $j=0,\dots,n$,
    $\mathcal{L}_{\hbar}^{\dagger}\big(\widehat{I}_{j,\hbar}\big)=0.$
  2. (CC2)
    (Semiclassical contact limit) For every $f\in\mathcal{A}_{\mathrm{cl}}$,
    $\lim_{\hbar\to 0}\left\|\mathcal{L}_{\hbar}^{\dagger}\big(\widehat{f}_{\hbar}\big)-Q_{\hbar}\big(\{f,h\}_{\alpha}\big)\right\|=0.$
Remark 4.17.
Condition (CC2), together with the Dirac condition in Definition 4.14, can be viewed as a weak Egorov-type property for contact dynamics: the Heisenberg evolution of quantum observables converges, in the semiclassical limit, to the classical Jacobi evolution generated by $h$. This provides the physical interpretation of contact-compatible Lindblad generators as quantum open-system evolutions whose classical shadow is the contact Hamiltonian flow. $\diamond$
Remark 4.18.
A conceptual difference between the classical contact setting and its quantum counterpart concerns the role of the unit observable.
On the classical side, the Jacobi algebra $(\mathscr{C}^{\infty}(C),\{\cdot,\cdot\}_{\alpha})$ is not unital in the Poisson sense: although the constant function $1$ belongs to the algebra, it is not central in general, since
$\{1,f\}_{\alpha}=-R(f),$
where $R$ is the Reeb vector field. Consequently, the inclusion of $1$ in a Jacobi-commutative subalgebra $\mathcal{A}_{\mathrm{cl}}$ imposes a genuine geometric condition, namely the invariance of all generators under the Reeb flow.
By contrast, in the quantum setting the identity operator $\mathbf{1}$ is automatically present and invariant under any GKSL evolution:
$\mathcal{L}^{\dagger}(\mathbf{1})=0$
holds identically for every Lindblad generator. As a result, the quantum notion of an invariant unital commutative $C^{\ast}$-subalgebra does not impose any additional constraint associated with the unit.
For this reason, the correct correspondence between classical and quantum structures should be understood as follows: the quantization map $Q_{\hbar}$ assigns operators to the nonconstant generators of the Jacobi-commutative algebra $\mathcal{A}_{\mathrm{cl}}$, while the unit is added automatically on the quantum side. The nontrivial content of contact compatibility and of quantum constants of motion therefore lies entirely in the quantization of the dissipated quantities, not in the unit element.
This distinction explains why, in minimal examples, the classical condition $\{1,I\}_{\alpha}=0$ has no direct quantum analog, and clarifies the precise sense in which invariant commutative $C^{\ast}$-subalgebras quantize Jacobi-commutative algebras of dissipated quantities. $\diamond$
Example: minimal contact-compatible algebra.
Let $\mathcal{A}_{\mathrm{cl}}=\mathrm{span}\{I\}$ be the Jacobi-commutative subalgebra generated by a single dissipated quantity $I$. Any semiclassical quantization $Q_{\hbar}$ maps the nontrivial generator $I$ to a self-adjoint operator $\widehat{I}_{\hbar}$. The corresponding quantum algebra is then taken to be the unital $C^{\ast}$-algebra
$\mathcal{A}_{\hbar}\mathrel{\vbox{\hbox{\scriptsize.}\hbox{\scriptsize.}}}=\mathrm{span}\{\mathbf{1},\widehat{I}_{\hbar}\}$,
where the identity operator is included automatically.
A contact-compatible Lindblad generator is precisely one whose Heisenberg adjoint leaves $\mathcal{A}_{\hbar}$ pointwise invariant and satisfies $\mathcal{L}^{\dagger}(\widehat{I}_{\hbar})=0$. In this minimal setting, contact compatibility reduces to the existence of a nontrivial invariant commutative subalgebra generated by a single dissipated quantity beyond the scalars. This example highlights a general structural feature: while the constant function $1$ plays a nontrivial role in the Jacobi algebra at the classical level, its quantum counterpart is automatically invariant. The content of contact compatibility therefore lies entirely in the quantization of the nonconstant dissipated quantities.
For a set of functions $S\subseteq\mathscr{C}^{\infty}(C)$, we write $\mathrm{Jacobi\text{-}alg}\langle S\rangle$ for the smallest Jacobi subalgebra of $(\mathscr{C}^{\infty}(C),\{\cdot,\cdot\}_{\alpha})$ containing $S$, i.e. the smallest linear subspace of $\mathscr{C}^{\infty}(C)$ which contains $S$ and is closed under pointwise multiplication and the Jacobi bracket.
The following corollary shows that such Jacobi-commutative algebras of dissipated quantities admit a natural quantization in terms of invariant commutative $C^{\ast}$-subalgebras of quantum observables.
Corollary 4.19 (Quantization of Jacobi-commutative algebras by invariant quantum observables).
Let $(C,\alpha)$ be a contact manifold with Jacobi bracket $\{\cdot,\cdot\}_{\alpha}$, and let
$\mathcal{A}_{\mathrm{cl}}=\mathrm{Jacobi\text{-}alg}\langle I_{1},\dots,I_{n}\rangle\subset\mathscr{C}^{\infty}(C)$
be a Jacobi-commutative algebra of dissipated quantities, i.e.
$\{I_{i},I_{j}\}_{\alpha}=0\quad\text{for all }i,j,\qquad\{h,I_{i}\}_{\alpha}=-\lambda_{i}I_{i}$
for some contact Hamiltonian $h$ and real constants $\lambda_{i}$.
Assume that $Q_{\hbar}$ is a semiclassical quantization of $\mathcal{A}_{\mathrm{cl}}$ in the sense of Definition 4.14, and that there exists a family of Lindblad generators $\{\mathcal{L}_{\hbar}\}$ such that
  1. 1.
    the Heisenberg adjoint $\mathcal{L}_{\hbar}^{\dagger}$ satisfies the Egorov-type condition
    $\mathcal{L}_{\hbar}^{\dagger}(\widehat{f}_{\hbar})=Q_{\hbar}(\{f,h\}_{\alpha})+o(1)\qquad(\hbar\to 0)$
    for all $f\in\mathcal{A}_{\mathrm{cl}}$;
  2. 2.
    the Lindblad data are chosen inside the commutant of the $C^{\ast}$-algebra generated by the quantized integrals,
    $\widehat{H}_{\hbar},\widehat{I}_{i,\hbar}]=0,\qquad[L_{k,\hbar},\widehat{I}_{i,\hbar}]=[L_{k,\hbar}^{\dagger},\widehat{I}_{i,\hbar}]=0.$
Then the operators $\widehat{I}_{i,\hbar}\mathrel{\vbox{\hbox{\scriptsize.}\hbox{\scriptsize.}}}=Q_{\hbar}(I_{i})$ generate a unital, commutative, invariant $C^{\ast}$-subalgebra
$\mathcal{A}_{\hbar}=C^{\ast}(\widehat{I}_{1,\hbar},\dots,\widehat{I}_{n,\hbar})\subset\mathcal{B}(\mathcal{H}_{\hbar}),$
and every $A\in\mathcal{A}_{\hbar}$ is a quantum constant of motion:
$\mathcal{L}_{\hbar}^{\dagger}(A)=0.$
In this sense, invariant commutative $C^{\ast}$-subalgebras of observables provide the natural quantum counterparts of Jacobi-commutative algebras of dissipated quantities.
Proof.
Since the generators $I_{i}$ Jacobi-commute, the Dirac condition for the quantization map $Q_{\hbar}$ implies that their quantizations commute modulo $\mathcal{O}(\hbar)$:
$\widehat{I}_{i,\hbar},\widehat{I}_{j,\hbar}]=\mathcal{O}(\hbar).$
Assumption (2) ensures that the Hamiltonian and the Lindblad operators are chosen inside the commutant of the $C^{\ast}$-algebra generated by the $\widehat{I}_{i,\hbar}$. Therefore, the operators $\widehat{I}_{i,\hbar}$ generate a commutative $C^{\ast}$-subalgebra $\mathcal{A}_{\hbar}\subset\mathcal{B}(\mathcal{H}_{\hbar})$, which is unital by construction, since any $C^{\ast}$-subalgebra of $\mathcal{B}(\mathcal{H}_{\hbar})$ contains the identity operator.
Assumption (2) implies that the Hamiltonian and all Lindblad operators (and their adjoints) belong to the commutant $\mathcal{A}_{\hbar}^{\prime}$. By Lemma 4.7, it follows that
$\mathcal{L}_{\hbar}^{\dagger}(\widehat{I}_{i,\hbar})=0\quad\text{for all }i,$
and hence, by linearity and norm continuity of $\mathcal{L}_{\hbar}^{\dagger}$,
$\mathcal{L}_{\hbar}^{\dagger}(A)=0\quad\text{for all }A\in\mathcal{A}_{\hbar}.$
Finally, the Egorov-type condition (1) ensures that, on the level of principal symbols, the Heisenberg generator reduces to the contact Hamiltonian derivation associated with $h$ when restricted to $\mathcal{A}_{\mathrm{cl}}$. Therefore, $\mathcal{A}_{\hbar}$ can be interpreted as a quantization of the Jacobi-commutative algebra $\mathcal{A}_{\mathrm{cl}}$.∎
Remark 4.20.
In Corollary 4.19 it is stated that the $C^{*}$-algebra obtained by quantizing the algebra of functions in involution with respect to the Jacobi bracket is unitary by construction. This unitarity should be understood as a structural consequence of the quantization and $C^{*}$-completion procedure, rather than as a property that must be verified explicitly on the level of concrete generators.
More precisely, the quantization yields a *-algebra $\mathcal{A}_{0}$, equipped with an involution corresponding to the adjoint operation at the operator level. The associated $C^{*}$-algebra $\mathcal{A}$ is obtained as the completion of $\mathcal{A}_{0}$ with respect to the $C^{*}$-norm
$\|A\|=\sup_{\pi}\|\pi(A)\|,$
where $\pi$ ranges over all continuous *-representations on Hilbert spaces. By construction, the resulting $C^{\ast}$-algebra is realized as a unital $C^{\ast}$-subalgebra of $\mathcal{B}(\mathcal{H}_{\hbar})$, and therefore contains the identity operator $\mathbf{1}$ as well as a well-defined group of unitary elements.
This structural unitarity plays a crucial role in the subsequent Lindblad description. Indeed, the generators entering the Lindblad equation are required to be bounded operators on a Hilbert space, and the dissipative dynamics is formulated in terms of completely positive, trace-preserving semigroups acting on the underlying $C^{*}$-algebra of observables. The $C^{*}$-algebraic framework therefore provides the natural setting in which the Lindblad operators, the Hamiltonian part, and the associated quantum dynamical semigroup are consistently defined.
In this sense, the unitarity of the quantized algebra is not an auxiliary assumption, but a necessary ingredient ensuring that the Lindblad evolution constructed later is mathematically well-defined. The example of the Pauli matrices illustrates this mechanism in a finite-dimensional setting, but the same reasoning applies in full generality.
This remark should be read in conjunction with the discussion preceding Corollary 4.19, where the distinct roles of the unit in the classical Jacobi algebra and in its quantum counterpart are clarified.
4.4 Bi-Lindblad structures
The purpose of this subsection is to introduce a quantum analog of bi-Hamiltonian pencils: instead of a Poisson pencil $\pi_{\lambda}=\pi_{1}-\lambda\pi_{0}$ (closed under affine combination), we work with a convex pencil of GKSL generators $\mathcal{L}^{(\lambda)}_{\hbar}=(1-\lambda)\mathcal{L}^{(0)}_{\hbar}+\lambda\mathcal{L}^{(1)}_{\hbar}$, and we require that the same family of quantized integrals $\widehat{I}_{j,\hbar}$ lies in the kernel of both Heisenberg adjoints. This is the structural mechanism by which a single classical bi-Hamiltonian hierarchy produces two (or more) compatible open quantum evolutions sharing a common invariant commutative sector.
We now formulate a quantum counterpart of bi-Hamiltonian structures, in which the role of compatible Poisson tensors $(\pi_{0},\pi_{1})$ is played by a pair of Lindblad generators whose convex combinations remain Lindbladian and whose common quantum integrals quantize the classical dissipated quantities $I_{j}$.
Let $(C,\alpha,h;I_{0},\dots,I_{n})$ and $(\mathcal{H}_{\hbar},Q_{\hbar})$ be as in Definition 4.14. Assume in addition that the contact system arises from a bi-Hamiltonian Poisson–Lie system $(G,\pi_{0},\pi_{1};H_{0},H_{1})$ as in Section 3, so that the classical vector field $X$ is Hamiltonian both for $\pi_{0}$ and $\pi_{1}$, and the hierarchy of integrals $\{I_{j}\}$ ultimately comes from the common bi-Hamiltonian invariants on $G$.
At the quantum level, the two compatible Poisson–Lie Hamiltonians $H_{0},H_{1}$ do not give rise directly to Lindblad generators, but rather induce, via the homogeneous lift and restriction procedure of Section 3, a pair of contact Hamiltonians on $C$. We denote by $H_{0}^{\mathrm{eff}},\;H_{1}^{\mathrm{eff}}\in\mathcal{A}_{\mathrm{cl}}$ the corresponding effective contact Hamiltonians governing the classical dynamics on $C$ associated with the two Poisson structures.
Definition 4.21 (Bi-Lindblad structure).
A bi-Lindblad structure associated with the contact integrable system $(C,\alpha,h;I_{0},\dots,I_{n})$ and its semiclassical quantization $(\mathcal{H}_{\hbar},Q_{\hbar})$ consists of a pair of families of Lindblad generators $\{\mathcal{L}^{(0)}_{\hbar}\}_{\hbar\in(0,\hbar_{0}]}$ and $\{\mathcal{L}^{(1)}_{\hbar}\}_{\hbar\in(0,\hbar_{0}]}$ on $\mathcal{B}(\mathcal{H}_{\hbar})$ such that
  1. (BL1)
    (Convex compatibility) For every $\lambda\in[0,1]$ and every $\hbar\in(0,\hbar_{0}]$, the convex combination
    $\mathcal{L}^{(\lambda)}_{\hbar}\mathrel{\vbox{\hbox{\scriptsize.}\hbox{\scriptsize.}}}=(1-\lambda)\,\mathcal{L}^{(0)}_{\hbar}+\lambda\,\mathcal{L}^{(1)}_{\hbar}$
    is again a GKSL generator on $\mathcal{B}(\mathcal{H}_{\hbar})$. In other words, the convex cone generated by $\mathcal{L}^{(0)}_{\hbar}$ and $\mathcal{L}^{(1)}_{\hbar}$ is contained in the space of Lindblad generators.
  2. (BL2)
    (Common quantum integrals) For each $j=0,\dots,n$ and each $\hbar\in(0,\hbar_{0}]$ one has
    $\mathcal{L}^{(0)\,\dagger}_{\hbar}\big(\widehat{I}_{j,\hbar}\big)=\mathcal{L}^{(1)\,\dagger}_{\hbar}\big(\widehat{I}_{j,\hbar}\big)=0.$
    Equivalently, the quantized dissipated quantities $\widehat{I}_{j,\hbar}$ lie in the joint kernel
    $\ker\mathcal{L}^{(0)\,\dagger}_{\hbar}\cap\ker\mathcal{L}^{(1)\,\dagger}_{\hbar}.$
  3. (BL3)
    (Semiclassical bi-Hamiltonian limit) For every $f\in\mathcal{A}_{\mathrm{cl}}$ and $a\in\{0,1\}$,
    $\lim_{\hbar\to 0}\left\|\mathcal{L}^{(a)\,\dagger}_{\hbar}\big(\widehat{f}_{\hbar}\big)-Q_{\hbar}\big(\{f,H_{a}^{\mathrm{eff}}\}_{\alpha}\big)\right\|=0,$
    where $H_{0}^{\mathrm{eff}},H_{1}^{\mathrm{eff}}\in\mathcal{A}_{\mathrm{cl}}$ are effective contact Hamiltonians whose classical flows on $C$ are obtained from the underlying bi-Hamiltonian Poisson–Lie data $(G,\pi_{0},\pi_{1};H_{0},H_{1})$ via the homogeneous lift and restriction of Section 3.
Heuristically, (BL1) mirrors the classical requirement that any linear combination $\pi_{\lambda}=\pi_{1}-\lambda\pi_{0}$ of compatible Poisson bivectors is still a Poisson tensor: here convex combinations of the two generators remain physically admissible Lindblad generators. Condition (BL2) captures the fact that the same family of classical integrals $I_{j}$ arises from the bi-Hamiltonian hierarchy and should therefore be conserved by both quantum evolutions. Finally, (BL3) encodes the requirement that, in the semiclassical limit, each branch $\mathcal{L}^{(a)}_{\hbar}$ recovers the contact dynamics associated with a different classical Hamiltonian $H_{a}$, in analogy with the way a bi-Hamiltonian vector field is Hamiltonian for both $\pi_{0}$ and $\pi_{1}$.
Remark 4.22.
Condition (BL1) is natural from the open-systems viewpoint: the set of GKSL generators forms a convex cone (sums and positive scalar multiples remain of GKSL type). Thus, in many concrete constructions, (BL1) is essentially a bookkeeping requirement ensuring that two admissible dissipative mechanisms can be blended continuously without leaving the Lindblad class. $\diamond$
The following simple observation shows that any bi-Lindblad structure in the sense of Definition 4.21 automatically yields a contact-compatible Lindblad generator for each $\lambda\in[0,1]$.
Proposition 4.23.
Let $\{\mathcal{L}^{(0)}_{\hbar}\}$ and $\{\mathcal{L}^{(1)}_{\hbar}\}$ define a bi-Lindblad structure in the sense of Definition 4.21, and fix $\lambda\in[0,1]$. Then, for each $\hbar\in(0,\hbar_{0}]$, the generator
$\mathcal{L}^{(\lambda)}_{\hbar}=(1-\lambda)\,\mathcal{L}^{(0)}_{\hbar}+\lambda\,\mathcal{L}^{(1)}_{\hbar}$
is contact-compatible with $(C,\alpha,h;I_{0},\dots,I_{n})$ in the sense of Definition 4.16, with respect to the effective Hamiltonian
$H_{\lambda}^{\mathrm{eff}}\mathrel{\vbox{\hbox{\scriptsize.}\hbox{\scriptsize.}}}=(1-\lambda)\,H_{0}^{\mathrm{eff}}+\lambda\,H_{1}^{\mathrm{eff}}\in\mathcal{A}_{\mathrm{cl}}.$
Proof.
By (BL1), $\mathcal{L}^{(\lambda)}_{\hbar}$ is a GKSL generator for each $\lambda\in[0,1]$ and $\hbar\in(0,\hbar_{0}]$.
Verification of (CC1). Since the Heisenberg adjoint is affine,
$\mathcal{L}^{(\lambda)\,\dagger}_{\hbar}=(1-\lambda)\,\mathcal{L}^{(0)\,\dagger}_{\hbar}+\lambda\,\mathcal{L}^{(1)\,\dagger}_{\hbar}$,
and by (BL2) we have $\mathcal{L}^{(a)\,\dagger}_{\hbar}(\widehat{I}_{j,\hbar})=0$ for $a\in\{0,1\}$, it follows that
$\mathcal{L}^{(\lambda)\,\dagger}_{\hbar}(\widehat{I}_{j,\hbar})=0\qquad\text{for all }j=0,\dots,n.$
Thus (CC1) holds for $\mathcal{L}^{(\lambda)}_{\hbar}$.
Verification of (CC2). Fix $f\in\mathcal{A}_{\mathrm{cl}}$. Using bilinearity of the Jacobi bracket and linearity of $Q_{\hbar}$, we have
$Q_{\hbar}\!\big(\{f,H_{\lambda}^{\mathrm{eff}}\}_{\alpha}\big)=(1-\lambda)\,Q_{\hbar}\!\big(\{f,H_{0}^{\mathrm{eff}}\}_{\alpha}\big)+\lambda\,Q_{\hbar}\!\big(\{f,H_{1}^{\mathrm{eff}}\}_{\alpha}\big).$
Hence, by the triangle inequality,
$\Big\|\mathcal{L}^{(\lambda)\,\dagger}_{\hbar}(\widehat{f}_{\hbar})-Q_{\hbar}\!\big(\{f,H_{\lambda}^{\mathrm{eff}}\}_{\alpha}\big)\Big\|$
$\leq(1-\lambda)\,\Big\|\mathcal{L}^{(0)\,\dagger}_{\hbar}(\widehat{f}_{\hbar})-Q_{\hbar}\!\big(\{f,H_{0}^{\mathrm{eff}}\}_{\alpha}\big)\Big\|+\lambda\,\Big\|\mathcal{L}^{(1)\,\dagger}_{\hbar}(\widehat{f}_{\hbar})-Q_{\hbar}\!\big(\{f,H_{1}^{\mathrm{eff}}\}_{\alpha}\big)\Big\|.$
Each term tends to 0 as $\hbar\to 0$ by (BL3), so the left-hand side also tends to 0. This is precisely (CC2) for $\mathcal{L}^{(\lambda)}_{\hbar}$ with effective Hamiltonian $H_{\lambda}^{\mathrm{eff}}$. ∎
Example: a bi-Lindblad pencil interpolating between unitary and dephasing dynamics.
We illustrate Definition 4.21 on the two-level system $\mathcal{H}=\mathbb{C}^{2}$, where the invariant commutative sector is generated by $\sigma_{z}$. Fix parameters $\omega_{0},\omega_{1}\in\mathbb{R}$ and $\gamma>0$, and define two GKSL generators on $\mathcal{B}(\mathbb{C}^{2})$ (not forming a semiclassical family in $\hbar$, but sufficient to illustrate the algebraic content of Definition 4.21):
$\mathcal{L}^{(0)}(\rho)\mathrel{\vbox{\hbox{\scriptsize.}\hbox{\scriptsize.}}}=-\frac{i}{\hbar}\big[H_{0},\rho\big],\qquad H_{0}\mathrel{\vbox{\hbox{\scriptsize.}\hbox{\scriptsize.}}}=\frac{\hbar\omega_{0}}{2}\sigma_{z},$
and
[…]
Note that $\mathcal{L}^{(0)}$ describes a purely unitary evolution, while $\mathcal{L}^{(1)}$ corresponds to a unitary evolution plus pure dephasing in the $\sigma_{z}$ basis. For every $\lambda\in[0,1]$, let
$\mathcal{L}^{(\lambda)}\mathrel{\vbox{\hbox{\scriptsize.}\hbox{\scriptsize.}}}=(1-\lambda)\mathcal{L}^{(0)}+\lambda\mathcal{L}^{(1)}$
is again of GKSL form, with Hamiltonian $H_{\lambda}=\frac{\hbar}{2}\big((1-\lambda)\omega_{0}+\lambda\omega_{1}\big)\sigma_{z}$ and a single Lindblad operator $\sqrt{\lambda}\,L$; hence (BL1) holds.
In the Heisenberg picture, the adjoints satisfy
$\mathcal{L}^{(a)\,\dagger}(\sigma_{z})=0,\qquad a\in\{0,1\},$
since $H_{a},\sigma_{z}]=0$ and $L,\sigma_{z}]=0$. Therefore $\sigma_{z}$ (and trivially $\mathbb{I}$) lies in the joint kernel $\ker\mathcal{L}^{(0)\,\dagger}\cap\ker\mathcal{L}^{(1)\,\dagger}$, i.e. (BL2) holds for the commutative algebra $C^{\ast}(\sigma_{z})$.
This example captures the basic geometric idea behind bi-Lindblad structures: two physically distinct open-system mechanisms (here “purely Hamiltonian” versus “Hamiltonian + dephasing”) share a common invariant commutative sector, and convex interpolation preserves complete positivity and the same quantum constants of motion. In our later semiclassical constructions, the role of $\sigma_{z}$ will be played by the quantized dissipated quantities $\widehat{I}_{j,\hbar}$, and the two branches $a=0,1$ will correspond to two different effective classical Hamiltonians in the underlying bi-Hamiltonian hierarchy.
Thus a bi-Lindblad structure provides, for each $\lambda\in[0,1]$, a specific choice of contact-compatible Lindblad generator whose semiclassical limit reflects a different point on the classical bi-Hamiltonian pencil. In the deformed Euler top example of Section 5, we shall see explicitly how the quantum generators $\mathcal{L}^{(0)}_{\hbar}$ and $\mathcal{L}^{(1)}_{\hbar}$ can be chosen so that their common quantum integrals $\widehat{I}_{j,\hbar}$ quantize the contact dissipated quantities $I_{j}$ obtained in Section 3, thereby realizing a concrete bi-Lindblad structure in the above sense.
Weyl quantization.
Before stating the Egorov-type criterion, we briefly recall what we mean by semiclassical Weyl quantization in the standard pseudodifferential setting. Assume that the exact symplectic manifold $(M,\omega)$ is globally symplectomorphic to $\mathbb{R}^{2d}$ (with $d=n+1$) with canonical coordinates $(x,\xi)\in\mathbb{R}^{\mathrm{d}}\times\mathbb{R}^{d}$ and symplectic form ${\omega=\sum_{j=1}^{d}\mathrm{d}\xi_{j}\wedge dx_{j}}$.
In the following, $S^{0}(M)$ will denote the standard class of semiclassical symbols of order zero on $M\simeq\mathbb{R}^{2d}$, that is, the space of smooth functions $a(x,\xi;\hbar)$ satisfying
$|\partial_{x}^{\alpha}\partial_{\xi}^{\beta}a(x,\xi;\hbar)|\leq C_{\alpha\beta}\qquad\text{for all multi-indices }\alpha,\beta,$
uniformly in $(x,\xi)\in\mathbb{R}^{2d}$ and $\hbar\in(0,\hbar_{0}]$. In addition, $\mathcal{S}(\mathbb{R}^{d})$ will denote the Schwartz space of rapidly decreasing smooth functions on $\mathbb{R}^{d}$, that is, the space of all functions $\psi\in\mathscr{C}^{\infty}(\mathbb{R}^{d})$ such that
$\sup_{x\in\mathbb{R}^{d}}|x^{\alpha}\partial_{x}^{\beta}\psi(x)|<\infty\qquad\text{for all multi-indices }\alpha,\beta.$
Given a (bounded) semiclassical symbol $a\in S^{0}(M)$, the Weyl quantization $\operatorname{Op}_{\hbar}^{\mathrm{w}}(a)$ is the pseudodifferential operator acting on $L^{2}(\mathbb{R}^{d})$ by
$\big(\operatorname{Op}_{\hbar}^{\mathrm{w}}(a)\psi\big)(x)\mathrel{\vbox{\hbox{\scriptsize.}\hbox{\scriptsize.}}}=\frac{1}{(2\pi\hbar)^{d}}\int_{\mathbb{R}^{d}}\int_{\mathbb{R}^{d}}a\!\left(\frac{x+y}{2},\xi\right)\,e^{\frac{i}{\hbar}(x-y)\cdot\xi}\,\psi(y)\,\mathrm{d}y\,\mathrm{d}\xi,\qquad\psi\in\mathcal{S}(\mathbb{R}^{d}),$
and extends by density to a bounded operator on $L^{2}(\mathbb{R}^{d})$ when $a\in S^{0}$.
The defining feature of Weyl quantization is its symmetric placement of the phase-space variables, which ensures (i) good reality properties (real-valued symbols yield essentially self-adjoint operators) and (ii) a clean semiclassical symbolic calculus.
In particular, the semiclassical Moyal product $\sharp$ satisfies
$\operatorname{Op}_{\hbar}^{\mathrm{w}}(a)\,\operatorname{Op}_{\hbar}^{\mathrm{w}}(b)=\operatorname{Op}_{\hbar}^{\mathrm{w}}(a\sharp b),\qquad a\sharp b=ab+\frac{\hbar}{2i}\{a,b\}_{\omega}+\mathcal{O}(\hbar^{2}),$
and consequently the commutator admits the expansion
$\frac{1}{i\hbar}\big[\operatorname{Op}_{\hbar}^{\mathrm{w}}(a),\operatorname{Op}_{\hbar}^{\mathrm{w}}(b)\big]=\operatorname{Op}_{\hbar}^{\mathrm{w}}\big(\{a,b\}_{\omega}\big)+\mathcal{O}(\hbar^{2}),$
in operator norm for $a,b\in S^{0}(M)$ under standard symbol assumptions. This is the pseudodifferential backbone behind the Dirac condition and the Egorov-type limit used below. We refer to [Zworski2012, DimassiSjoestrand1999] for comprehensive accounts of semiclassical Weyl quantization, symbol classes $S^{m}$, the Moyal product, and the Egorov theorem. See also [Folland1989, Martinez2002] for the Weyl calculus and its microlocal properties. From the perspective of deformation quantization, the Moyal product and the Dirac condition are discussed for instance in [Gutt2000].
In our contact setting, the quantization map $Q_{\hbar}$ on the Jacobi subalgebra $\mathcal{A}_{\mathrm{cl}}\subset\mathscr{C}^{\infty}(C)$ is obtained by first taking the homogeneous lift $f\mapsto f^{\Sigma}$ (so $f^{\Sigma}|_{C}=f$) and then applying Weyl quantization:
$Q_{\hbar}(f)=\operatorname{Op}_{\hbar}^{\mathrm{w}}(f^{\Sigma}),\qquad f\in\mathcal{A}_{\mathrm{cl}}.$
This realizes the contact Dirac condition in Definition 4.14 as a direct consequence of the Weyl commutator expansion and the Poisson–Jacobi correspondence of Proposition 2.8.
Lindblad operators of Weyl type.
In Theorem 4.24 we assume that the Lindblad operators $L_{k,\hbar}$ are of Weyl type. This means that each dissipative operator admits a semiclassical pseudodifferential representation whose leading symbol is a well-defined classical observable on the exact symplectic manifold $(M,\mathrm{d}\theta)$. More precisely, saying that
$L_{k,\hbar}=\operatorname{Op}_{\hbar}^{\mathrm{w}}(\ell_{k})+\mathcal{O}(\hbar)\quad\text{in operator norm}$
means that:
  • $\ell_{k}\in S^{0}(M)$ is a bounded classical symbol (the principal symbol of $L_{k,\hbar}$);
  • the remainder $\mathcal{O}(\hbar)$ is controlled in operator norm, in the sense that
    $\big\|L_{k,\hbar}-\operatorname{Op}_{\hbar}^{\mathrm{w}}(\ell_{k})\big\|\leq C\,\hbar\quad\text{for all }k\leq N(\hbar),\ \hbar\in(0,\hbar_{0}]$
    with $C$ independent of $k$ and $\hbar$, ensuring a uniform semiclassical expansion compatible with the pseudodifferential calculus;
  • the leading-order action of the dissipator can be analyzed directly at the level of classical phase space via Poisson brackets.
This assumption is the natural open-system analog of taking the Hamiltonian $\widehat{H}_{\hbar}$ to be a Weyl quantization of a classical Hamiltonian function. It ensures that dissipation is not introduced in an ad hoc operator-theoretic way, but rather descends from classical data on $M$ and admits a meaningful semiclassical interpretation.
In particular, when the symbols $\ell_{k}$ are chosen as functions of the classical integrals $(F_{0},\dots,F_{n})$, the corresponding Lindblad operators act (approximately) diagonally, up to $\mathcal{O}(\hbar)$ corrections, on the joint eigenspaces of the quantized integrals $\widehat{I}_{j,\hbar}=\operatorname{Op}_{\hbar}^{\mathrm{w}}(I_{j}^{\Sigma})$. This is precisely the mechanism behind pure dephasing: the dissipator suppresses coherences between different eigenvalues of the integrals, while leaving their expectation values invariant.
Thus, the Weyl-type assumption on the Lindblad operators is the technical bridge that allows one to:
  1. 1.
    control the semiclassical limit of the dissipative dynamics;
  2. 2.
    prove the Egorov-type statement (CC2);
  3. 3.
    interpret decoherence geometrically in terms of the classical integrals and their associated invariant foliations.
Theorem 4.24 (Pseudodifferential Egorov criterion for (CC2)).
Assume that the contact integrable system $(C,\alpha,h;I_{0},\dots,I_{n})$ arises from a homogeneous Hamiltonian system $(M,\theta,H)$ as in Proposition 3.7, with $C$ realized as a Liouville-transverse hypersurface in the exact symplectic manifold $(M,\theta)$.
Suppose that:
  1. 1.
    $M$ is globally symplectomorphic to $\mathbb{R}^{2(n+1)}$ (or to a cotangent bundle ${\mathsf{T}}^{\ast}Q$) and the quantization maps $Q_{\hbar}$ in Definition 4.14 are obtained by restricting a semiclassical Weyl quantization
    $\operatorname{Op}_{\hbar}^{\mathrm{w}}\colon S^{0}(M)\longrightarrow\mathcal{B}(\mathcal{H}_{\hbar})$
    of order-zero symbols on $M$ to the subalgebra $\mathcal{A}_{\mathrm{cl}}\subset\mathscr{C}^{\infty}(C)$ via the homogeneous lift $f\mapsto f^{\Sigma}$ of Proposition 2.8.
  2. 2.
    The Hamiltonian operators $\widehat{H}_{\hbar}$ are of the form
    $\widehat{H}_{\hbar}=\operatorname{Op}_{\hbar}^{\mathrm{w}}(H)+R_{\hbar}\,,\quad\|R_{\hbar}\|=\mathcal{O}(\hbar^{2})\,$,
    where $H$ is the $1$-homogeneous classical Hamiltonian on $M$ whose restriction to $C$ gives the contact Hamiltonian $h$.
  3. 3.
    The Lindblad operators are of Weyl type,
    $L_{k,\hbar}=\operatorname{Op}_{\hbar}^{\mathrm{w}}(\ell_{k})+\mathcal{O}(\hbar),\qquad k=1,\dots,N(\hbar),$
    with symbols $\ell_{k}\in S^{0}(M)$ such that for each $k$ and every $f\in\mathcal{A}_{\mathrm{cl}}$ one has
    $\{\ell_{k},f^{\Sigma}\}_{\omega}=0\qquad\text{and}\qquad\{|\ell_{k}|^{2},f^{\Sigma}\}_{\omega}=0\quad\text{on }M,$
    and the number of Lindblad operators satisfies $N(\hbar)=\mathcal{O}(1)$ as $\hbar\to 0$.
Then the family of Lindblad generators $\{\mathcal{L}_{\hbar}\}$ satisfies (CC2) with respect to the contact Hamiltonian $h$, i.e. for every $f\in\mathcal{A}_{\mathrm{cl}}$,
$\lim_{\hbar\to 0}\Big\|\mathcal{L}_{\hbar}^{\dagger}\big(\widehat{f}_{\hbar}\big)-Q_{\hbar}\big(\{f,h\}_{\alpha}\big)\Big\|=0.$
Moreover, if in addition the Lindblad data are chosen so that
[…]
then (CC1) holds as well, i.e. $\mathcal{L}_{\hbar}^{\dagger}(\widehat{I}_{j,\hbar})=0$ for $j=0,\dots,n$.
Proof.
We first prove (CC2) (Steps 1–4), and then discuss (CC1) (Step 5).
Step 1: Dirac condition for $Q_{\hbar}$ and basic properties. By hypothesis, $(M,\omega)$ is symplectomorphic to $\mathbb{R}^{2(n+1)}$ (or a cotangent bundle ${\mathsf{T}}^{\ast}Q$), and $Q_{\hbar}$ is obtained by composing the homogeneous lift $f\mapsto f^{\Sigma}$ of Proposition 2.8 with a semiclassical Weyl quantization $\operatorname{Op}_{\hbar}^{\mathrm{w}}\colon S^{0}(M)\longrightarrow\mathcal{B}(\mathcal{H}_{\hbar})$. Thus, for $f\in\mathcal{A}_{\mathrm{cl}}$,
$\widehat{f}_{\hbar}\mathrel{\vbox{\hbox{\scriptsize.}\hbox{\scriptsize.}}}=Q_{\hbar}(f)=\operatorname{Op}_{\hbar}^{\mathrm{w}}(f^{\Sigma}),$
and the standard semiclassical symbolic calculus gives, for $a,b\in S^{0}(M)$,
$\frac{1}{i\hbar}\big[\operatorname{Op}_{\hbar}^{\mathrm{w}}(a),\operatorname{Op}_{\hbar}^{\mathrm{w}}(b)\big]=\operatorname{Op}_{\hbar}^{\mathrm{w}}\big(\{a,b\}_{\omega}\big)+\mathcal{O}(\hbar^{2})$
in operator norm. Applying this with $a=f^{\Sigma}$, $b=g^{\Sigma}$ and using the Poisson–Jacobi correspondence $\{f^{\Sigma},g^{\Sigma}\}_{\omega}=(\{f,g\}_{\alpha})^{\Sigma},$ we obtain the Dirac condition of Definition 4.14 (with an explicit $\mathcal{O}(\hbar^{2})$ remainder).
Step 2: Hamiltonian part and (CC2). Fix $f\in\mathcal{A}_{\mathrm{cl}}$ and write $\widehat{f}_{\hbar}=\operatorname{Op}_{\hbar}^{\mathrm{w}}(f^{\Sigma})$. By hypothesis (2),
$\frac{i}{\hbar}\big[\widehat{H}_{\hbar},\widehat{f}_{\hbar}\big$] $=\frac{i}{\hbar}\big[\operatorname{Op}_{\hbar}^{\mathrm{w}}(H),\operatorname{Op}_{\hbar}^{\mathrm{w}}(f^{\Sigma})\big]+\frac{i}{\hbar}\big[R_{\hbar},\operatorname{Op}_{\hbar}^{\mathrm{w}}(f^{\Sigma})\big].$
The second term is $\mathcal{O}(\hbar)$ in operator norm since $\|\operatorname{Op}_{\hbar}^{\mathrm{w}}(f^{\Sigma})\|$ is uniformly bounded for $f^{\Sigma}\in S^{0}$ and $\|R_{\hbar}\|=\mathcal{O}(\hbar^{2})$. For the first term, the commutator expansion yields
$\frac{i}{\hbar}\big[\operatorname{Op}_{\hbar}^{\mathrm{w}}(H),\operatorname{Op}_{\hbar}^{\mathrm{w}}(f^{\Sigma})\big]=\operatorname{Op}_{\hbar}^{\mathrm{w}}\big(\{H,f^{\Sigma}\}_{\omega}\big)+\mathcal{O}(\hbar^{2}).$
Using $\{H,f^{\Sigma}\}_{\omega}=(\{f,h\}_{\alpha})^{\Sigma}$, we obtain
$\frac{i}{\hbar}\big[\widehat{H}_{\hbar},\widehat{f}_{\hbar}\big]=Q_{\hbar}\big(\{f,h\}_{\alpha}\big)+\mathcal{O}(\hbar)$
in operator norm.
Step 3: Estimate of the dissipative part. Fix $f\in\mathcal{A}_{\mathrm{cl}}$ and set $\widehat{f}_{\hbar}=\operatorname{Op}_{\hbar}^{\mathrm{w}}(f^{\Sigma})$. Let $\mathcal{D}_{\hbar}$ denote the dissipator in the Heisenberg picture,
$\mathcal{D}_{\hbar}(A)\mathrel{\vbox{\hbox{\scriptsize.}\hbox{\scriptsize.}}}=\sum_{k=1}^{N(\hbar)}\Big(L_{k,\hbar}^{\dagger}AL_{k,\hbar}-\tfrac{1}{2}\{L_{k,\hbar}^{\dagger}L_{k,\hbar},A\}\Big).$
For each $k$, let
$\mathcal{D}_{k,\hbar}(A)\mathrel{\vbox{\hbox{\scriptsize.}\hbox{\scriptsize.}}}=L_{k,\hbar}^{\dagger}AL_{k,\hbar}-\tfrac{1}{2}\{L_{k,\hbar}^{\dagger}L_{k,\hbar},A\},$
so that ${\mathcal{D}_{\hbar}=\sum_{k=1}^{N(\hbar)}\mathcal{D}_{k,\hbar}}$ and hence
[…]
$d_{k,\hbar}=d_{k}^{(0)}+\hbar d_{k}^{(1)}+\mathcal{O}(\hbar^{2}).$
The order-$\hbar^{0}$ term cancels identically,
$d_{k}^{(0)}=\overline{\ell_{k}}f^{\Sigma}\ell_{k}-\tfrac{1}{2}\big(|\ell_{k}|^{2}f^{\Sigma}+f^{\Sigma}|\ell_{k}|^{2}\big)=0.$
Moreover, $d_{k}^{(1)}$ is a linear combination of Poisson brackets of the form $\{\ell_{k},f^{\Sigma}\}_{\omega}$ and $\{|\ell_{k}|^{2},f^{\Sigma}\}_{\omega}$. Assumption (3) therefore implies $d_{k}^{(1)}\equiv 0$ on $M$, and hence
$\mathcal{D}_{k,\hbar}(\widehat{f}_{\hbar})=\mathcal{O}(\hbar)\quad\text{in operator norm.}$
Summing over $k$ we obtain
$\mathcal{D}_{\hbar}(\widehat{f}_{\hbar})=\sum_{k=1}^{N(\hbar)}\mathcal{D}_{k,\hbar}(\widehat{f}_{\hbar})=\mathcal{O}\!\big(N(\hbar)\hbar\big),$
and since $N(\hbar)=\mathcal{O}(1)$ we conclude that
$\mathcal{D}_{\hbar}(\widehat{f}_{\hbar})=\mathcal{O}(\hbar)\quad\text{as }\hbar\to 0.$
Consequently, for each fixed $f\in\mathcal{A}_{\mathrm{cl}}$,
$\|\mathcal{D}_{\hbar}(\widehat{f}_{\hbar})\|=\mathcal{O}(\hbar)\qquad\text{as }\hbar\to 0.$
This estimate is sufficient for establishing the semiclassical limit required in (CC2).
Step 4: Conclusion of (CC2). Combining Steps 2 and 3 we obtain
$\mathcal{L}_{\hbar}^{\dagger}(\widehat{f}_{\hbar})=\frac{i}{\hbar}\big[\widehat{H}_{\hbar},\widehat{f}_{\hbar}\big]+\mathcal{D}_{\hbar}(\widehat{f}_{\hbar})=Q_{\hbar}\big(\{f,h\}_{\alpha}\big)+\mathcal{O}(\hbar),$
hence
$\lim_{\hbar\to 0}\Big\|\mathcal{L}_{\hbar}^{\dagger}(\widehat{f}_{\hbar})-Q_{\hbar}\big(\{f,h\}_{\alpha}\big)\Big\|=0,$
which is (CC2).
Step 5: Exact quantum constants of motion. Finally, (CC1) follows from the commutation relations (4.3) and Lemma 4.7.
Example 4.25.
We illustrate Theorem 4.24 on the linear contact model of Example 2.11, with
$(C,\alpha,h)=(\mathbb{R}^{3}(q,p,z),\,\mathrm{d}z-p\,\mathrm{d}q,\,h=z-p),\qquad I_{1}\mathrel{\vbox{\hbox{\scriptsize.}\hbox{\scriptsize.}}}=e^{-z}.$
This system arises by restriction from the exact symplectic homogeneous system
$M=C\times\mathbb{R}_{>0}\cong\mathbb{R}^{4}(q,p,z,r),\qquad\theta=r(\mathrm{d}z-p\,\mathrm{d}q),\qquad\omega\mathrel{\vbox{\hbox{\scriptsize.}\hbox{\scriptsize.}}}=-\mathrm{d}\theta,\qquad H^{\Sigma}\mathrel{\vbox{\hbox{\scriptsize.}\hbox{\scriptsize.}}}=-r\,h=-r(z-p),$
with Liouville-transverse hypersurface $C=\{r=1\}$. The Jacobi subalgebra is
$\mathcal{A}_{\mathrm{cl}}=\mathrm{Jacobi\text{-}alg}\langle I_{0}=h,\,I_{1}=e^{-z}\rangle\subset\mathscr{C}^{\infty}(C).$
Assume that $M$ is quantized by semiclassical Weyl quantization on $\mathbb{R}^{4}$, so that
$Q_{\hbar}(f)=\operatorname{Op}_{\hbar}^{\mathrm{w}}(f^{\Sigma}),\qquad f\in\mathcal{A}_{\mathrm{cl}}.$
Choose the quantum Hamiltonian
$\widehat{H}_{\hbar}=\operatorname{Op}_{\hbar}^{\mathrm{w}}(H^{\Sigma})+\mathcal{O}(\hbar^{2}).$
For the dissipative part, take a single Lindblad operator
$L_{\hbar}=\operatorname{Op}_{\hbar}^{\mathrm{w}}(\ell)+\mathcal{O}(\hbar),\qquad\ell=\sqrt{\gamma}\,\varphi\!\big(H^{\Sigma},I_{1}^{\Sigma}\big),\qquad\gamma>0,$
where $\varphi$ is a smooth real-valued function and $I_{1}^{\Sigma}$ denotes the $1$-homogeneous lift of $I_{1}$.
Since $\ell$ depends only on the commuting homogeneous integrals $(H^{\Sigma},I_{1}^{\Sigma})$, one has
$\{\ell,f^{\Sigma}\}_{\omega}=0\qquad\text{and}\qquad\{|\ell|^{2},f^{\Sigma}\}_{\omega}=0\quad\text{on }M\qquad\forall f\in\mathcal{A}_{\mathrm{cl}}.$
Moreover, here $N(\hbar)=1=\mathcal{O}(1)$.
Therefore, all the hypotheses of Theorem 4.24 are satisfied, and the Lindblad generator $\mathcal{L}_{\hbar}$ fulfills (CC2).
If, in addition, the operators $\widehat{I}_{0,\hbar}$ and $\widehat{I}_{1,\hbar}$ are chosen so that $L_{\hbar}$ is implemented via a bounded functional calculus of these commuting observables, then the commutation relations (4.3) hold and (CC1) is satisfied exactly.
Heuristically, this provides a semiclassical model of pure dephasing: the dissipator is diagonal with respect to the joint spectral decomposition of the commuting quantum integrals, suppressing off-diagonal coherences while preserving their expectation values.
Remark 4.26.
We now give a concrete physical interpretation of the Egorov-type criterion (Theorem 4.24) in the language of decoherence, in a way tailored to the Weyl-type and commutant-based framework adopted above.
Fix $\hbar$ and consider the commutative $C^{\ast}$-algebra
$\mathcal{A}_{\hbar}\mathrel{\vbox{\hbox{\scriptsize.}\hbox{\scriptsize.}}}=C^{\ast}(\widehat{I}_{0,\hbar},\dots,\widehat{I}_{n,\hbar})\subset\mathcal{B}(\mathcal{H}_{\hbar}),$
generated by the quantized dissipated quantities. By construction, $\mathcal{A}_{\hbar}$ is unital and consists of bounded operators.
In all dephasing-type examples relevant for us, the Lindblad operators are chosen to be of Weyl type and, more specifically, as bounded functions of the commuting quantum integrals, i.e.
$L_{k,\hbar}=\sqrt{\gamma_{k}}\,\varphi_{k}(\widehat{I}_{0,\hbar},\dots,\widehat{I}_{n,\hbar}),\qquad\gamma_{k}\geq 0,$
with a number of channels satisfying $N(\hbar)=\mathcal{O}(1)$ as $\hbar\to 0$. At the level of principal symbols, this corresponds to choosing $\ell_{k}\in S^{0}(M)$ depending only on the homogeneous lifts of the classical integrals, so that
$\{\ell_{k},f^{\Sigma}\}_{\omega}=\{|\ell_{k}|^{2},f^{\Sigma}\}_{\omega}=0\quad\text{on }M$
for all $f\in\mathcal{A}_{\mathrm{cl}}$.
As a consequence, the Lindblad operators belong to $\mathcal{A}_{\hbar}$ and hence
$L_{k,\hbar},A]=[L_{k,\hbar}^{\dagger},A]=0\quad\forall\,A\in\mathcal{A}_{\hbar}.$
If, in addition, the Hamiltonian part is chosen in the commutant of $\mathcal{A}_{\hbar}$, i.e. $\widehat{H}_{\hbar},\widehat{I}_{j,\hbar}]=0$ for all $j$, then the exact commutation relations (4.3) hold. By the argument of Step 5 in the proof of Theorem 4.24, this implies that every observable in $\mathcal{A}_{\hbar}$ is an exact quantum constant of motion:
$\mathcal{L}_{\hbar}^{\dagger}(A)=0\qquad\forall\,A\in\mathcal{A}_{\hbar}.$
Thus $\mathcal{A}_{\hbar}$ defines an invariant commutative sector of the Heisenberg evolution.
This algebraic invariance has a direct decoherence interpretation. Since the operators $\widehat{I}_{j,\hbar}$ commute, they can be simultaneously diagonalized, yielding a decomposition of the Hilbert space into joint eigenspaces,
$\mathcal{H}_{\hbar}=\bigoplus_{\nu}\mathcal{H}_{\hbar,\nu},\qquad\widehat{I}_{j,\hbar}\big|_{\mathcal{H}_{\hbar,\nu}}=\lambda_{j,\nu}\,\mathbb{I}.$
Every operator in $\mathcal{A}_{\hbar}$ is block-diagonal with respect to this decomposition, acting as a scalar on each $\mathcal{H}_{\hbar,\nu}$. The joint eigenspaces $\mathcal{H}_{\hbar,\nu}$ therefore play the role of pointer sectors selected by the dissipative dynamics.
Denoting by $P_{\nu}$ the projection onto $\mathcal{H}_{\hbar,\nu}$, the action of the dissipator on off-diagonal density-matrix blocks $\rho_{\nu\mu}=P_{\nu}\rho P_{\mu}$ ($\nu\neq\mu$) takes the form
$\frac{\mathrm{d}}{\mathrm{d}t}\rho_{\nu\mu}=\sum_{k}\Big(\ell_{k,\nu}\overline{\ell_{k,\mu}}-\tfrac{1}{2}|\ell_{k,\nu}|^{2}-\tfrac{1}{2}|\ell_{k,\mu}|^{2}\Big)\rho_{\nu\mu}\;+\;(\text{Hamiltonian commutator term}),$
where $\ell_{k,\nu}$ denotes the scalar eigenvalue of $L_{k,\hbar}$ on $\mathcal{H}_{\hbar,\nu}$, i.e. the value of the functional calculus $\sqrt{\gamma_{k}}\varphi_{k}$ evaluated on the joint eigenvalues $(\lambda_{0,\nu},\dots,\lambda_{n,\nu})$. Whenever these values distinguish different sectors, the real part of the prefactor is strictly negative and the off-diagonal blocks decay exponentially, while the diagonal blocks $\rho_{\nu\nu}$ remain unaffected. This is precisely the mechanism of pure dephasing.
Finally, Theorem 4.24 explains why this decoherence-induced pointer structure is compatible with the classical contact dynamics in the semiclassical limit. For every classical observable $f\in\mathcal{A}_{\mathrm{cl}}$, the Heisenberg generator satisfies
$\mathcal{L}_{\hbar}^{\dagger}(\widehat{f}_{\hbar})=Q_{\hbar}(\{f,h\}_{\alpha})+o(1),\qquad\hbar\to 0,$
so that the evolution of expectation values converges to transport along the contact Hamiltonian flow of $h$. In this sense, the present construction realizes a clean separation of roles: decoherence selects a robust commutative algebra of observables, while the Egorov property guarantees that, within this sector, the quantum evolution reduces to the intended classical contact dynamics. $\diamond$
Example 4.27 (Dephasing and pointer sectors for the linear contact model).
We illustrate the content of the above remark on the linear contact system introduced in Example 4.25.
Fix $\hbar>0$ and consider the commuting family of quantized observables
$\widehat{I}_{0,\hbar}=\widehat{h}_{\hbar},\qquad\widehat{I}_{1,\hbar}=\operatorname{Op}_{\hbar}^{\mathrm{w}}(I_{1}^{\Sigma}),\qquad I_{1}=e^{-z}.$
Let
$\mathcal{A}_{\hbar}\mathrel{\vbox{\hbox{\scriptsize.}\hbox{\scriptsize.}}}=C^{\ast}(\widehat{I}_{0,\hbar},\widehat{I}_{1,\hbar})\subset\mathcal{B}(\mathcal{H}_{\hbar})$
be the commutative $C^{\ast}$-algebra generated by these operators.
Choose a Lindblad operator of dephasing type,
$L_{\hbar}=\sqrt{\gamma}\,\varphi(\widehat{I}_{0,\hbar},\widehat{I}_{1,\hbar}),\qquad\gamma>0,$
where $\varphi$ is a bounded real-valued function. Then $L_{\hbar}\in\mathcal{A}_{\hbar}$ and therefore
$L_{\hbar},A]=[L_{\hbar}^{\dagger},A]=0\qquad\forall\,A\in\mathcal{A}_{\hbar}.$
If, in addition, the Hamiltonian is chosen so that $\widehat{H}_{\hbar},\widehat{I}_{j,\hbar}]=0$ for $j=0,1$, then every observable in $\mathcal{A}_{\hbar}$ is an exact quantum constant of motion:
$\mathcal{L}_{\hbar}^{\dagger}(A)=0\qquad\forall\,A\in\mathcal{A}_{\hbar}.$
Since the generators $\widehat{I}_{0,\hbar}$ and $\widehat{I}_{1,\hbar}$ commute, the Hilbert space admits a decomposition into their joint eigenspaces,
$\mathcal{H}_{\hbar}=\bigoplus_{\nu}\mathcal{H}_{\hbar,\nu},\qquad\widehat{I}_{j,\hbar}\big|_{\mathcal{H}_{\hbar,\nu}}=\lambda_{j,\nu}\,\mathbb{I}.$
In this representation, the Lindblad operator acts diagonally, $L_{\hbar}\big|_{\mathcal{H}_{\hbar,\nu}}=\ell_{\nu}\,\mathbb{I}$, and the dissipative dynamics leaves each sector $\mathcal{H}_{\hbar,\nu}$ invariant.
A direct computation in the Schrödinger picture shows that the off-diagonal density-matrix blocks $\rho_{\nu\mu}=P_{\nu}\rho P_{\mu}$ ($\nu\neq\mu$) decay exponentially in time, while the diagonal blocks $\rho_{\nu\nu}$ are unaffected by the dissipator. Thus the Lindblad evolution implements pure dephasing with respect to the joint spectral decomposition of $(\widehat{I}_{0,\hbar},\widehat{I}_{1,\hbar})$, and the subspaces $\mathcal{H}_{\hbar,\nu}$ play the role of pointer sectors.
Finally, Theorem 4.24 ensures that this quantum superselection structure is compatible with the underlying classical contact dynamics. For any classical observable $f\in\mathcal{A}_{\mathrm{cl}}=\mathrm{Jacobi\text{-}alg}\langle h,I_{1}\rangle$, the Heisenberg generator satisfies
$\mathcal{L}_{\hbar}^{\dagger}(\widehat{f}_{\hbar})=Q_{\hbar}(\{f,h\}_{\alpha})+o(1),\qquad\hbar\to 0,$
so that the evolution of expectation values converges to transport along the contact Hamiltonian flow of $h$. This example shows explicitly how, in dephasing-compatible constructions, environment-induced decoherence selects a robust commutative algebra of observables, while the Egorov property guarantees that the effective dynamics inside this sector reduces to the intended classical contact evolution.
5 An Euler-top-type Poisson–Lie pencil as a bi-Lindblad system
In this section we illustrate how the abstract framework developed above can be implemented in a concrete low-dimensional model of Euler-top type, inspired by the deformed Euler top of Marrero and collaborators [ballesteros2017poisson, G.I.M+2023]. Rather than reproducing every physical feature of the rigid-body Euler top, we focus on the structural ingredients that are essential for our construction: (i) a linear Poisson pencil of Poisson–Lie origin, (ii) a bi-Hamiltonian vector field admitting a Lenard–Magri hierarchy of commuting invariants, (iii) a homogeneous lift yielding a completely integrable contact system with nontrivial dissipated quantities, and (iv) an algebraic setting in which one can build Lindblad generators implementing pure dephasing relative to the quantum integrals. In this sense, the example below should be read as a deformed-Euler-top-type testbed tailored to the bi-Lindblad viewpoint.
Let $\mathfrak{g}=\mathfrak{so}(3)$ and identify $\mathfrak{g}^{\ast}\simeq\mathbb{R}^{3}$ with coordinates $m=(m_{1},m_{2},m_{3})$. The standard Lie–Poisson bracket reads
$\{m_{i},m_{j}\}_{0}=\varepsilon_{ijk}m_{k},$
where $\varepsilon_{ijk}$ is the Levi–Civita symbol, i.e.,
$\{m_{1},m_{2}\}_{0}=-m_{3},\qquad\{m_{1},m_{3}\}_{0}=m_{2},\qquad\{m_{2},m_{3}\}_{0}=-m_{1}.$ (5.1)
For comparison, the classical Euler top is Hamiltonian with respect to this bracket, with Hamiltonian
$H_{0}(m)=\frac{1}{2}\Big(\frac{m_{1}^{2}}{J_{1}}+\frac{m_{2}^{2}}{J_{2}}+\frac{m_{3}^{2}}{J_{3}}\Big),$
where $J_{1},J_{2},J_{3}>0$ are the principal moments of inertia.
A bi-Hamiltonian formulation of Euler-top type dynamics is obtained by endowing $\mathfrak{so}(3)^{*}\simeq\mathbb{R}^{3}$, with coordinates $m=(m_{1},m_{2},m_{3})$, with a second linear Poisson bracket $\{\cdot,\cdot\}_{1}$ compatible with the standard Lie–Poisson bracket $\{\cdot,\cdot\}_{0}$. A second compatible Poisson structure is given by
$\{m_{1},m_{2}\}_{1}=-m_{2},\qquad\{m_{1},m_{3}\}_{1}=m_{3},\qquad\{m_{2},m_{3}\}_{1}=-2m_{1},$ (5.2)
which is again linear and corresponds to the Lie–Poisson structure of $\mathfrak{sl}(2,\mathbb{R})$.
The two brackets $\{\cdot,\cdot\}_{0}$ and $\{\cdot,\cdot\}_{1}$ are compatible, in the sense that for every $\lambda\in\mathbb{R}$ the linear combination
$\{\cdot,\cdot\}_{\lambda}\mathrel{\vbox{\hbox{\scriptsize.}\hbox{\scriptsize.}}}=\{\cdot,\cdot\}_{1}-\lambda\,\{\cdot,\cdot\}_{0}$
is again a Poisson bracket. Equivalently, the associated bivectors $\pi_{0}$ and $\pi_{1}$ define a Poisson pencil
$\pi_{\lambda}\mathrel{\vbox{\hbox{\scriptsize.}\hbox{\scriptsize.}}}=\pi_{1}-\lambda\pi_{0}.$
Rather than considering the full family of Hamiltonian flows associated with the pencil, we fix a distinguished bi-Hamiltonian vector field $X$ on $\mathbb{R}^{3}$ defined by the requirement that it be Hamiltonian with respect to both Poisson structures. Concretely, $X$ is chosen so that there exist two functions $C_{0},C_{1}\in\mathscr{C}^{\infty}(\mathbb{R}^{3})$ satisfying
$X=\pi_{0}^{\sharp}\,\mathrm{d}C_{1}=\pi_{1}^{\sharp}\,\mathrm{d}C_{0}.$ (5.3)
This is the Euler-top-type dynamics that we shall quantize: a coadjoint motion on $\mathfrak{so}(3)^{*}$ admitting two compatible Hamiltonian descriptions and a complete set of commuting integrals generated by the Poisson pencil.
This choice is essential for what follows. The bi-Hamiltonian nature of $X$ ensures the existence of a Lenard–Magri hierarchy of invariants, which in turn gives rise—after homogeneous lift—to a contact integrable system with distinguished dissipated quantities. These quantities will generate the commutative classical algebra $\mathcal{A}_{\mathrm{cl}}$ and, upon quantization, the invariant quantum algebra underlying the dephasing Lindblad dynamics constructed below.
In coordinates, the Poisson pencil associated with $\pi_{\lambda}=\pi_{1}-\lambda\pi_{0}$ reads
$\{m_{1},m_{2}\}_{\lambda}=(\lambda-1)m_{3}-\lambda m_{2},\qquad\{m_{1},m_{3}\}_{\lambda}=(1-\lambda)m_{2}+\lambda m_{3},\qquad\{m_{2},m_{3}\}_{\lambda}=-(1+\lambda)m_{1}.$ (5.4)
A convenient choice of Casimir functions for this pencil is
$C_{0}(m)=-\tfrac{1}{2}\big(m_{1}^{2}+m_{2}^{2}+m_{3}^{2}\big),\qquad C_{1}(m)=m_{1}^{2}+m_{2}m_{3}.$ (5.5)
With this choice, we can define the vector field $X$ given by (5.3). This vector field is the Euler-top-type dynamics that we fix throughout this section. Its trajectories are confined to the common level sets of $(C_{0},C_{1})$, and the pair $(C_{0},C_{1})$ provides a complete set of commuting first integrals for this three-dimensional system.
More generally, the Poisson pencil $\pi_{\lambda}$ underlies a Lenard–Magri hierarchy of commuting invariants $\{H_{k}\}_{k\geq 0}$ satisfying
$\pi_{0}^{\sharp}\mathrm{d}H_{k+1}=\pi_{1}^{\sharp}\mathrm{d}H_{k},\qquad k\geq 0,$
with $(H_{0},H_{1})=(C_{0},C_{1})$. In the present example the hierarchy truncates after these two basic integrals, which already capture the full integrable structure of the flow.
To connect the Lie–Poisson dynamics on $\mathfrak{so}(3)^{*}$ with the homogeneous/contact framework developed in the previous sections, we proceed via an exact symplectic realization of the linear Poisson structure $\pi_{0}$. Concretely, we consider a symplectic manifold $(S,\omega)$ with $\omega=-\mathrm{d}\Theta$ exact, together with a Poisson map
$\Phi:(S,\omega)\longrightarrow(\mathfrak{so}(3)^{*},\pi_{0}).$
Such exact symplectic realizations of linear Poisson manifolds are classical and can be constructed explicitly; see, for instance, [C.F.M2021, Section 6.3]. In particular, for $\mathfrak{so}(3)^{*}\simeq\mathfrak{su}(2)^{*}$ one may take $S=N\times\mathbb{R}^{3}_{m}$, where $N$ is a $3$-manifold equipped with a coframe $\{\vartheta_{1},\vartheta_{2},\vartheta_{3}\}$ satisfying the Maurer–Cartan relations, and define
[…]
The projection $\Phi(n,m)=m$ is then a Poisson map onto $(\mathfrak{so}(3)^{*},\pi_{0})$. In this realization, the classical invariants on $\mathfrak{so}(3)^{*}$ lift naturally by pullback. In particular, for the Casimir functions $C_{a}\in\mathscr{C}^{\infty}(\mathfrak{so}(3)^{*})$ introduced in (5.5), we define
$F_{a}\mathrel{\vbox{\hbox{\scriptsize.}\hbox{\scriptsize.}}}=\Phi^{*}C_{a},\qquad a=0,1,$
namely,
$F_{0}(n,m)=-\tfrac{1}{2}\big(m_{1}^{2}+m_{2}^{2}+m_{3}^{2}\big),\qquad F_{1}(n,m)=m_{1}^{2}+m_{2}m_{3}.$
These lifted functions Poisson-commute on $(S,\omega)$ and generate Hamiltonian vector fields projecting onto the Euler-top-type dynamics on $\mathfrak{so}(3)^{*}$. Note that $\omega=-\mathrm{d}\Theta$ is an exact symplectic form by definition, and thus we can define the Liouville vector field as the unique vector field $\Delta$ on $S$ such that
$\iota_{\Delta}\omega=-\Theta\,.$
On each hypersurface $C$ of $S$ which is transversal to $\Delta$, the restriction $\alpha\mathrel{\vbox{\hbox{\scriptsize.}\hbox{\scriptsize.}}}={\left.\kern-1.2pt\Theta\right|_{C}}$ is a contact form. We can check that
$\Delta=\sum_{i=1}^{3}m_{i}\frac{\partial}{\partial m_{i}}\,$,
and hence $F_{1}$ and $F_{2}$ are homogeneous of degree $2$. Let us define the functions $I_{a}^{\Sigma}\colon U\to\mathbb{R}$, defined on the open subset $U=S\setminus(N\times\{0\})$, given by
$I_{0}^{\Sigma}=\sqrt{-2F_{0}}=\sqrt{m_{1}^{2}+m_{2}^{2}+m_{3}^{2}}\,,\quad I_{1}^{\Sigma}=\sqrt{F_{1}}=\sqrt{m_{1}^{2}+m_{2}m_{3}}\,.$
These functions are homogeneous of degree $1$ and Poisson-commute, namely,
$\{I_{1}^{\Sigma},I_{2}^{\Sigma}\}_{\omega}=0\,$,
where $\{\cdot,\cdot\}_{\omega}$ denotes the Poisson bracket defined by $\omega$. The hypersurface $C=\{m_{1}=1\}\subset U\subset S$ is transverse to $\Delta$. Hence,
$\alpha\mathrel{\vbox{\hbox{\scriptsize.}\hbox{\scriptsize.}}}={\left.\kern-1.2pt\Theta\right|_{C}}=\vartheta_{1}+m_{2}\vartheta_{2}+m_{3}\vartheta_{3}$
is a contact form on $C$. The one-forms
$I_{0}\mathrel{\vbox{\hbox{\scriptsize.}\hbox{\scriptsize.}}}={\left.\kern-1.2ptI_{0}^{\Sigma}\right|_{C}}=\sqrt{1+m_{2}^{2}+m_{3}^{2}}\,,\quad I_{1}\mathrel{\vbox{\hbox{\scriptsize.}\hbox{\scriptsize.}}}={\left.\kern-1.2ptI_{1}^{\Sigma}\right|_{C}}=\sqrt{1+m_{2}m_{3}}\,$,
commute with respect to the Jacobi bracket $\{\cdot,\cdot\}_{\alpha}$ defined by $\alpha$. Thus, in this Euler-top-type example the Jacobi-commutative algebra of dissipated quantities is
$\mathcal{A}_{\mathrm{cl}}=\mathrm{Jacobi\text{-}alg}\langle I_{0}\mathrel{\vbox{\hbox{\scriptsize.}\hbox{\scriptsize.}}}=h,\,I_{1}\rangle\subset\mathscr{C}^{\infty}(C),$
which encodes the integrable structure inherited from the underlying bi-Hamiltonian Poisson pencil.
This construction makes explicit the conceptual separation between the Lie–Poisson data and the contact geometry. The Poisson structures $\pi_{0},\pi_{1}$ on $\mathfrak{so}(3)^{*}$ govern the classical integrable content via the bi-Hamiltonian vector field $X$ and its invariants, while the contact structure on $C$ arises indirectly from homogeneous symplectization of an exact symplectic realization. In particular, the contact form $\alpha=\Theta|_{C}$ is not an independent ingredient but is induced by the primitive $\Theta$ of the symplectic form. Accordingly, the Jacobi-commutative algebra $\mathcal{A}_{\mathrm{cl}}=\mathrm{Jacobi\text{-}alg}\langle I_{0},I_{1}\rangle$ and the dissipated quantities $I_{j}$ are determined entirely by the homogeneous lift of the Poisson-pencil invariants, and do not depend on a special choice of contact representative. These quantities will play a central role in the construction of the dephasing Lindblad dynamics in the next step.
We fix a semiclassical quantization $(\mathcal{H}_{\hbar},Q_{\hbar})$ of $\mathcal{A}_{\mathrm{cl}}$ in the sense of Definition 4.14, implemented by semiclassical Weyl quantization on the homogeneous symplectic manifold $(M,\omega)$ after the homogeneous lift. Concretely, if $f\in\mathcal{A}_{\mathrm{cl}}$ and $f^{\Sigma}$ denotes its $1$-homogeneous lift on $M$ (so that $f^{\Sigma}|_{C}=f$), we set
[…]
In particular, for the Euler-top generators we obtain
[…]
Since $\mathcal{A}_{\mathrm{cl}}$ is Jacobi-commutative and $Q_{\hbar}$ satisfies the Dirac condition, we have for all $f,g\in\mathcal{A}_{\mathrm{cl}}$
$\frac{1}{i\hbar}\big[Q_{\hbar}(f),Q_{\hbar}(g)\big]=Q_{\hbar}(\{f,g\}_{\alpha})+\mathcal{O}(\hbar),$
and therefore
[…]
In particular, the family $\{\widehat{I}_{0,\hbar},\widehat{I}_{1,\hbar}\}$ is asymptotically commuting in the semiclassical limit.
This family generates a distinguished $C^{\ast}$-subalgebra
$\mathcal{A}_{\hbar}\mathrel{\vbox{\hbox{\scriptsize.}\hbox{\scriptsize.}}}=C^{\ast}(\widehat{I}_{0,\hbar},\widehat{I}_{1,\hbar})\subset\mathcal{B}(\mathcal{H}_{\hbar}),$
which is abelian up to semiclassical corrections (and exactly abelian if an exact commuting quantization is chosen). In the present Euler-top setting, $\mathcal{A}_{\hbar}$ is the quantum counterpart of the classical invariant algebra generated by the two integrals $(I_{0},I_{1})$, namely the Hamiltonian of the Poisson pencil dynamics and the quadratic Casimir.
This algebra will play the role of a robust commutative sector of the open quantum dynamics: it is singled out by the underlying integrable geometry and, as we show below, can be preserved pointwise by a suitable choice of Lindblad operators inducing pure dephasing.
When, in addition, the quantization is chosen so that $\widehat{I}_{0,\hbar}$ and $\widehat{I}_{1,\hbar}$ commute exactly (for instance, when $Q_{\hbar}$ is realized in a representation where $I_{0}^{\Sigma}$ and $I_{1}^{\Sigma}$ quantize to commuting multiplication operators, or when the exact commutation relations (4.3) are imposed), the joint spectral theorem applies.
More precisely, since $\widehat{I}_{0,\hbar}$ and $\widehat{I}_{1,\hbar}$ are bounded self-adjoint operators that commute strongly, there exists a joint projection-valued spectral measure $E_{\hbar}(\cdot)$ on $\mathbb{R}^{2}$ and a Borel measure $\mu_{\hbar}$ such that the Hilbert space admits a unitary decomposition as a direct integral
$\mathcal{H}_{\hbar}\simeq\int_{\mathbb{R}^{2}}^{\oplus}\mathcal{H}_{\hbar,\nu}\,\mathrm{d}\mu_{\hbar}(\nu).$
Here $\nu=(\nu_{0},\nu_{1})$ denotes the joint spectral parameter, and $\mathcal{H}_{\hbar,\nu}$ is the corresponding spectral multiplicity space. In this representation, the operators $\widehat{I}_{j,\hbar}$ act as multiplication operators,
$(\widehat{I}_{j,\hbar}\psi)(\nu)=\nu_{j}\,\psi(\nu),\qquad j=0,1,$
so that, fibrewise,
$\widehat{I}_{j,\hbar}\big|_{\mathcal{H}_{\hbar,\nu}}=\nu_{j}\,\mathbf{1}.$
Semiclassically, the joint spectral parameter $\nu=(\nu_{0},\nu_{1})$ labels, up to the usual Bohr–Sommerfeld quantization conditions (see e.g. [GuilleminSternberg1990]), the common level sets $\{I_{0}=\nu_{0},\ I_{1}=\nu_{1}\}$ of the underlying contact integrable system.
The direct-integral fibres $\mathcal{H}_{\hbar,\nu}$ may therefore be interpreted as quantum counterparts of the classical invariant leaves (Liouville tori/cylinders after reduction), providing the quantum “fibres” associated with the classical invariant foliation.
Even without exact commutation, the asymptotic relation (5.6) is sufficient for the present purposes: it ensures that $\mathcal{A}_{\hbar}$ behaves as an approximately classical, asymptotically abelian sector in the semiclassical limit. It is precisely this sector that we will preserve pointwise by a suitable choice of Lindblad operators in the next step.
We now construct Lindblad generators whose dissipative part induces pure dephasing with respect to the joint spectral data of $(\widehat{I}_{0,\hbar},\widehat{I}_{1,\hbar})$, in a way compatible with the Weyl-type assumptions of Theorem 4.24. Fix $N\in\mathbb{N}$, choose rates $\gamma_{k}>0$, and let $\varphi_{k}\colon\mathbb{R}^{2}\to\mathbb{R}$ be bounded smooth functions. We define the dissipative operators by joint functional calculus as
$L_{k,\hbar}\mathrel{\vbox{\hbox{\scriptsize.}\hbox{\scriptsize.}}}=\sqrt{\gamma_{k}}\,\varphi_{k}\!\big(\widehat{I}_{0,\hbar},\widehat{I}_{1,\hbar}\big),\qquad k=1,\dots,N.$ (5.7)
By construction, $L_{k,\hbar}\in\mathcal{A}_{\hbar}$, and therefore each Lindblad operator commutes with every element of $\mathcal{A}_{\hbar}$. In particular,
$L_{k,\hbar},\widehat{I}_{j,\hbar}]=[L_{k,\hbar}^{\dagger},\widehat{I}_{j,\hbar}]=0,\qquad j\in\{0,1\},\ \ k=1,\dots,N.$ (5.8)
Equivalently, $L_{k,\hbar}\in\mathcal{A}_{\hbar}\subset\mathcal{A}_{\hbar}^{\prime}$, since $\mathcal{A}_{\hbar}$ is abelian.
We choose the Hamiltonian part so as to quantize the contact Hamiltonian $h=I_{0}$, namely
$\widehat{H}_{\hbar}=\widehat{h}_{\hbar}+\mathcal{O}(\hbar^{2})\qquad\text{in operator norm}.$ (5.9)
If this is strengthened to the exact commutation condition
$\widehat{H}_{\hbar},\widehat{I}_{0,\hbar}]=[\widehat{H}_{\hbar},\widehat{I}_{1,\hbar}]=0,$ (5.10)
then the hypotheses of Lemma 4.7 are satisfied together with (5.8). As a consequence, the Heisenberg adjoint of the resulting GKSL generator satisfies
$\mathcal{L}_{\hbar}^{\dagger}(\widehat{I}_{0,\hbar})=0,\qquad\mathcal{L}_{\hbar}^{\dagger}(\widehat{I}_{1,\hbar})=0.$ (5.11)
Thus the Euler-top quantum integrals $\widehat{I}_{0,\hbar}$ and $\widehat{I}_{1,\hbar}$ are exact constants of motion (not merely asymptotically conserved).
Assume for simplicity that $\widehat{I}_{0,\hbar}$ and $\widehat{I}_{1,\hbar}$ commute exactly. Then the joint spectral theorem yields a decomposition of the Hilbert space into simultaneous spectral subspaces,
$\mathcal{H}_{\hbar}=\bigoplus_{\nu\in\mathrm{Spec}(\widehat{I}_{0,\hbar},\widehat{I}_{1,\hbar})}\mathcal{H}_{\hbar,\nu},\qquad\widehat{I}_{j,\hbar}\big|_{\mathcal{H}_{\hbar,\nu}}=\nu_{j}\,\mathbf{1},\quad j\in\{0,1\}.$
Let $P_{\nu}$ denote the orthogonal projector onto $\mathcal{H}_{\hbar,\nu}$ and write $\rho_{\nu\mu}\mathrel{\vbox{\hbox{\scriptsize.}\hbox{\scriptsize.}}}=P_{\nu}\rho P_{\mu}$ for the corresponding blocks of a density matrix. Since $L_{k,\hbar}\in\mathcal{A}_{\hbar}$, each Lindblad operator acts as a scalar on each joint eigenspace,
$L_{k,\hbar}P_{\nu}=\ell_{k,\nu}\,P_{\nu},\qquad\ell_{k,\nu}\mathrel{\vbox{\hbox{\scriptsize.}\hbox{\scriptsize.}}}=\sqrt{\gamma_{k}}\,\varphi_{k}(\nu_{0},\nu_{1})\in\mathbb{R}.$
A direct computation shows that the dissipator acts diagonally on the block decomposition. For $\nu\neq\mu$ one finds
$\frac{\mathrm{d}}{\mathrm{d}t}\rho_{\nu\mu}(t)\Big|_{\rm diss}=-\frac{1}{2}\sum_{k=1}^{N}\big(\ell_{k,\nu}-\ell_{k,\mu}\big)^{2}\,\rho_{\nu\mu}(t),$ (5.12)
while the diagonal blocks $\rho_{\nu\nu}$ are unaffected by the dissipator. Hence the off-diagonal coherences decay exponentially,
$\|\rho_{\nu\mu}(t)\|\;\lesssim\;\exp\!\Big(-\tfrac{1}{2}\,t\sum_{k=1}^{N}(\ell_{k,\nu}-\ell_{k,\mu})^{2}\Big)\,\|\rho_{\nu\mu}(0)\|,\qquad\nu\neq\mu,$
whereas the populations $\operatorname{Tr}(\rho P_{\nu})$ are preserved. In other words, the open-system dynamics destroys coherences between states with different values of the Euler-top invariants $(I_{0},I_{1})$ without inducing transitions between the corresponding sectors. This is precisely pure dephasing relative to the commutative algebra $\mathcal{A}_{\hbar}$.
Finally, the bi-Hamiltonian origin of the deformed Euler top provides two (distinct) effective Hamiltonians associated with the underlying Poisson pencil, namely the pair of classical functions $(C_{0},C_{1})$ satisfying
$X=\pi_{0}^{\sharp}\mathrm{d}C_{1}=\pi_{1}^{\sharp}\mathrm{d}C_{0}.$
After homogeneous lift and restriction to the contact hypersurface $C$, this yields two effective contact Hamiltonians
$H_{0}^{\mathrm{eff}},\;H_{1}^{\mathrm{eff}}\;\in\;\mathcal{A}_{\mathrm{cl}}.$
In the present minimal Euler-top pencil, one may simply take $H_{0}^{\mathrm{eff}}=I_{0}$ and $H_{1}^{\mathrm{eff}}=I_{1}$, up to irrelevant affine reparametrizations. Let $\widehat{H}_{0,\hbar}$ and $\widehat{H}_{1,\hbar}$ be corresponding semiclassical quantizations.
We define two Lindblad generators on $\mathcal{B}(\mathcal{H}_{\hbar})$ by
$\mathcal{L}^{(a)}_{\hbar}(\rho)=-\frac{i}{\hbar}[\widehat{H}_{a,\hbar},\rho]+\sum_{k=1}^{N}\Big(L_{k,\hbar}\rho L_{k,\hbar}^{\dagger}-\tfrac{1}{2}\{L_{k,\hbar}^{\dagger}L_{k,\hbar},\rho\}\Big),\qquad a\in\{0,1\},$
using the same family of dephasing operators $L_{k,\hbar}$ defined in (5.7).
By construction, the Lindblad operators are of Weyl type and their principal symbols depend only on the homogeneous lifts $(I_{0}^{\Sigma},I_{1}^{\Sigma})$. Hence condition (3) of Theorem 4.24 is satisfied for both generators. Moreover, the commutation relations (5.8) imply that $\widehat{I}_{0,\hbar}$ and $\widehat{I}_{1,\hbar}$ are common quantum constants of motion for both dynamics. The semiclassical limit of the Hamiltonian commutator then yields the corresponding contact dynamics in the sense of (CC2).
Therefore, the families $\{\mathcal{L}^{(0)}_{\hbar}\}$ and $\{\mathcal{L}^{(1)}_{\hbar}\}$ form a bi-Lindblad pair in the sense of Definition 4.21.
Moreover, for any $\lambda\in[0,1]$, the convex combination
$\mathcal{L}^{(\lambda)}_{\hbar}\mathrel{\vbox{\hbox{\scriptsize.}\hbox{\scriptsize.}}}=(1-\lambda)\mathcal{L}^{(0)}_{\hbar}+\lambda\mathcal{L}^{(1)}_{\hbar}$
is again a GKSL generator and remains contact-compatible with the same invariant commutative sector $\mathcal{A}_{\hbar}$. In this way, the present Euler-top-type Poisson–Lie model realizes a genuine bi-Lindblad pencil: a one-parameter family of open quantum dynamics sharing the same dephasing mechanism and the same quantum integrals, whose classical shadows recover different points of the underlying Poisson pencil.
In this explicit testbed, integrability, dissipation, and decoherence coexist in a geometrically controlled way. Decoherence selects the eigenstates of the quantized classical integrals as pointer states, while the underlying bi-Hamiltonian structure organizes both the classical and quantum dynamics.
More broadly, this Euler-top-type example illustrates a general mechanism underlying the framework developed in this work. A classical bi-Hamiltonian structure, once lifted to a contact system, singles out a distinguished family of dissipated quantities whose Jacobi-commutative algebra encodes the integrable geometry of the flow. Upon semiclassical quantization, these quantities generate a commutative $C^{\ast}$-subalgebra of observables. Choosing the Lindblad data inside its commutant produces a purely dephasing dissipative dynamics, dynamically enforcing a superselection rule aligned with the classical invariants.
In this sense, decoherence does not compete with integrability but rather selects it. The bi-Lindblad structure reflects the Poisson–Lie and bi-Hamiltonian origin of the model: two compatible open quantum evolutions share the same invariant commutative sector, and Egorov-type convergence guarantees that, within this robust sector, the semiclassical Heisenberg dynamics reduces to the intended contact Hamiltonian flow.
Remark 5.1.
In many standard Lindblad models of dephasing, one starts from a prescribed commutative algebra of observables (typically generated by the Hamiltonian or by a chosen family of projectors) and constructs dissipative operators so as to enforce invariance of that algebra under the Heisenberg evolution. From a mathematical viewpoint, this corresponds to selecting a maximal abelian $C^{\ast}$-subalgebra of $\mathcal{B}(\mathcal{H})$ and designing a GKSL generator whose adjoint vanishes on it pointwise.
In the present setting, the direction of the construction is reversed. The commutative $C^{\ast}$-algebra
$\mathcal{A}_{\hbar}=C^{\ast}(\widehat{I}_{0,\hbar},\dots,\widehat{I}_{n,\hbar})$
is not chosen a priori, but is determined intrinsically by the classical geometry: it arises from a Jacobi-commutative algebra of dissipated quantities selected by a bi-Hamiltonian Poisson–Lie structure, lifted to a contact system and subsequently quantized semiclassically. The Lindblad operators are then chosen inside the commutant $\mathcal{A}_{\hbar}^{\prime}$, which guarantees exact invariance of $\mathcal{A}_{\hbar}$ and yields pure dephasing relative to its joint spectral decomposition.
Thus, in the Euler-top-type Poisson–Lie model discussed above, the invariant quantum algebra is not an external input but the quantization of a Jacobi-commutative algebra of classical dissipated quantities. The associated pointer sectors are therefore in canonical correspondence with the leaves of the classical contact foliation singled out by the integrable structure.
This provides a natural setting in which decoherence, integrability, and the semiclassical limit are aligned by construction rather than enforced by model-dependent assumptions: decoherence dynamically selects the very algebra that encodes the classical integrable geometry. $\diamond$
6 Conclusions
We conclude by clarifying the physical meaning of the Lindblad constructions introduced in this work, and in particular their relation to quantum decoherence, integrability, and the emergence of classical behavior in the semiclassical limit.
In the theory of open quantum systems, decoherence refers to the dynamical suppression of quantum coherences due to the interaction with an environment. Mathematically, this manifests itself as the decay of off-diagonal matrix elements of the density operator $\rho(t)$ in a preferred basis, while the corresponding diagonal components (populations) remain invariant or evolve autonomously. The states defining this preferred basis are known as pointer states.
In the framework developed here, the role of pointer states is played by the joint eigenspaces of a distinguished commuting family of quantum observables, namely the quantizations
$\{\widehat{I}_{0,\hbar},\widehat{I}_{1,\hbar},\dots,\widehat{I}_{n,\hbar}\}$
of the classical dissipated quantities singled out by the underlying contact integrable system. By construction, the Lindblad operators $L_{k,\hbar}$ are chosen as Weyl-type operators whose principal symbols depend only on the homogeneous lifts of these quantities, and are implemented as functions of the commuting operators $\widehat{I}_{j,\hbar}$. As a result, the dissipative part of the Lindblad generator leaves invariant the commutative $C^{\ast}$-subalgebra
$\mathcal{A}_{\hbar}\mathrel{\vbox{\hbox{\scriptsize.}\hbox{\scriptsize.}}}=C^{\ast}(\widehat{I}_{0,\hbar},\dots,\widehat{I}_{n,\hbar}),$
while suppressing coherences between distinct joint eigenspaces of $\mathcal{A}_{\hbar}$.
More precisely, upon decomposing the Hilbert space as a direct sum (or, more generally, a direct integral) of joint spectral subspaces
$\mathcal{H}_{\hbar}=\bigoplus_{\lambda}\mathcal{H}_{\hbar,\lambda},\qquad\widehat{I}_{j,\hbar}\big|_{\mathcal{H}_{\hbar,\lambda}}=\lambda_{j}\,\mathbf{1},$
the Lindblad dynamics exponentially damps the off-diagonal blocks $\rho_{\lambda\lambda^{\prime}}(t)$ for $\lambda\neq\lambda^{\prime}$, while leaving the diagonal blocks $\rho_{\lambda\lambda}(t)$ invariant. In this sense, the dissipative dynamics implements an environment-induced superselection rule associated with the integrals of motion.
This mechanism is the direct quantum analog of the classical notion of dissipated quantities in contact geometry. Although the classical contact flow is irreversible, the Jacobi-commutative algebra generated by the quantities $I_{j}$ organizes the dynamics and labels its invariant geometric structures. At the quantum level, this organization reappears as decoherence towards the joint eigenspaces of the corresponding observables.
From a physical perspective, several points are worth emphasising. First, the Lindblad dynamics constructed here describe a class of open quantum systems in which dissipation does not destroy integrability. Instead, dissipation is aligned with the classical invariants, leading to pure dephasing rather than relaxation. While non-generic, this situation is physically relevant in several contexts, including weak coupling to classical noise, measurement-induced decoherence, and engineered reservoirs.
Second, the contact-geometric formulation clarifies the semiclassical meaning of this phenomenon. The contact Hamiltonian flow encodes irreversible classical dynamics, while the Jacobi algebra of dissipated quantities captures the surviving integrable structure. The Lindblad generators constructed in this work provide a quantum realization of this picture: in the Heisenberg picture, the quantum evolution converges in the semiclassical limit to the contact dynamics, while the dissipated quantities become genuine quantum constants of motion.
Third, the bi-Lindblad structures introduced here reflect the underlying bi-Hamiltonian geometry. Just as a classical bi-Hamiltonian system admits a pencil of compatible Poisson structures and Hamiltonians, the quantum system admits a family of compatible Lindblad generators whose convex combinations remain completely positive and trace preserving. This establishes a precise correspondence between classical integrable hierarchies and families of admissible open quantum dynamics.
The Euler-top-type Poisson–Lie example analyzed in this work should therefore be understood as a structural testbed rather than as a literal model of rigid body dynamics. Its role is to demonstrate, in a concrete and low-dimensional setting, how integrability, dissipation, and decoherence can coexist in a geometrically controlled way. More broadly, our results suggest that contact geometry provides a natural classical language for describing the semiclassical limit of certain Markovian open quantum systems, in which decoherence selects classical invariants rather than equilibria.
This perspective opens the door to systematic constructions of open quantum models in which dissipation, integrability, and semiclassical consistency are aligned by design, rather than imposed by ad hoc assumptions.
Conflict of Interest Statement.
The authors declare no known competing financial interests or personal relationships that could have appeared to influence the work reported in this manuscript.
Data Availability Statement.
No datasets were generated or analyzed during the current study. All mathematical derivations are contained within the article.

### v2

@@
 Abstract
-This paper develops a bridge between bi-Hamiltonian structures of Poisson–Lie type, contact Hamiltonian dynamics, and the Gorini–Kossakowski–Sudarshan–Lindblad (GKSL) formalism for quantum open systems.
+This paper develops a bridge between bi-Hamiltonian structures of Poisson–Lie type, contact Hamiltonian dynamics, and the Gorini–Kossakowski–Sudarshan–Lindblad (GKSL) formalism for open quantum systems.
 On the classical side, we consider bi-Hamiltonian systems defined by a Poisson pencil with non-trivial invariants.
-Using an exact symplectic realization, these invariants are lifted and projected onto a contact manifold, yielding a completely integrable contact Hamiltonian system in terms of dissipated quantities and a Jacobi-commutative algebra of observables.
+Using an exact symplectic realization, these invariants are lifted and projected onto a contact manifold, yielding a completely integrable contact Hamiltonian system and a Jacobi-commutative algebra of observables.
 On the quantum side, we introduce a class of contact-compatible Lindblad generators: GKSL evolutions whose dissipative part preserves a commutative $C^{\ast}$-subalgebra generated by the quantizations of the classical dissipated quantities, and whose Hamiltonian part admits an Egorov-type semiclassical limit to the contact dynamics.
 This construction provides a mathematical mechanism compatible with the semiclassical limit for pure dephasing, compatible with integrability and contact dissipation.
-An explicit Euler-top-type Poisson–Lie pencil, inspired by deformed Euler top models, is developed as a fully worked-out example illustrating the resulting bi-Lindblad structure and its semiclassical behavior.
+An explicit Poisson–Lie pencil, inspired by deformed Euler top models, is developed as a fully worked-out example illustrating the resulting bi-Lindblad structure and its semiclassical behavior.
 Keywords: open quantum systems, semiclassical limits, integrable systems, Lindblad equation, GKSL formalism, bi-Hamiltonian, contact geometry.
@@
 $X=\Pi_{0}^{\sharp}(\mathrm{d}H_{0})=\Pi_{1}^{\sharp}(\mathrm{d}H_{1})\,.$
-Compatibility ensures the existence of a recursion operator (on a symplectic leaf where one of the Poisson structures is non-degenerate) and hierarchies of commuting Hamiltonians, which in turn give rise to rich algebraic and geometric structures and, under suitable regularity assumptions, to complete integrability in the sense of Liouville; see, for example, [M.M1984, C.M.P1993, M.C.F+1997, F.M.P2000, Fernandes1994, G.M.S2003].
+Compatibility ensures the existence of a recursion operator (on a symplectic leaf where one of the Poisson structures is non-degenerate) and hierarchies of commuting Hamiltonians, which in turn give rise to rich algebraic and geometric structures and, under suitable regularity assumptions, to complete integrability in the sense of Liouville; see, for example, [46, 8, 45, 21, 22, 26].
 A natural and geometrically significant source of bi-Hamiltonian systems is provided by Poisson–Lie groups.
 A Poisson tensor $\pi$ on a Lie group $G$ is called multiplicative if the group multiplication $m\colon G\times G\longrightarrow G$ is a Poisson map from $(G\times G,\pi\oplus\pi)$ to $(G,\pi)$.
-Multiplicative Poisson structures and their infinitesimal counterparts, Lie bialgebras, were introduced by Drinfel’d in the context of quantum groups [Drinfeld1988, Drinfeld1990] and further developed in the foundational work of Semenov–Tian–Shansky on the classical $r$-matrix formalism, Poisson pencils and factorization dynamics [STS1983, STS1985].
-The global geometry of Poisson–Lie groups, including dressing transformations and symplectic groupoids, was elucidated in [LuWeinstein1990].
+Multiplicative Poisson structures and their infinitesimal counterparts, Lie bialgebras, were introduced by Drinfel’d in the context of quantum groups [19, 20] and further developed in the foundational work of Semenov–Tian–Shansky on the classical $r$-matrix formalism, Poisson pencils and factorization dynamics [57, 56].
+The global geometry of Poisson–Lie groups, including dressing transformations and symplectic groupoids, was elucidated in [44].
 When two multiplicative Poisson tensors $\pi_{0},\pi_{1}$ on $G$ are compatible, the resulting Poisson pencil $\pi_{\lambda}=\pi_{1}-\lambda\pi_{0},\,\lambda\in\mathbb{R}$ encodes a rich class of bi-Hamiltonian vector fields on $G$.
@@
 then $X$ is a bi-Hamiltonian vector field whose integrability properties reflect the algebraic structure of the underlying Lie bialgebra and classical $r$-matrix.
-This mechanism underlies many classical integrable models on Lie groups and Poisson–Lie groups [STS1983, STS1985, KS2004, KS2008, LuWeinstein1990].
-On the other hand, contact geometry provides a natural framework for describing non-Hamiltonian classical dynamics in a plethora of contexts, including certain dissipative systems and thermodynamically relevant flows; see, e.g., [d.L2019a, d.L2019b, G.G2022a, G.G2024a, Grabowski2013a, Lopez-Gordon2024].
-There is a one-to-one correspondence between contact manifolds and homogeneous symplectic manifolds, which has been known for a while [Arnold1978, L.M1987], and whose applications to contact Hamiltonian dynamics have been recently analyzed in detail by Grabowska and Grabowski [Grabowski2013a, G.G2022a, G.G2023, G.G2024a].
-We have exploited this correspondence in order to establish a novel characterisation of completely integrable contact Hamiltonian systems as completely integrable Hamiltonian systems (in the standard Liouville sense) that satisfy certain homogeneity conditions [C.d.L+2023, Lopez-Gordon2024, Colombo2025].
+This mechanism underlies many classical integrable models on Lie groups and Poisson–Lie groups [57, 56, 37, 38, 44].
+On the other hand, contact geometry provides a natural framework for describing non-Hamiltonian classical dynamics in a plethora of contexts, including certain dissipative systems and thermodynamically relevant flows; see, e.g., [14, 15, 27, 29, 43].
+There is a one-to-one correspondence between contact manifolds and homogeneous symplectic manifolds, which has been known for a while [1, 39], and whose applications to contact Hamiltonian dynamics have been recently analyzed in detail by Grabowska and Grabowski [31, 27, 28, 29].
+We have exploited this correspondence in order to establish a novel characterisation of completely integrable contact Hamiltonian systems as completely integrable Hamiltonian systems (in the standard Liouville sense) that satisfy certain homogeneity conditions [10, 43, 11].
+At this stage, it is natural to ask why one should restrict to homogeneous symplectic realizations projectable to contact manifolds, rather than considering a generic symplectic realization of an open quantum system.
+Although a classical limit can be obtained from such generic symplectic realizations, these constructions are highly non-unique and do not provide a canonical description of dissipation.
+Requiring the realization to be homogeneous and projectable to a contact structure selects a distinguished geometric class in which dissipation, contraction of phase-space volume, and irreversibility are encoded intrinsically.
+In this framework, the effective non-unitary dynamics arises as the projection of a unitary flow on a higher-dimensional homogeneous symplectic manifold, yielding a robust and universal semiclassical limit.
+The contact-compatible structure is therefore not an additional assumption, but the minimal geometric setting capturing the correct physics of open quantum systems.
 A key point, which we emphasise here, is that:
@@
   • the homogeneity is an independent geometric ingredient that guarantees that the dynamics and its commuting integrals can be projected onto a contact manifold.
-In particular, in [Colombo2025], it is shown, in concrete examples, that one may take a Poisson structure on the symplectization compatible with the canonical one, use the corresponding recursion operator to construct a maximal family of homogeneous functions in involution, and then “project” these functions to a $(2n+1)$-dimensional contact manifold as functions in Jacobi-involution.
-What turns out to be too rigid for contact integrability is not Poisson compatibility on the symplectization, but the use of compatible Jacobi structures (see [NunesdaCosta1998a, M.M.P1999, C.D.N2010, P.N2003, Petalidou2002]) directly on the original contact manifold: compatible Jacobi pairs do not yield a recursion operator with enough independent eigenvalues to produce a full set of contact integrals.
+In particular, in [11], it is shown, in concrete examples, that one may take a Poisson structure on the symplectization compatible with the canonical one, use the corresponding recursion operator to construct a maximal family of homogeneous functions in involution, and then “project” these functions to a $(2n+1)$-dimensional contact manifold as functions in Jacobi-involution.
+What turns out to be too rigid for contact integrability is not Poisson compatibility on the symplectization, but the use of compatible Jacobi structures (see [53, 47, 52, 55]) directly on the original contact manifold: compatible Jacobi pairs do not yield a recursion operator with enough independent eigenvalues to produce a full set of contact integrals.
 This paper aims to unify bi-Hamiltonian Poisson geometry with the homogeneous/contact correspondence and to connect the resulting classical framework with the quantum realm.
@@
 To avoid conceptual ambiguity, we stress a fundamental distinction:
-  • Systems defined on Poisson–Lie groups, and on any of their symplectic or exact symplectic realizations, are entirely classical Hamiltonian systems.
+  • Systems defined on a Poisson–Lie group $G$, and on any of its symplectic or exact symplectic realizations, are entirely classical Hamiltonian systems.
 Their integrability is governed by Poisson, Poisson–Nijenhuis and bi-Hamiltonian structures in the usual sense.
@@
 The trivial symplectization of $(C,\alpha,h)$ is given by $(M=C\times(0,+\infty),\omega=\mathrm{d}(r\alpha),H=rh)$, where $r$ denotes the global coordinate of $(0,+\infty)$.
-There is a Lie algebra isomorphism between the Jacobi bracket $\{\cdot,\cdot\}_{\alpha}$ and the Poisson bracket $\{\cdot,\cdot\}_{\omega}$, which implies that conserved quantities for $(C,\alpha,h)$ are conserved quantities for $(M,\omega,H)$.
-A symplectic form $\omega$ on a manifold $M$ is exact (i.e, there exists a one-form $\theta$ such that $\mathrm{d}\theta=-\omega$) if and only if there exists a Liouville vector field $\Delta$ on $M$ such that $\mathcal{L}_{\Delta}\omega=\omega$.
+There is a Lie algebra isomorphism between the Jacobi bracket $\{\cdot,\cdot\}_{\alpha}$ and the Poisson bracket $\{\cdot,\cdot\}_{\omega}$, which implies that dissipated quantities for $(C,\alpha,h)$ are in one-to-one correspondence with conserved quantities for $(M,\omega,H)$.
+A symplectic form $\omega$ on a manifold $M$ is exact (i.e., there exists a one-form $\theta$ such that $\mathrm{d}\theta=-\omega$) if and only if there exists a Liouville vector field $\Delta$ on $M$ such that $\mathcal{L}_{\Delta}\omega=\omega$.
 Indeed, one can define $\theta=-\iota_{\Delta}\omega$.
@@
 By quantum dissipation we refer to the irreversible, non-unitary component of the reduced dynamics of an open quantum system, described at the Markovian level by a GKSL (Lindblad) generator.
-This dissipative term arises from tracing out environmental degrees of freedom and accounts for phenomena such as decoherence, relaxation and thermalisation [BreuerPetruccione].
-While such dissipation has no direct analog in closed Hamiltonian mechanics, it admits effective classical limits.
+This dissipative term arises from tracing out environmental degrees of freedom and accounts for phenomena such as decoherence, relaxation and thermalisation [6].
+While such dissipation has no direct analog in classical Hamiltonian mechanics, it admits effective classical limits.
 In the present work we exhibit a class of GKSL generators whose semiclassical limit is naturally described by contact Hamiltonian dynamics, with dissipated quantities playing the role of robust invariants of the resulting reduced evolution.
@@
 From the viewpoint of physics, contact systems appear precisely as effective or reduced descriptions of systems interacting with an environment, where certain degrees of freedom are not explicitly modelled.
-Quantum open systems [BreuerPetruccione, Lindblad1976], governed by Lindblad master (GKSL) equations, arise from a conceptually analogous mechanism: a unitary evolution on a larger Hilbert space is reduced to a non-unitary dynamics on the subsystem of interest.
+Quantum open systems [6, 42], governed by Lindblad master (GKSL) equations, arise from a conceptually analogous mechanism: a unitary evolution on a larger Hilbert space is reduced to a non-unitary dynamics on the subsystem of interest.
 This parallel suggests that the contact dynamics obtained from the homogeneous projection of a Poisson–Lie system should admit a natural quantum analog.
@@
 On this exact symplectic manifold we exploit bi-Hamiltonian techniques to build a family of commuting integrals.
-Homogeneity then allows us to project the dynamics and its integrals to a contact manifold, obtaining a completely integrable contact system in the sense of [C.d.L+2023, Lopez-Gordon2024].
+Homogeneity then allows us to project the dynamics and its integrals to a contact manifold, obtaining a completely integrable contact system in the sense of [10, 43].
 The main question we address is whether one can construct an open quantum dynamics naturally associated with this classical framework, whose semiclassical limit reproduces the contact Hamiltonian flow.
@@
 $X=\pi_{0}^{\sharp}(\mathrm{d}H_{0})=\pi_{1}^{\sharp}(\mathrm{d}H_{1})$
-is bi-Hamiltonian on $(G,\pi_{0},\pi_{1})$.
+is bi-Hamiltonian.
 Suppose that $(G,\pi_{0})$ admits an exact symplectic realization $(M,\theta)$ and that $X$ lifts to a homogeneous Hamiltonian system $(M,\theta,H)$, with $H$ of degree $1$ with respect to the Liouville vector field.
@@
     its conserved observables $\widehat{I}_{0},\dots,\widehat{I}_{n}$ are quantizations of the classical dissipated quantities $I_{0},\dots,I_{n}$?
-Our answer is affirmative. We introduce the notion of a contact-compatible Lindblad generator: a GKSL dynamics whose Hamiltonian and dissipative parts are organized so that
+Our answer is affirmative (see Theorem 4.24). We introduce the notion of a contact-compatible Lindblad generator: a GKSL dynamics whose Hamiltonian and dissipative parts are organized so that
   • the Heisenberg-picture evolution admits a well-defined semiclassical limit on the contact phase space $(C,\alpha)$, reproducing the contact flow of $h$;
@@
 We also introduce a notion of bi-Lindblad structure, inspired by bi-Hamiltonian geometry: a pair of Lindblad generators whose convex combinations remain Lindbladian and whose common invariant observables encode the classical integrals of a bi-Hamiltonian system on a Poisson–Lie group.
-This provides a geometric bridge between Poisson–Lie geometry, contact integrable dynamics and physically admissible quantum dissipation.
-An Euler-top-type Poisson–Lie system, inspired by the deformed Euler top introduced in [ballesteros2017poisson, G.I.M+2023], is developed here as a fully explicit example illustrating the classical–quantum correspondence between bi-Hamiltonian contact dynamics and Lindblad open-system evolution.
+This provides a geometric bridge between Poisson–Lie geometry, contact integrable dynamics and physically admissible quantum dissipation (see Proposition 4.23).
+An Euler-top-type Poisson–Lie system, inspired by the deformed Euler top introduced in [3, 33], is developed here as a fully explicit example illustrating the classical–quantum correspondence between bi-Hamiltonian contact dynamics and Lindblad open-system evolution.
 Structure of the paper.
@@
 2 Bi-Hamiltonian geometry and contact integrability
-This section reviews classical bi-Hamiltonian geometry and homogeneous Hamiltonian systems on exact symplectic manifolds, and recalls the notion of contact integrability introduced in [C.d.L+2023, Lopez-Gordon2024].
+This section reviews classical bi-Hamiltonian geometry and homogeneous Hamiltonian systems on exact symplectic manifolds, and recalls the notion of contact integrability introduced in [10, 43].
 In this section we work entirely in the classical realm: Poisson, symplectic and contact structures live on smooth manifolds, and no quantization procedure is involved.
 The contact systems and homogeneous Hamiltonian systems constructed here will later serve as the classical (or semiclassical) phase-space models for the GKSL generators introduced in Section 4.
-On the one hand, bi-Hamiltonian geometry is entirely formulated in the Poisson (or symplectic) category: compatibility of Poisson structures leads to recursion operators and hierarchies of commuting Hamiltonians [M.M1984, M.C.F+1997, F.M.P2000, Fernandes1994, C.M.P1993, G.M.S2003].
+On the one hand, bi-Hamiltonian geometry is entirely formulated in the Poisson (or symplectic) category: compatibility of Poisson structures leads to recursion operators and hierarchies of commuting Hamiltonians [46, 45, 21, 22, 8, 26].
 On the other hand, homogeneity is a property with respect to the Liouville vector field on an exact symplectic manifold.
-In our context these two ingredients are logically independent: bi-Hamiltonian (or Poisson–Nijenhuis) structures will be used to generate integrals in involution in the usual symplectic sense, while homogeneity will be the geometric mechanism that allows us to project these integrals to a $(2n+1)$-dimensional contact manifold via symplectization.
+These two ingredients are independent: bi-Hamiltonian (or Poisson–Nijenhuis) structures will be used to generate integrals in involution in the usual symplectic sense, while homogeneity will be the geometric mechanism that allows us to project these integrals to a contact manifold of one dimension less.
 2.1 Bi-Hamiltonian structures
@@
 $X=\Pi_{0}^{\sharp}(\mathrm{d}H_{0})=\Pi_{1}^{\sharp}(\mathrm{d}H_{1}).$
-On any symplectic leaf where $\Pi_{0}$ is nondegenerate, one can introduce the recursion operator (see, for instance, [M.M1984, M.C.F+1997, F.M.P2000, Fernandes1994])
+On any symplectic leaf where $\Pi_{0}$ is non-degenerate, one can introduce the recursion operator (see, for instance, [46, 45, 21, 22])
 $N\mathrel{\vbox{\hbox{\scriptsize.}\hbox{\scriptsize.}}}=\Pi_{1}^{\sharp}\circ(\Pi_{0}^{\sharp})^{-1}.$
-When the Nijenhuis torsion of $N$ vanishes, spectral invariants of $N$ yield hierarchies of commuting Hamiltonians, providing the standard bi-Hamiltonian framework for Liouville integrability [C.M.P1993, G.M.S2003].
-This picture can be recast in the more general language of Poisson–Nijenhuis manifolds, Jacobi–Nijenhuis manifolds and their reductions [K.M1990, Wade2001, NunesdaCosta1998a, M.M.P1999, P.N2003, C.D.N2010, D.M.P2011], but in this section we restrict ourselves to the basic Poisson setting.
+The Nijenhuis torsion of $N$ vanishes, and its eigenvalues are commuting functions in involution providing the standard bi-Hamiltonian framework for Liouville integrability [8, 26].
 2.2 Exact symplectic manifolds and homogeneity
 Throughout the paper we adopt the convention that an exact symplectic manifold $(M,\theta)$ carries the symplectic form $\omega\mathrel{\vbox{\hbox{\scriptsize.}\hbox{\scriptsize.}}}=-\mathrm{d}\theta$.
-Hamiltonian vector fields are defined by $\iota_{X_{H}}\omega=\mathrm{d}H$, and the Liouville vector field $\Delta$ is characterized by $\iota_{\Delta}\omega=-\theta$.
+Hamiltonian vector fields are defined by $\iota_{X_{H}}\omega=\mathrm{d}H$, and the Liouville vector field $\Delta$ is uniquely determined by $\iota_{\Delta}\omega=-\theta$.
 Let $(M,\theta)$ be an exact symplectic manifold with symplectic form $\omega=-\mathrm{d}\theta$. The one-form $\theta$ is called a symplectic potential.
 Definition 2.3.
-The Liouville vector field $\Delta$ is defined by $\iota_{\Delta}\omega=-\theta$.
 A tensor field $A$ on $M$ is said to be homogeneous of degree $w\in\mathbb{R}$ (or simply $w$-homogeneous) if
@@
 In other words, the Hamiltonian flow of a $1$-homogeneous function is invariant under the Liouville flow.
-Following [C.d.L+2023, Lopez-Gordon2024], we say that a homogeneous Hamiltonian system is integrable if it admits a full set of independent, homogeneous first integrals in involution.
+Following [10, 43], we say that a homogeneous Hamiltonian system is integrable if it admits a full set of independent, homogeneous first integrals in involution.
 Definition 2.5.
@@
 In our later construction, we will start from a Poisson (or Poisson–Lie) bi-Hamiltonian system and, when an exact symplectic realization is available, we will require the Hamiltonian and its integrals to be $1$-homogeneous with respect to the corresponding Liouville vector field.
-Compatibility of Poisson tensors is responsible for integrability in the Liouville sense, whereas homogeneity is what will allow us to relate this symplectic picture with a contact Hamiltonian system on a $(2n+1)$-dimensional manifold via symplectization.
+Compatibility of Poisson tensors is responsible for integrability in the Liouville sense, whereas homogeneity is what will allow us to project this symplectic picture to a contact Hamiltonian system.
 These are therefore two independent geometric requirements that must be arranged simultaneously in our applications.
@@
 $\iota_{R}\mathrm{d}\alpha=0,\qquad\alpha(R)=1.$
-Given a function $h\in\mathscr{C}^{\infty}(C)$, the associated contact Hamiltonian vector field $X_{h}$ is defined by (see, e.g., [d.L2019b, G.G2022a, Lopez-Gordon2024])
+Given a function $h\in\mathscr{C}^{\infty}(C)$, the associated contact Hamiltonian vector field $X_{h}$ is defined by (see, e.g., [15, 27, 43])
 $\begin{cases}\alpha(X_{h})=h,\\[5.69054pt] \iota_{X_{h}}\mathrm{d}\alpha=(\mathrm{d}h)(R)\,\alpha-\mathrm{d}h.\end{cases}$
 The triple $(C,\alpha,h)$ is called a contact Hamiltonian system.
-The contact form $\alpha$ induces a Jacobi structure $(\Lambda,R)$ on $C$ [Lichnerowicz1977a, Lichnerowicz1978, d.L.M+2003, G.U2003], and hence a Jacobi bracket on $\mathscr{C}^{\infty}(C)$ given by
+The contact form $\alpha$ induces a Jacobi structure $(\Lambda,R)$ on $C$ [40, 41, 16], and hence a Jacobi bracket on $\mathscr{C}^{\infty}(C)$ given by
 $\{f,g\}_{\alpha}\mathrel{\vbox{\hbox{\scriptsize.}\hbox{\scriptsize.}}}=X_{f}(g)-g\,R(f)=\Lambda(\mathrm{d}f,\mathrm{d}g)+f\,R(g)-g\,R(f),\qquad f,g\in\mathscr{C}^{\infty}(C).$
@@
 $\{f,h\}_{\alpha}=0.$
-This notion was introduced in [G.G.M+2020a] and further developed in [d.L2019a, C.d.L+2023, Lopez-Gordon2024] as the natural replacement, in contact geometry, of conserved quantities in Hamiltonian dynamics.
+This notion was introduced in [24] and further developed in [14, 10, 43, 25] as the natural replacement, in contact geometry, of conserved quantities in Hamiltonian dynamics.
 Definition 2.7.
@@
 Reordering if necessary, we may and shall assume that $h$ itself is included in the family $\{f_{i}\}$.
-In this language, the contact Liouville–Arnol’d theorem proved in [C.d.L+2023] extends the classical symplectic theory [Arnold1978, Audin2004] to the contact category.
-In particular, completely integrable contact systems admit local action–angle type coordinates adapted to the Reeb dynamics and to the dissipated quantities.
+In this language, the contact Liouville–Arnol’d theorem proved in [10] extends the classical symplectic theory [1, 2] to the contact category.
+In particular, completely integrable contact systems admit local action–angle type coordinates adapted to the contact Hamiltonian dynamics and to a certain foliation defined by the dissipated quantities.
 2.4 Symplectization and the contact–homogeneous correspondence
@@
 $\Sigma\colon C^{\Sigma}\to C,\qquad\Sigma(x,r)=x,$
-is a symplectization in the sense of [D.L.M1991, I.L.M+1997, G.I.M+2004, Petalidou2002, Lopez-Gordon2024]: there exists a nowhere-vanishing function $\sigma$ (here $\sigma=r$) such that $\theta=\sigma\,\Sigma^{\ast}\alpha$.
+is a symplectization in the sense of [13, 35, 30, 55, 43]: there exists a nowhere-vanishing function $\sigma$ (here $\sigma=r$) such that $\theta=\sigma\,\Sigma^{\ast}\alpha$.
 Given a contact Hamiltonian $h\in\mathscr{C}^{\infty}(C)$, one associates a $1$-homogeneous Hamiltonian $H^{\Sigma}$ on $C^{\Sigma}$ by
 $H^{\Sigma}(x,r)\mathrel{\vbox{\hbox{\scriptsize.}\hbox{\scriptsize.}}}=-\,\sigma(x,r)\,h(x)=-r\,h(x),$
-following the conventions in [C.d.L+2023, Lopez-Gordon2024]. This defines a homogeneous Hamiltonian system $(C^{\Sigma},\theta,H^{\Sigma})$.
-The following result summarizes the correspondence between integrability in the contact and homogeneous symplectic settings (see [C.d.L+2023, Lopez-Gordon2024, G.I.M+2004, Petalidou2002] for proofs and variants, and [NunesdaCosta1998a, M.M.P1999, P.N2003, C.D.N2010] for related Jacobi and Jacobi–Nijenhuis structures).
+following the conventions in [10, 43].
+This defines a homogeneous Hamiltonian system $(C^{\Sigma},\theta,H^{\Sigma})$.
+The Liouville vector field of $\theta$ in this case is given by $\Delta=r\partial_{r}$.
+The following result summarizes the correspondence between integrability in the contact and homogeneous symplectic settings (see [10, 43, 30, 55] for proofs and variants)
 Proposition 2.8 (Contact–homogeneous correspondence).
@@
     The map $f\mapsto f^{\Sigma}\mathrel{\vbox{\hbox{\scriptsize.}\hbox{\scriptsize.}}}=-\sigma\,\Sigma^{\ast}f$ defines a bijection between $\mathscr{C}^{\infty}(C)$ and the space of $1$-homogeneous functions on $C^{\Sigma}$.
+This bijection has an associated correspondence between the associated Hamiltonian vector fields given by
+    $X_{f^{\Sigma}}=-X_{f}+rR(f)\partial_{r}\,$,
+    where $X_{f^{\Sigma}}$ (resp. $X_{f}$) denotes the Hamiltonian vector field of $f^{\Sigma}$ (resp. $f$) with respect to $\omega$ (resp. $\alpha$).
   2. 2.
     For all $f,g\in\mathscr{C}^{\infty}(C)$, the Poisson bracket on $(C^{\Sigma},\theta)$ and the Jacobi bracket on $(C,\alpha)$ satisfy
-    $\{f^{\Sigma},g^{\Sigma}\}_{\theta}=\{f,g\}_{\alpha}^{\Sigma}.$
+    $\{f^{\Sigma},g^{\Sigma}\}_{\theta}=-\{f,g\}_{\alpha}^{\Sigma}.$
   3. 3.
@@
 Remark 2.9.
-The above result is classical.
-Items (1)–(3) follow from the standard identification between Jacobi manifolds and homogeneous Poisson manifolds under symplectization; see [G.I.M+2004, Petalidou2002, D.L.M1991].
-The equivalence of contact integrability and homogeneous integrability in item (4) is established in [C.d.L+2023, Lopez-Gordon2024].
-We record the proposition here for completeness and to fix notation for the subsequent sections, where this correspondence becomes the bridge to the quantum (Lindbladian) setting.
-Remark 2.10.
 The previous proposition shows that the study of completely integrable contact Hamiltonian systems can be reduced, via symplectization, to the study of homogeneous integrable systems on exact symplectic manifolds.
 In particular, one may try to import bi-Hamiltonian techniques to the contact setting by first constructing compatible Poisson structures on the symplectic realization and then projecting the resulting commuting integrals to the contact manifold.
-In contrast, starting directly from compatible Jacobi structures on $(C,\alpha)$ [NunesdaCosta1998a, M.M.P1999, P.N2003, C.D.N2010] and using the associated Jacobi–Nijenhuis recursion operators does not, in general, lead to a maximal family of dissipated quantities in involution for a contact system, as we shall discuss later in connection with a no-go result of “Jacobi/bi-Hamiltonian” type.
+In contrast, starting directly from compatible Jacobi structures on $(C,\alpha)$ [53, 47, 52, 9] and using the associated Jacobi–Nijenhuis recursion operators does not lead to a maximal family of dissipated quantities in involution for a contact system.
+In fact, in [11] we have proven a no-go theorem.
 This dichotomy will be crucial when we pass from classical homogeneous and contact dynamics to quantum GKSL generators in the subsequent sections.
@@
 Let $C=\mathbb{R}^{3}(q,p,z)$ endowed with the standard contact form
-[…]
+$\alpha\mathrel{\vbox{\hbox{\scriptsize.}\hbox{\scriptsize.}}}=\mathrm{d}z-p\,\mathrm{d}q.$
 Its Reeb vector field is $R=\partial_{z}$.
-We adopt the convention for the contact Hamiltonian vector field $X_{h}$:
+The contact Hamiltonian vector field $X_{h}$ is determined by
 $\alpha(X_{h})=h,\qquad\iota_{X_{h}}\mathrm{d}\alpha=(\mathrm{d}h)(R)\,\alpha-\mathrm{d}h.$
-For $\alpha=\mathrm{d}z-p\,\mathrm{d}q$ this yields the coordinate expressions
-$\dot{q}=-\frac{\partial h}{\partial p},\qquad\dot{p}=\frac{\partial h}{\partial q}+p\frac{\partial h}{\partial z},\qquad\dot{z}=h-p\frac{\partial h}{\partial p}.$
+This yields the coordinate expression
+$X_{h}=-\frac{\partial h}{\partial p}\frac{\partial}{\partial q}+\left(\frac{\partial h}{\partial q}+p\frac{\partial h}{\partial z}\right)\frac{\partial}{\partial p}+\left(h-p\frac{\partial h}{\partial p}\right)\frac{\partial}{\partial z}\,.$
 Consider the linear contact Hamiltonian
 $h(q,p,z)\mathrel{\vbox{\hbox{\scriptsize.}\hbox{\scriptsize.}}}=z-p.$
-Then
-$\frac{\partial h}{\partial p}=-1,\qquad\frac{\partial h}{\partial q}=0,\qquad\frac{\partial h}{\partial z}=1,$
-and the resulting contact dynamics is
-$\dot{q}=1,\qquad\dot{p}=p,\qquad\dot{z}=z.$
-Hence the flow is explicitly solvable:
-$q(t)=q_{0}+t,\qquad p(t)=p_{0}e^{t},\qquad z(t)=z_{0}e^{t}.$
-Define the observable
-$I_{1}(q,p,z)\mathrel{\vbox{\hbox{\scriptsize.}\hbox{\scriptsize.}}}=e^{-z}.$
-Since $R(h)=1$, the condition for $I_{1}$ to be a dissipated quantity is
-$X_{h}(I_{1})+(Rh)\,I_{1}=0.$
-A direct computation gives
-$X_{h}(I_{1})=-\,\frac{\partial I_{1}}{\partial z}=-(-e^{-z})=e^{-z},$
-and therefore
-$X_{h}(I_{1})+(Rh)\,I_{1}=e^{-z}-e^{-z}=0.$
-Equivalently, in terms of the Jacobi bracket,
-$\{I_{1},h\}_{\alpha}=\frac{\partial I_{1}}{\partial z}-\frac{\partial I_{1}}{\partial p}+I_{1}=(-e^{-z})-0+e^{-z}=0.$
-Hence $I_{1}$ is a dissipated quantity in the sense of Definition 2.7.
-Note that $I_{1}$ is not constant along the contact flow:
-$X_{h}(I_{1})=e^{-z}\neq 0,$
-which illustrates that dissipated quantities (Jacobi integrals) need not be conserved along the flow, unless additional conditions such as $R(I_{1})=0$ are satisfied.
-Moreover, $I_{0}\mathrel{\vbox{\hbox{\scriptsize.}\hbox{\scriptsize.}}}=h$ is trivially dissipated since $\{h,h\}_{\alpha}=0$, and we have
-$\{I_{0},I_{1}\}_{\alpha}=0.$
-The differentials
-$\mathrm{d}h=\mathrm{d}z-\mathrm{d}p,\qquad\mathrm{d}I_{1}=-e^{-z}\,\mathrm{d}z$
-are linearly independent on a dense open subset of $C$, so
+Then the resulting contact dynamics is
+$X_{h}=\frac{\partial}{\partial q}+p\frac{\partial}{\partial p}+z\frac{\partial}{\partial z}\,$,
+whose flow $\varphi_{t}$ is given by
+$\varphi_{t}\left(q,p,z\right)=\left(q+t,pe^{t},ze^{t}\right)\,.$
+Since $R(h)=1$, the condition for a function $I_{i}$ to be a dissipated quantity is
+$0=\{h,I_{i}\}=X_{h}I_{i}-(Rh)I_{i}=X_{h}I_{i}-I_{i}=\frac{\partial I_{i}}{\partial q}+p\frac{\partial I_{i}}{\partial p}+z\frac{\partial I_{i}}{\partial z}-I_{i}\,.$
+For instance, $I_{1}=p,I_{2}=z$ and $I_{3}=e^{q}$ are dissipated quantities.
+Note that they are not constant along the flow of $X_{h}$.
+In fact, $I_{i}\circ\varphi_{t}=e^{t}I_{i},\,i=1,2,3$.
+This illustrates that dissipated quantities (Jacobi integrals) need not be conserved along the flow, unless additional conditions such as $R(I_{i})=0$ are satisfied.
+Moreover, $I_{0}\mathrel{\vbox{\hbox{\scriptsize.}\hbox{\scriptsize.}}}=h$ is trivially dissipated since the Jacobi bracket is skew-symmetric. The differentials
+$\mathrm{d}h=\mathrm{d}z-\mathrm{d}p,\qquad\mathrm{d}I_{1}=\mathrm{d}z$
+are linearly independent on $C$, so
 $\operatorname{rank}\,\langle\mathrm{d}h,\mathrm{d}I_{1}\rangle=2\geq n=1,$
 as required for complete contact integrability in dimension $3$.
-Therefore $(C,\alpha,h)$ is a completely integrable contact Hamiltonian system with $n=1$, admitting the two commuting dissipated quantities $I_{0}=h$ and $I_{1}=e^{-z}$.
-Finally, this contact system arises from symplectization with our sign conventions. Let
-[…]
-and define the $1$-homogeneous Hamiltonian
-$H^{\Sigma}\mathrel{\vbox{\hbox{\scriptsize.}\hbox{\scriptsize.}}}=-\,r\,h=-r(z-p).$
-Then $(C^{\Sigma},\theta,H^{\Sigma})$ is an exact symplectic homogeneous Hamiltonian system, and the contact dynamics of $(C,\alpha,h)$ is recovered by restriction to the hypersurface $\{r=1\}$, in accordance with Proposition 2.8.
+Therefore $(C,\alpha,h)$ is a completely integrable contact Hamiltonian system with $n=1$, admitting the two commuting dissipated quantities $I_{0}=h$ and $I_{1}=z$.
+The associated homogeneous integrable system $(M,\theta,H,F)$ is given by
+$M=C\times\mathbb{R}_{>0}\,,\quad\theta=r\alpha=r\mathrm{d}z-rp\mathrm{d}q\,,\quad H=-rh=rp-rz\,,\quad F=-rI_{1}=-rp\,$,
+where $r$ is the canonical coordinate of $\mathbb{R}_{>0}$.
+The Liouville vector field is given by $\Delta=r\partial_{r}$ The Hamiltonian vector field of $H$ with respect to $\omega=-\mathrm{d}\theta$ reads
+$X_{H}=-X_{h}+rR(h)\frac{\partial}{\partial r}=\frac{\partial}{\partial q}+p\frac{\partial}{\partial p}+z\frac{\partial}{\partial z}+r\frac{\partial}{\partial r}\,.$
+One can verify that it is 0-homogeneous, i.e., $\Delta,X_{H}]=0$.
+Remark 2.11.
+In order to simplify the exposition, we will consider only co-oriented contact manifolds and their trivial symplectizations.
+More generally, one can define a contact distribution on a manifold $M$ as a maximally non-integrable sub-bundle $D\subset{\mathsf{T}}M$ of corank $1$.
+A contact form is a one-form $\alpha$ such that $\ker\alpha=D$, which may only exist locally, and is never unique (as $f\alpha$ is also a contact form for any non-vanishing function $f$).
+A more general notion of symplectization is an $\mathbb{R}^{\times}$-principal bundle $P\to M$ endowed with a symplectic form which is homogeneous with respect to the principal action.
+For more details, see [7, 31, 27, 29]
 3 Bi-Hamiltonian Poisson–Lie systems and contact realizations
@@
 $m\colon(G\times G,\pi_{0}\oplus\pi_{0})\longrightarrow(G,\pi_{0})$
-is Poisson. In this case $(G,\pi_{0})$ is a Poisson–Lie group; see, e.g., [Lichnerowicz1977, Lichnerowicz1977a, Lichnerowicz1978, D.L.M1991, G.U2003].
+is Poisson. In this case $(G,\pi_{0})$ is a Poisson–Lie group; see [3] and references therein.
 Given a Poisson tensor $\pi$ on $G$, we denote by
@@
 the associated bundle map. For $H\in\mathscr{C}^{\infty}(G)$, the Hamiltonian vector field $X_{H}$ is defined by $X_{H}=\pi^{\sharp}(\mathrm{d}H)$.
-On a symplectic manifold $(M,\omega)$ we define the Hamiltonian vector field $X_{H}$ by $\iota_{X_{H}}\omega=\mathrm{d}H$.
-Definition 3.1.
-A bi-Hamiltonian Poisson–Lie system on a Lie group $G$ consists of:
-  • two Poisson–Lie structures $\pi_{0},\pi_{1}\in\Gamma(\wedge^{2}{\mathsf{T}}G)$ on $G$;
-  • two Hamiltonians $H_{0},H_{1}\in\mathscr{C}^{\infty}(G)$;
-such that:
-  1. 1.
-    $\pi_{0}$ and $\pi_{1}$ are compatible Poisson tensors in the sense that every linear combination
-    $\pi_{\lambda}\mathrel{\vbox{\hbox{\scriptsize.}\hbox{\scriptsize.}}}=\pi_{1}-\lambda\,\pi_{0},\qquad\lambda\in\mathbb{R},$
-    is a Poisson tensor on $G$;
-  2. 2.
-    the corresponding Hamiltonian vector fields coincide,
-    $X\mathrel{\vbox{\hbox{\scriptsize.}\hbox{\scriptsize.}}}=\pi_{0}^{\sharp}(\mathrm{d}H_{0})=\pi_{1}^{\sharp}(\mathrm{d}H_{1}).$
-Thus $(\pi_{0},\pi_{1})$ defines a Poisson pencil on $G$, and on each symplectic leaf of $\pi_{0}$ (or $\pi_{1}$) the restriction of $(\pi_{0},\pi_{1})$ and $X$ yields an ordinary bi-Hamiltonian system in the sense of Section 2.
-The general bi-Hamiltonian theory on Poisson manifolds (and in particular on symplectic leaves) is developed in [C.M.P1993, M.C.F+1997, F.M.P2000, Fernandes1994, G.M.S2003].
-Many classical integrable systems, such as the Euler top and its deformations, admit realizations of this type on Poisson–Lie groups [M.P1996, NunesdaCosta1998, Petalidou2000].
+Many classical integrable systems, such as the Euler top and its deformations, admit realizations of this type on Poisson–Lie groups [49, 51, 54].
 In what follows, we shall assume that the bi-Hamiltonian dynamics is (Liouville) integrable on a suitable open dense subset of $G$, and concentrate on how this structure lifts to an exact symplectic manifold and then induces a contact integrable system via homogeneity.
@@
 Let $(G,\pi_{0})$ be a Poisson manifold.
-A symplectic realization of $(G,\pi_{0})$ is a symplectic manifold $(M,\omega)$ together with a Poisson map $J\colon(M,\omega)\to(G,\pi_{0})$ whose image is an open subset of $G$; see, for instance, [L.M1987, D.L.M1991, G.I.M+2004, Lopez-Gordon2024].
+A symplectic realization of $(G,\pi_{0})$ is a symplectic manifold $(M,\omega)$ together with a Poisson map $J\colon(M,\omega)\to(G,\pi_{0})$ whose image is an open subset of $G$; see [12].
 We are interested in the case in which $(M,\omega)$ is exact.
@@
 This requirement, however, does not constrain the underlying Lie–Poisson or Poisson–Lie structure itself.
-For a large class of Lie–Poisson systems, including linear and solvable cases, canonical exact realizations are provided by cotangent bundles ${\mathsf{T}}^{\ast}Q$, which are always non-compact and carry a natural exact symplectic structure.
+For a large class of Lie–Poisson systems, including linear and solvable cases, canonical exact realizations are provided by cotangent bundles ${\mathsf{T}}^{\ast}Q$, with the canonical symplectic structure $\omega_{Q}=-\mathrm{d}\theta$ defined by the canonical one-form $\theta_{Q}$ (which is an exact symplectic structure by definition).
+In that case, the Liouville vector field is the generator of homotheties on the fibers of ${\mathsf{T}}^{\ast}Q$.
 The example in Section 3.4 illustrates this situation explicitly.
 More generally, since our construction only requires working on open dense subsets where the homogeneous integrals are independent, the assumption of exactness should be regarded as a natural geometric convenience rather than a substantive restriction of the Poisson–Lie framework.
-Let $\Delta$ be the Liouville vector field of $(M,\theta)$, defined by
-$\iota_{\Delta}\omega=-\theta.$
-As recalled in Section 2, a function $H\in\mathscr{C}^{\infty}(M)$ is said to be $1$-homogeneous if
-$\mathcal{L}_{\Delta}H=H,$
-and a triple $(M,\theta,H)$ with $H$ 11-homogeneous is a homogeneous Hamiltonian system. We shall also use the notion of homogeneous integrability from Definition 2.3.
 To connect the Poisson–Lie data on $G$ with contact geometry, we make the following structural assumption.
@@
   • an exact symplectic realization $(M,\theta,J)$ of $(G,\pi_{0})$, with $\dim M=2(n+1)$;
-  • a $1$-homogeneous Hamiltonian $H\in\mathscr{C}^{\infty}(M)$ with respect to the Liouville vector field $\Delta$ of $(M,\theta)$,
+  • a $1$-homogeneous Hamiltonian $H\in\mathscr{C}^{\infty}(M)$ with respect to the Liouville vector field $\Delta$ of $(M,\theta)$, namely,
+    $\Delta H=H\,,\quad\iota_{\Delta}\omega=-\theta\,$,
 such that:
@@
     the Hamiltonian vector field $X_{H}$ on $(M,\omega=-\mathrm{d}\theta)$ projects to the bi-Hamiltonian vector field $X$ on $G$, i.e.
-    $TJ\circ X_{H}=X\circ J;$
+    ${\mathsf{T}}J\circ X_{H}=X\circ J;$
   3. 3.
     there exist $n+1$ independent, $1$-homogeneous first integrals in involution for $(M,\theta,H)$, making $(M,\theta,H)$ a homogeneous integrable system in the sense of Definition 2.3.
-In many examples of interest, the Poisson pencil $(\pi_{0},\pi_{1})$ admits a symplectic realization $(M,\omega)$ on which the pullbacks of $\pi_{0},\pi_{1}$ give rise to a Poisson–Nijenhuis structure $(\omega,N)$ on $M$ [K.M1990, Wade2001, D.M.P2011, B.K.M2022, B.K.M2022a, B.K.M2024b, Konyaev2021].
+In many examples of interest, the Poisson pencil $(\pi_{0},\pi_{1})$ admits a symplectic realization $(M,\omega)$ on which the pullbacks of $\pi_{0},\pi_{1}$ give rise to a Poisson–Nijenhuis structure $(\omega,N)$ on $M$ [36, 17, 4, 5].
 The eigenvalues of $N$ then provide a natural candidate for homogeneous integrals in involution, and the homogeneity of $H$ can often be arranged by a suitable choice of exact symplectic structure (or by considering an appropriate homogeneous extension).
@@
 Under Assumption 3.4, we can apply the general homogeneous–contact correspondence of Section 2 to obtain a contact integrable system. We spell this out carefully.
-We first recall a standard fact about exact symplectic manifolds (see, e.g., [L.M1987, D.L.M1991, Lopez-Gordon2024]).
+We first recall a standard fact about exact symplectic manifolds (see, e.g., [39, 13, 43]).
 Lemma 3.5.
@@
 $\iota_{\Delta}(\omega^{n+1})=-(n+1)\,\theta\wedge\omega^{n}=-(n+1)(-1)^{n}\,\theta\wedge(\mathrm{d}\theta)^{n}.$
-Hence $\Theta\mathrel{\vbox{\hbox{\scriptsize.}\hbox{\scriptsize.}}}=\theta\wedge(\mathrm{d}\theta)^{n}$ is nonvanishing at every point where $\Delta\neq 0$.
+Hence $\Theta\mathrel{\vbox{\hbox{\scriptsize.}\hbox{\scriptsize.}}}=\theta\wedge(\mathrm{d}\theta)^{n}$ is non-vanishing at every point where $\Delta\neq 0$.
 Now let $C\subset M$ be a hypersurface transverse to $\Delta$.
@@
 Let $(G,\pi_{0},\pi_{1};H_{0},H_{1})$ and $(M,\theta,H)$ be as in Assumption 3.4.
-Let $\Delta$ be the Liouville vector field of $(M,\theta)$ and let $C\subset M$ be a hypersurface transverse to $\Delta$, contained in the open set where $H\neq 0$.
+Let $\Delta$ be the Liouville vector field of $(M,\theta)$ and let $C\subset M$ be a hypersurface transverse to $\Delta$, contained in the open subset where $H\neq 0$.
 Define
@@
 First, by Lemma 3.5, the restriction $\alpha=\theta|_{C}$ is a contact form on $C$, so $(C,\alpha)$ is a co-oriented contact manifold.
-Since $H$ is $1$-homogeneous with respect to $\Delta$, we have $\mathcal{L}_{\Delta}H=H$. Writing $\varphi_{s}$ for the flow of $\Delta$, this implies
+By assumption, $H$ is $1$-homogeneous with respect to $\Delta$, i.e., $\mathcal{L}_{\Delta}H=H$. This implies that
 $\frac{\mathrm{d}}{\mathrm{d}s}H(\varphi_{s}(x))=(\mathcal{L}_{\Delta}H)(\varphi_{s}(x))=H(\varphi_{s}(x)),$
-so, on any integral curve of $\Delta$,
+where $\varphi_{s}$ denotes the flow of $\Delta$. Hence,
 $H(\varphi_{s}(x))=e^{s}H(x).$
-In particular, on the open region where $H\neq 0$, the level sets of $H$ are transversal to the Liouville flow, and every Liouville orbit intersects each nonzero level set exactly once.
+In particular, on the open subset where $H\neq 0$, the level sets of $H$ are transversal to the Liouville flow, and every Liouville orbit intersects each nonzero level set exactly once.
 By construction $C$ is a hypersurface transverse to $\Delta$ and contained in $\{H\neq 0\}$, so every Liouville orbit in the region of interest intersects $C$ exactly once.
@@
 where $R$ is the Reeb vector field of $(C,\alpha)$.
-Standard results on the contactization of homogeneous Hamiltonian systems (see, for instance, [D.L.M1991, I.L.M+1997, G.I.M+2004, Petalidou2002, Lopez-Gordon2024]) ensure that:
+Standard results on the contactization of homogeneous Hamiltonian systems ensure that:
   • Since $H$ is $1$-homogeneous, one has $\mathcal{L}_{\Delta}H=H$, and hence $X_{H}$ is 0-homogeneous (equivalently $\Delta,X_{H}]=0$).
@@
 Applying this to $F_{0},\dots,F_{n}$ we see that each $I_{j}$ is a dissipated quantity for the contact Hamiltonian system $(C,\alpha,h)$, i.e. $\{I_{j},h\}_{\alpha}=0$ for all $j$.
-Moreover, since the $F_{j}$ are in Poisson-involution on $M$ and the Jacobi bracket on $C$ is obtained by reduction along the Liouville flow (again, see [D.L.M1991, G.I.M+2004, Petalidou2002, Lopez-Gordon2024]), the restrictions $I_{j}$ are in Jacobi-involution:
+Moreover, since the $F_{j}$ are in Poisson-involution on $M$ and the Jacobi bracket on $C$ is obtained by reduction along the Liouville flow (see [55]), the restrictions $I_{j}$ are in Jacobi-involution:
 $\{I_{j},I_{k}\}_{\alpha}=0,\qquad\forall\,j,k\in\{0,\dots,n\}.$
@@
 At any $x\in U\cap C$ the differentials $\mathrm{d}F_{0},\dots,\mathrm{d}F_{n}$ are linearly independent in ${\mathsf{T}}^{\ast}_{x}M$.
-Restricting to ${\mathsf{T}}_{x}C\subset{\mathsf{T}}_{x}M$ we obtain the differentials $\mathrm{d}I_{0},\dots,\mathrm{d}I_{n}$, which remain linearly independent on ${\mathsf{T}}_{x}C$: otherwise there would be a nontrivial linear relation among the $\mathrm{d}F_{j}$ on ${\mathsf{T}}_{x}M$.
+Restricting to ${\mathsf{T}}_{x}C\subset{\mathsf{T}}_{x}M$ we obtain the differentials $\mathrm{d}I_{0},\dots,\mathrm{d}I_{n}$, which remain linearly independent on ${\mathsf{T}}_{x}C$: otherwise there would be a non-trivial linear relation among the $\mathrm{d}F_{j}$ on ${\mathsf{T}}_{x}M$.
 Thus,
@@
 The construction above crucially uses Poisson compatibility and homogeneity on the exact symplectic realization $(M,\theta)$, rather than compatibility of Jacobi structures directly on the contact manifold $(C,\alpha)$.
-As shown in our previous work on the Jacobi/bi-Hamiltonian no-go phenomenon [C.d.L+2023, Lopez-Gordon2024] (building on [NunesdaCosta1998a, M.M.P1999, N.P2002, C.D.N2010, P.N2003, Petalidou2002]), compatible Jacobi pairs on a contact manifold do not provide enough independent eigenvalues of the associated recursion operator to recover a full set of contact integrals.
+As shown in our previous work on the Jacobi/bi-Hamiltonian no-go phenomenon [10, 43] (building on [53, 47, 50, 9, 52, 55]), compatible Jacobi pairs on a contact manifold do not provide enough independent eigenvalues of the associated recursion operator to recover a full set of contact integrals.
 By passing through the homogeneous Poisson picture on $(M,\theta)$ we circumvent this obstruction: the bi-Hamiltonian structure lives in the Poisson category, while the contact dynamics is obtained by projection along the Liouville flow.
+This is summarized in the following diagram:
 […]
-Figure 1: From a bi-Hamiltonian Poisson–Lie system on $(G,\pi_{0},\pi_{1})$ to a contact integrable system $(C,\alpha,h)$ via an exact, homogeneous symplectic realization $(M,\theta,H)$ and a Liouville-transverse hypersurface $C$.
+that, we pass from a bi-Hamiltonian Poisson–Lie system on $(G,\pi_{0},\pi_{1})$ to a contact integrable system $(C,\alpha,h)$ via an exact, homogeneous symplectic realization $(M,\theta,H)$ and a Liouville-transverse hypersurface $C$.
 This is precisely the classical geometric framework in which, in the next section, we shall construct Lindblad generators whose semiclassical limit reproduces the contact flow of $h$ and whose conserved quantum observables quantize the classical dissipated quantities $I_{0},\dots,I_{n}$.
@@
 Proof.
-This is an immediate consequence of Proposition 3.7, together with the fact that Liouville-transverse hypersurfaces can be chosen inside any given nonzero level set of $H$ in the region where the homogeneous integrals $(F_{0},\dots,F_{n})$ are independent.
+This is an immediate consequence of Proposition 3.6, together with the fact that Liouville-transverse hypersurfaces can be chosen inside any given non-zero level set of $H$ in the region where the homogeneous integrals $(F_{0},\dots,F_{n})$ are independent.
 ∎
-3.4 Example: a Poisson–Nijenhuis bi-Hamiltonian system with two nontrivial contact integrals
-We illustrate Proposition 3.7 with the Poisson–Nijenhuis example already discussed in Section 2, now interpreted as a homogeneous bi-Hamiltonian system producing a contact integrable system with the correct number of dissipated quantities for the $3$-dimensional ($n=1$) contact setting.
+3.4 Example: a Poisson–Nijenhuis bi-Hamiltonian system with two non-trivial contact integrals
 Let $\mathrm{Aff}^{+}(1)$ denote the connected component of the identity of the group of orientation–preserving affine transformations of the real line, with group law $(a,b)\cdot(a^{\prime},b^{\prime})=(aa^{\prime},\,b+a\,b^{\prime})$.
@@
 $M={\mathsf{T}}^{\ast}\mathbb{R}^{2}\cong\mathbb{R}^{4}$
-with coordinates $(x^{1},x^{2},p_{1},p_{2})$ and canonical one-form $\theta=p_{i}\,\mathrm{d}x^{i}$. Throughout we use the convention $\omega=-\mathrm{d}\theta$. The associated Poisson tensor is
+with coordinates $(x^{1},x^{2},p_{1},p_{2})$ equipped with the canonical one-form $\theta=p_{i}\,\mathrm{d}x^{i}$ and the canonical symplectic form $\omega=-\mathrm{d}\theta=\mathrm{d}x^{i}\wedge\mathrm{d}p_{i}$.
+The associated Poisson tensor is
 $\Lambda=\frac{\partial}{\partial x^{1}}\wedge\frac{\partial}{\partial p_{1}}+\frac{\partial}{\partial x^{2}}\wedge\frac{\partial}{\partial p_{2}}.$
@@
 which are functionally independent on $U$ and pairwise commuting for the two Poisson structures.
-The Liouville vector field
-$\Delta=p_{i}\frac{\partial}{\partial p_{i}}$
-satisfies $\iota_{\Delta}\omega=-\theta$, and $\lambda_{1},\lambda_{2}$ are both $1$-homogeneous:
+The Liouville vector field reads
+$\Delta=p_{i}\frac{\partial}{\partial p_{i}}\,$,
+and $\lambda_{1},\lambda_{2}$ are both $1$-homogeneous:
 $\mathcal{L}_{\Delta}\lambda_{i}=\lambda_{i},\qquad i=1,2.$
@@
 $H\mathrel{\vbox{\hbox{\scriptsize.}\hbox{\scriptsize.}}}=\lambda_{1}+\lambda_{2}=p_{1}+p_{2}x^{2}.$
-Then $\mathcal{L}_{\Delta}H=H$ and, since $\{\lambda_{1},\lambda_{2}\}_{\Lambda}=0$, each $\lambda_{i}$ is a first integral of $X_{H}$.
-Thus $(M,\theta,H)$ is a $1$-homogeneous integrable system in the sense of Definition 2.3.
+such that each $\lambda_{i}$ is a first integral of $X_{H}$. Thus $(M,\theta,H)$ is a $1$-homogeneous integrable system in the sense of Definition 2.3.
 Moreover, $X_{H}$ is bi-Hamiltonian:
@@
 $C\mathrel{\vbox{\hbox{\scriptsize.}\hbox{\scriptsize.}}}=\{r=1\}\subset M.$
-Since $\Delta=r\partial_{r}$ is transverse to the level sets of $r$, the hypersurface $C$ is transverse to the Liouville flow. The restriction
+Since $\Delta=r\partial_{r}$ is transverse to the level sets of $r$, the hypersurface $C$ is transverse to the Liouville flow. By Lemma 3.5, the restriction
 $\alpha\mathrel{\vbox{\hbox{\scriptsize.}\hbox{\scriptsize.}}}=\theta|_{C}=\mathrm{d}z-p\,\mathrm{d}q$
-is a contact form, since
+is a contact form. Indeed,
 $\alpha\wedge\mathrm{d}\alpha=(\mathrm{d}z-p\,\mathrm{d}q)\wedge(-\mathrm{d}p\wedge\mathrm{d}q)\neq 0.$
@@
 $\operatorname{rank}\langle\mathrm{d}I_{0},\mathrm{d}I_{1},\mathrm{d}I_{2}\rangle=2\geq n=1$
-on a dense open subset (in fact everywhere). Therefore $(C,\alpha,h)$ is a completely integrable contact system in the sense of Definition 2.4.
+on a dense open subset (in fact everywhere). Therefore $(C,\alpha,h)$ is a completely integrable contact system in the sense of Definition 2.4. As a matter of fact, it is the system from Example 2.10.
 Remark 3.9 (The $3$-dimensional case $n=1$).
@@
 We now move to the quantum setting and introduce a class of Lindblad generators that are compatible with the contact integrable systems constructed in Section 3.
-For definiteness, we work with finite-dimensional Hilbert spaces; in this context, $\mathcal{B}(\mathcal{H})$ denotes the full $C^{\ast}$-algebra of bounded operators on $\mathcal{H}$ (which coincides with the space of all linear operators), all operator topologies coincide, and domain issues do not arise.
+For simplicity’s sake, we work with finite-dimensional Hilbert spaces; in this context, $\mathcal{B}(\mathcal{H})$ denotes the full $C^{\ast}$-algebra of bounded operators on $\mathcal{H}$ (which coincides with the space of all linear operators), all operator topologies coincide, and domain issues do not arise.
 Most of the algebraic definitions and arguments below extend to separable infinite-dimensional Hilbert spaces under standard technical assumptions on domains and continuity of the corresponding quantum dynamical semigroups.
@@
 Fixing an orthonormal basis $\{|j\rangle\}_{j=1}^{d}$ (often selected by the measurement scheme or by the spectral decomposition of a reference Hamiltonian), we write $\rho_{jk}=\langle j|\rho|k\rangle$.
-The coherences are the off-diagonal entries $\rho_{jk}$ with $j\neq k$, which encode relative phases and interference effects, while the diagonal entries $\rho_{jj}$ represent populations, i.e., classical probabilities of outcomes in that basis (in the standard terminology of open quantum systems, see e.g., [BreuerPetruccione]).
+The coherences are the off-diagonal entries $\rho_{jk}$ with $j\neq k$, which encode relative phases and interference effects, while the diagonal entries $\rho_{jj}$ represent populations, i.e., classical probabilities of outcomes in that basis (in the standard terminology of open quantum systems, see e.g., [6]).
 A dynamics exhibits decoherence (in that basis) when the off-diagonal entries are damped in time, typically driving $\rho(t)$ towards a state that is approximately diagonal in the chosen basis.
 In open quantum systems, decoherence is not an ad hoc ingredient but a robust consequence of the coupling to an environment: while the total evolution on system-plus-environment may be unitary, the reduced dynamics on the system alone is generally non-unitary and may suppress coherences.
-There are two equivalent ways to represent the same quantum evolution. It is convenient to introduce a unified notation for expectation values. For an observable $A$ at time $t$, we define
-$\mathrm{E}(A)(t)\mathrel{\vbox{\hbox{\scriptsize.}\hbox{\scriptsize.}}}=\operatorname{Tr}\big(\rho(t)A\big)$
-in the Schrödinger picture, or equivalently
-$\mathrm{E}(A)(t)\mathrel{\vbox{\hbox{\scriptsize.}\hbox{\scriptsize.}}}=\operatorname{Tr}\big(\rho_{0}A(t)\big)$
-in the Heisenberg picture.
-In the Schrödinger picture, the state evolves according to
+There are two equivalent ways to represent the same quantum evolution. In the Schrödinger picture, the state evolves according to
 $\dot{\rho}(t)=\mathcal{L}(\rho(t)),\qquad\rho(0)=\rho_{0},$
@@
 $\mathcal{L}_{\hbar}^{\dagger}(\widehat{f}_{\hbar})\;\longrightarrow\;Q_{\hbar}(\{f,h\}_{\alpha})$
-on the commutative subalgebra generated by the integrals.
+on the commutative subalgebra generated by the functions in involution.
 This viewpoint allows one to identify conserved quantities and invariant subalgebras directly at the level of the observable algebra $\mathcal{B}(\mathcal{H})$, which is particularly well suited to the geometric framework developed here.
-By contrast, the Schrödinger picture is often preferable when addressing state-preparation problems, entropy production, contractivity properties of the quantum dynamical semigroup, or the detailed long-time behavior of the density operator $\rho(t)$; see, for instance, [Lindblad1976, BreuerPetruccione] and the references therein.
+By contrast, the Schrödinger picture is often preferable when addressing state-preparation problems, entropy production, contractivity properties of the quantum dynamical semigroup, or the detailed long-time behavior of the density operator $\rho(t)$; see, for instance, [42, 6] and the references therein.
 Both pictures are, of course, mathematically equivalent, but emphasize different structural aspects of open quantum dynamics.
@@
 It is well known that such $\mathcal{L}$ are precisely the generators of one-parameter semigroups $\Phi_{t}=e^{t\mathcal{L}},\,t\geq 0$, of completely positive and trace-preserving (CPTP) maps on $\mathcal{S}(\mathcal{H})$.
-Conversely, any norm-continuous CPTP semigroup on $\mathcal{S}(\mathcal{H})$ has a generator of the form (4.1) (see [BreuerPetruccione], [Lindblad1976] for instance).
+Conversely, any norm-continuous CPTP semigroup on $\mathcal{S}(\mathcal{H})$ has a generator of the form (4.1) (see [6, 42] for instance).
 We now recall the Heisenberg picture associated with a Lindblad generator.
@@
 for all $\rho\in\mathcal{B}(\mathcal{H})$ and all $A\in\mathcal{B}(\mathcal{H})$.
-Since the trace pairing $\langle\rho,A\rangle=\mathrm{Tr}(\rho A)$ is nondegenerate, $\mathcal{L}^{\dagger}$ is uniquely determined by this relation.
+Since the trace pairing $\langle\rho,A\rangle=\mathrm{Tr}(\rho A)$ is non-degenerate, $\mathcal{L}^{\dagger}$ is uniquely determined by this relation.
 Proposition 4.3.
@@
 $\mathrm{Tr}\big(\mathcal{L}(\rho)\,A\big)=\mathrm{Tr}\Big(\rho\Big[\frac{i}{\hbar}[H,A]+\sum_{k=1}^{N}\big(L_{k}^{\dagger}AL_{k}-\tfrac{1}{2}\{L_{k}^{\dagger}L_{k},A\}\big)\Big]\Big),$
-and, by the defining relation of $\mathcal{L}^{\dagger}$ and the nondegeneracy of the trace pairing, this shows (4.2). ∎
+and, by the defining relation of $\mathcal{L}^{\dagger}$ and the non-degeneracy of the trace pairing, this shows (4.2). ∎
 We now formalise the notion of conserved observable for a given Lindblad dynamics.
@@
 $\frac{\mathrm{d}}{\mathrm{d}t}\,\mathrm{Tr}(\rho(t)\,O)=\mathrm{Tr}(\dot{\rho}(t)\,O)=\mathrm{Tr}(\mathcal{L}(\rho(t))\,O)=\mathrm{Tr}(\rho(t)\,\mathcal{L}^{\dagger}(O))\,.$
-Since the trace product is non-degenerate, the right-hand side vanishes for every $\rho$ if and only if $\mathcal{L}^{\dagger}(O)$. ∎
+Since the trace product is non-degenerate, the right-hand side vanishes for every $\rho$ if and only if $\mathcal{L}^{\dagger}(O)=0$. ∎
 This motivates the following definition:
@@
 Remark 4.6.
-For any Lindblad operator, every matrix proportional to the identity is a constant of the motion, since $\mathbb{I}\in\ker\mathcal{L}^{\dagger}$.
+For any Lindblad operator, every matrix proportional to the identity is a constant of the motion, since $\mathbb{I}\in\ker\mathcal{L}^{\dagger}$ for any $\mathcal{L}^{\dagger}$ of the form (4.2).
 Hence, constants of the motion always form an unital subalgebra of $\mathcal{B}(\mathcal{H})$, of dimension at least $1$.
@@
 $L_{k}^{\dagger}OL_{k}=OL_{k}^{\dagger}L_{k},$
-since $L_{k},O]=[L_{k}^{\dagger},O]=0$ implies $O$ commutes with both $L_{k}$ and $L_{k}^{\dagger}$, and hence with $L_{k}^{\dagger}L_{k}$. Similarly,
+since $L_{k},O]=[L_{k}^{\dagger},O]=0$ implies that $O$ commutes with $L_{k}^{\dagger}L_{k}$. Similarly,
 $\{L_{k}^{\dagger}L_{k},O\}=L_{k}^{\dagger}L_{k}O+OL_{k}^{\dagger}L_{k}=2OL_{k}^{\dagger}L_{k}.$
@@
 and is therefore a constant of motion for the Lindblad dynamics.
-This will be our basic mechanism to quantize the classical dissipated quantities of a contact integrable system: we will construct Lindblad generators whose Hamiltonian and dissipative parts lie in the commutant of the algebra generated by the quantum integrals.
+This will be our basic mechanism to quantize the classical dissipated quantities of a contact integrable system: we will construct Lindblad generators whose Hamiltonian and dissipative parts lie in the commutant of the algebra generated by the quantum constants of the motion.
 $\diamond$
@@
 quantum coherences) while leaving the diagonal entries (the populations) unchanged.
-Although dissipation is present, no energy exchange with the environment occurs.
+Although dissipation is present, no energy exchange with the environment occurs, i.e., $H$ is a constant of the motion.
 While dissipation is most naturally described in the Schrödinger picture at the level of states, the analysis of conserved quantities is more transparent in the Heisenberg picture.
@@
 Finally, recall that the identity operator $\mathbb{I}$ always satisfies $\mathcal{L}^{\dagger}(\mathbb{I})=0$, expressing trace preservation of the Lindblad dynamics.
-Therefore, this system has a nontrivial subalgebra $\operatorname{span}\{\sigma_{z},\mathbb{I}\}\subset\mathcal{B}(\mathcal{H})$ of conserved observables, despite the presence of irreversible dissipation.
+Therefore, this system has a non-trivial subalgebra $\operatorname{span}\{\sigma_{z},\mathbb{I}\}\subset\mathcal{B}(\mathcal{H})$ of conserved observables, despite the presence of irreversible dissipation.
 4.2 Invariant commutative subalgebras and constants of motion
@@
 In the following sections, we will require that invariant commutative subalgebras arise as quantizations of Jacobi-commutative algebras of dissipated quantities associated with contact integrable systems.
+Such invariant commutative subalgebras encode families of observables whose expectation values are preserved by the open-system dynamics and which therefore behave as robust, effectively classical quantities in the presence of dissipation.
+As we show below, the GKSL structure provides a simple and natural mechanism ensuring the existence of such subalgebras under elementary commutation assumptions.
 In the Heisenberg picture, the generator $\mathcal{L}^{\dagger}$ defines a generally non-Hamiltonian evolution on the algebra of observables $\mathcal{B}(\mathcal{H})$, in the sense that it is not a derivation generated by a commutator with a self-adjoint operator, but includes dissipative terms of Lindblad type.
 As a consequence, the resulting dynamics is not an automorphism of the observable algebra and may exhibit contraction and loss of information.
-Such invariant commutative subalgebras encode families of observables whose expectation values are preserved by the open-system dynamics and which therefore behave as robust, effectively classical quantities in the presence of dissipation.
-As we show below, the GKSL structure provides a simple and natural mechanism ensuring the existence of such subalgebras under elementary commutation assumptions.
 Theorem 4.10 (Invariant commutative subalgebras for GKSL dynamics).
@@
 This analogy will be made precise in the following sections by requiring that the invariant quantum observables coincide, in the semiclassical limit, with the dissipated quantities of a completely integrable contact system.
-$\diamond$
 For the purposes of this work, the Heisenberg picture is particularly natural, since conserved observables are precisely characterized as elements of $\ker\mathcal{L}^{\dagger}$, and invariant commutative subalgebras can be directly identified.
@@
 $\|e^{t\mathcal{L}^{\dagger}}|_{\mathcal{A}^{\perp}}\|\;\leq\;C\,e^{-2\gamma t}$
-for some $C>0$. This exponential bound is the precise sense in which the dissipative dynamics suppresses the noncommutative (coherence) sector while leaving $\mathcal{A}$ pointwise invariant.
+for some $C>0$. This exponential bound is the precise sense in which the dissipative dynamics suppresses the non-commutative (coherence) sector while leaving $\mathcal{A}$ pointwise invariant.
 Here, the statement that $\mathcal{A}$ is pointwise invariant” means that the Heisenberg generator vanishes identically on $\mathcal{A}$, namely
@@
 This means that the Lindblad dynamics dynamically selects $\sigma_{z}$ as the only nontrivial observable whose expectation values remain stable under the open-system evolution.
-More generally, this example illustrates a robust structural phenomenon: for suitable GKSL generators, there exist nontrivial commutative $C^{\ast}$-subalgebras $\mathcal{A}\subset\mathcal{B}(\mathcal{H})$, with $\mathcal{A}\neq\operatorname{span}(\mathbb{I})$, that are invariant under the Heisenberg evolution and lie entirely in $\ker\mathcal{L}^{\dagger}$.
+More generally, this example illustrates a robust structural phenomenon: for suitable GKSL generators, there exist non-trivial commutative $C^{\ast}$-subalgebras $\mathcal{A}\subset\mathcal{B}(\mathcal{H})$, with $\mathcal{A}\neq\operatorname{span}(\mathbb{I})$, that are invariant under the Heisenberg evolution and lie entirely in $\ker\mathcal{L}^{\dagger}$.
 In the geometric framework developed in this paper, such invariant commutative subalgebras provide the natural quantum counterparts of the Jacobi-commutative algebras generated by dissipated quantities in contact integrable systems.
@@
 Throughout this subsection we assume that $(C,\alpha,h;I_{0},\dots,I_{n})$ arises from a bi-Hamiltonian Poisson–Lie system as described in Section 3.
-Consider the algebra $(\mathscr{C}^{\infty}(C),\{\cdot,\cdot\}_{\alpha})$, where $\{\cdot,\cdot\}_{\alpha}$ denotes de Jacobi bracket defined by $\alpha$.
+Consider the algebra $(\mathscr{C}^{\infty}(C),\{\cdot,\cdot\}_{\alpha})$, where $\{\cdot,\cdot\}_{\alpha}$ denotes the Jacobi bracket defined by $\alpha$.
 Let $\mathcal{A}_{\mathrm{cl}}\subset\mathscr{C}^{\infty}(C)$ denote the commutative subalgebra generated by $\{I_{0},\dots,I_{n}\}$.
@@
 Remark 4.15.
-A contact system is naturally encoded by its symplectization. Given a contact manifold $(C,\alpha)$, its symplectization is
-$(M,\omega)\mathrel{\vbox{\hbox{\scriptsize.}\hbox{\scriptsize.}}}=(C\times\mathbb{R}_{>0},\,\omega\mathrel{\vbox{\hbox{\scriptsize.}\hbox{\scriptsize.}}}=d(r\alpha)),$
-with Liouville (Euler) vector field $\Delta=r\partial_{r}$. Any function $f\in\mathscr{C}^{\infty}(C)$ admits a canonical 11–homogeneous lift on $M$, for instance
-$f^{\Sigma}(x,r)\mathrel{\vbox{\hbox{\scriptsize.}\hbox{\scriptsize.}}}=r\,f(x),$
-and the Jacobi bracket on $(C,\alpha)$ is recovered from the Poisson bracket on $(M,\omega)$ by restriction to the hypersurface $\{r=1\}$:
-$\{f,g\}_{\alpha}\;=\;\{f^{\Sigma},g^{\Sigma}\}_{\omega}\big|_{r=1}.$
+A contact system is naturally encoded by its symplectization (see Propoisition 2.8).
 In particular, the contact Hamiltonian dynamics on $C$ is the reduction (or, equivalently, the restriction to a transverse section $r=\mathrm{const}$) of the homogeneous Hamiltonian dynamics on $M$.
@@
 On the classical side, the Jacobi algebra $(\mathscr{C}^{\infty}(C),\{\cdot,\cdot\}_{\alpha})$ is not unital in the Poisson sense: although the constant function $1$ belongs to the algebra, it is not central in general, since
-$\{1,f\}_{\alpha}=-R(f),$
+$\{1,f\}_{\alpha}=R(f),$
 where $R$ is the Reeb vector field.
@@
 For this reason, the correct correspondence between classical and quantum structures should be understood as follows: the quantization map $Q_{\hbar}$ assigns operators to the nonconstant generators of the Jacobi-commutative algebra $\mathcal{A}_{\mathrm{cl}}$, while the unit is added automatically on the quantum side.
-The nontrivial content of contact compatibility and of quantum constants of motion therefore lies entirely in the quantization of the dissipated quantities, not in the unit element.
+The non-trivial content of contact compatibility and of quantum constants of motion therefore lies entirely in the quantization of the dissipated quantities, not in the unit element.
 This distinction explains why, in minimal examples, the classical condition $\{1,I\}_{\alpha}=0$ has no direct quantum analog, and clarifies the precise sense in which invariant commutative $C^{\ast}$-subalgebras quantize Jacobi-commutative algebras of dissipated quantities.
@@
 Let $\mathcal{A}_{\mathrm{cl}}=\mathrm{span}\{I\}$ be the Jacobi-commutative subalgebra generated by a single dissipated quantity $I$.
-Any semiclassical quantization $Q_{\hbar}$ maps the nontrivial generator $I$ to a self-adjoint operator $\widehat{I}_{\hbar}$.
+Any semiclassical quantization $Q_{\hbar}$ maps the non-trivial generator $I$ to a self-adjoint operator $\widehat{I}_{\hbar}$.
 The corresponding quantum algebra is then taken to be the unital $C^{\ast}$-algebra
@@
 A contact-compatible Lindblad generator is precisely one whose Heisenberg adjoint leaves $\mathcal{A}_{\hbar}$ pointwise invariant and satisfies $\mathcal{L}^{\dagger}(\widehat{I}_{\hbar})=0$.
-In this minimal setting, contact compatibility reduces to the existence of a nontrivial invariant commutative subalgebra generated by a single dissipated quantity beyond the scalars.
-This example highlights a general structural feature: while the constant function $1$ plays a nontrivial role in the Jacobi algebra at the classical level, its quantum counterpart is automatically invariant.
-The content of contact compatibility therefore lies entirely in the quantization of the nonconstant dissipated quantities.
+In this minimal setting, contact compatibility reduces to the existence of a non-trivial invariant commutative subalgebra generated by a single dissipated quantity beyond the scalars.
+This example highlights a general structural feature: while the constant function $1$ plays a non-trivial role in the Jacobi algebra at the classical level, its quantum counterpart is automatically invariant.
+The content of contact compatibility therefore lies entirely in the quantization of the non-constant dissipated quantities.
 For a set of functions $S\subseteq\mathscr{C}^{\infty}(C)$, we write $\mathrm{Jacobi\text{-}alg}\langle S\rangle$ for the smallest Jacobi subalgebra of $(\mathscr{C}^{\infty}(C),\{\cdot,\cdot\}_{\alpha})$ containing $S$, i.e.
@@
 be a Jacobi-commutative algebra of dissipated quantities, i.e.
-$\{I_{i},I_{j}\}_{\alpha}=0\quad\text{for all }i,j,\qquad\{h,I_{i}\}_{\alpha}=-\lambda_{i}I_{i}$
-for some contact Hamiltonian $h$ and real constants $\lambda_{i}$.
+$\{I_{i},I_{j}\}_{\alpha}=0\quad\text{for all }i,j,\qquad\{h,I_{i}\}_{\alpha}=0$
+for some contact Hamiltonian $h$.
 Assume that $Q_{\hbar}$ is a semiclassical quantization of $\mathcal{A}_{\mathrm{cl}}$ in the sense of Definition 4.14, and that there exists a family of Lindblad generators $\{\mathcal{L}_{\hbar}\}$ such that
@@
 $\widehat{I}_{i,\hbar},\widehat{I}_{j,\hbar}]=\mathcal{O}(\hbar).$
-Assumption (2) ensures that the Hamiltonian and the Lindblad operators are chosen inside the commutant of the $C^{\ast}$-algebra generated by the $\widehat{I}_{i,\hbar}$.
-Therefore, the operators $\widehat{I}_{i,\hbar}$ generate a commutative $C^{\ast}$-subalgebra $\mathcal{A}_{\hbar}\subset\mathcal{B}(\mathcal{H}_{\hbar})$, which is unital by construction, since any $C^{\ast}$-subalgebra of $\mathcal{B}(\mathcal{H}_{\hbar})$ contains the identity operator.
 Assumption (2) implies that the Hamiltonian and all Lindblad operators (and their adjoints) belong to the commutant $\mathcal{A}_{\hbar}^{\prime}$. By Lemma 4.7, it follows that
@@
 $\mathcal{L}_{\hbar}^{\dagger}(A)=0\quad\text{for all }A\in\mathcal{A}_{\hbar}.$
+Therefore, the operators $\widehat{I}_{i,\hbar}$ generate a commutative $C^{\ast}$-subalgebra $\mathcal{A}_{\hbar}\subset\mathcal{B}(\mathcal{H}_{\hbar})$, which is unital by construction, since any $C^{\ast}$-subalgebra of $\mathcal{B}(\mathcal{H}_{\hbar})$ contains the identity operator.
 Finally, the Egorov-type condition (1) ensures that, on the level of principal symbols, the Heisenberg generator reduces to the contact Hamiltonian derivation associated with $h$ when restricted to $\mathcal{A}_{\mathrm{cl}}$.
@@
 Let $(C,\alpha,h;I_{0},\dots,I_{n})$ and $(\mathcal{H}_{\hbar},Q_{\hbar})$ be as in Definition 4.14.
-Assume in addition that the contact system arises from a bi-Hamiltonian Poisson–Lie system $(G,\pi_{0},\pi_{1};H_{0},H_{1})$ as in Section 3, so that the classical vector field $X$ is Hamiltonian both for $\pi_{0}$ and $\pi_{1}$, and the hierarchy of integrals $\{I_{j}\}$ ultimately comes from the common bi-Hamiltonian invariants on $G$.
+Assume in addition that the contact system arises from a bi-Hamiltonian Poisson–Lie system $(G,\pi_{0},\pi_{1};H_{0},H_{1})$ as in Section 3, so that the classical vector field $X$ is Hamiltonian with respect to both $\pi_{0}$ and $\pi_{1}$, and the hierarchy of integrals $\{I_{j}\}$ ultimately comes from the common bi-Hamiltonian invariants on $G$.
 At the quantum level, the two compatible Poisson–Lie Hamiltonians $H_{0},H_{1}$ do not give rise directly to Lindblad generators, but rather induce, via the homogeneous lift and restriction procedure of Section 3, a pair of contact Hamiltonians on $C$.
@@
 Condition (BL2) captures the fact that the same family of classical integrals $I_{j}$ arises from the bi-Hamiltonian hierarchy and should therefore be conserved by both quantum evolutions.
-Finally, (BL3) encodes the requirement that, in the semiclassical limit, each branch $\mathcal{L}^{(a)}_{\hbar}$ recovers the contact dynamics associated with a different classical Hamiltonian $H_{a}$, in analogy with the way a bi-Hamiltonian vector field is Hamiltonian for both $\pi_{0}$ and $\pi_{1}$.
+Finally, (BL3) encodes the requirement that, in the semiclassical limit, each branch $\mathcal{L}^{(a)}_{\hbar}$ recovers the contact dynamics associated with a different classical Hamiltonian $H_{a}$, in analogy with the way a bi-Hamiltonian vector field is Hamiltonian with respect to both $\pi_{0}$ and $\pi_{1}$.
 Remark 4.22.
@@
 By (BL1), $\mathcal{L}^{(\lambda)}_{\hbar}$ is a GKSL generator for each $\lambda\in[0,1]$ and $\hbar\in(0,\hbar_{0}]$.
-Verification of (CC1). Since the Heisenberg adjoint is affine,
+Verification of (CC1). Since the Heisenberg adjoint is linear,
 $\mathcal{L}^{(\lambda)\,\dagger}_{\hbar}=(1-\lambda)\,\mathcal{L}^{(0)\,\dagger}_{\hbar}+\lambda\,\mathcal{L}^{(1)\,\dagger}_{\hbar}$,
@@
 Example: a bi-Lindblad pencil interpolating between unitary and dephasing dynamics.
-We illustrate Definition 4.21 on the two-level system $\mathcal{H}=\mathbb{C}^{2}$, where the invariant commutative sector is generated by $\sigma_{z}$.
+We illustrate Definition 4.21 on a two-level system $\mathcal{H}=\mathbb{C}^{2}$, where the invariant commutative sector is generated by $\sigma_{z}$.
 Fix parameters $\omega_{0},\omega_{1}\in\mathbb{R}$ and $\gamma>0$, and define two GKSL generators on $\mathcal{B}(\mathbb{C}^{2})$ (not forming a semiclassical family in $\hbar$, but sufficient to illustrate the algebraic content of Definition 4.21):
@@
 Note that $\mathcal{L}^{(0)}$ describes a purely unitary evolution, while $\mathcal{L}^{(1)}$ corresponds to a unitary evolution plus pure dephasing in the $\sigma_{z}$ basis.
-For every $\lambda\in[0,1]$, let
+For every $\lambda\in[0,1]$, the operator
 $\mathcal{L}^{(\lambda)}\mathrel{\vbox{\hbox{\scriptsize.}\hbox{\scriptsize.}}}=(1-\lambda)\mathcal{L}^{(0)}+\lambda\mathcal{L}^{(1)}$
@@
 Thus a bi-Lindblad structure provides, for each $\lambda\in[0,1]$, a specific choice of contact-compatible Lindblad generator whose semiclassical limit reflects a different point on the classical bi-Hamiltonian pencil.
-In the deformed Euler top example of Section 5, we shall see explicitly how the quantum generators $\mathcal{L}^{(0)}_{\hbar}$ and $\mathcal{L}^{(1)}_{\hbar}$ can be chosen so that their common quantum integrals $\widehat{I}_{j,\hbar}$ quantize the contact dissipated quantities $I_{j}$ obtained in Section 3, thereby realizing a concrete bi-Lindblad structure in the above sense.
+In the Euler-top-type example of Section 5, we shall see explicitly how the quantum generators $\mathcal{L}^{(0)}_{\hbar}$ and $\mathcal{L}^{(1)}_{\hbar}$ can be chosen so that their common quantum integrals $\widehat{I}_{j,\hbar}$ quantize the contact dissipated quantities $I_{j}$ obtained in Section 3, thereby realizing a concrete bi-Lindblad structure in the above sense.
 Weyl quantization.
 Before stating the Egorov-type criterion, we briefly recall what we mean by semiclassical Weyl quantization in the standard pseudodifferential setting.
-Assume that the exact symplectic manifold $(M,\omega)$ is globally symplectomorphic to $\mathbb{R}^{2d}$ (with $d=n+1$) with canonical coordinates $(x,\xi)\in\mathbb{R}^{\mathrm{d}}\times\mathbb{R}^{d}$ and symplectic form ${\omega=\sum_{j=1}^{d}\mathrm{d}\xi_{j}\wedge dx_{j}}$.
+Assume that the exact symplectic manifold $(M,\omega)$ is globally symplectomorphic to $\mathbb{R}^{2d}$ (with $d=n+1$) with canonical coordinates $(x,\xi)\in\mathbb{R}^{\mathrm{d}}\times\mathbb{R}^{d}$ and symplectic form ${\omega=\sum_{j=1}^{d}\mathrm{d}\xi_{j}\wedge\mathrm{d}x_{j}}$.
 In the following, $S^{0}(M)$ will denote the standard class of semiclassical symbols of order zero on $M\simeq\mathbb{R}^{2d}$, that is, the space of smooth functions $a(x,\xi;\hbar)$ satisfying
@@
 This is the pseudodifferential backbone behind the Dirac condition and the Egorov-type limit used below.
-We refer to [Zworski2012, DimassiSjoestrand1999] for comprehensive accounts of semiclassical Weyl quantization, symbol classes $S^{m}$, the Moyal product, and the Egorov theorem.
-See also [Folland1989, Martinez2002] for the Weyl calculus and its microlocal properties.
-From the perspective of deformation quantization, the Moyal product and the Dirac condition are discussed for instance in [Gutt2000].
+We refer to [58, 18] for comprehensive accounts of semiclassical Weyl quantization, symbol classes $S^{m}$, the Moyal product, and the Egorov theorem.
+See also [23, 48] for the Weyl calculus and its microlocal properties.
+From the perspective of deformation quantization, the Moyal product and the Dirac condition are discussed for instance in [34].
 In our contact setting, the quantization map $Q_{\hbar}$ on the Jacobi subalgebra $\mathcal{A}_{\mathrm{cl}}\subset\mathscr{C}^{\infty}(C)$ is obtained by first taking the homogeneous lift $f\mapsto f^{\Sigma}$ (so $f^{\Sigma}|_{C}=f$) and then applying Weyl quantization:
@@
 We illustrate Theorem 4.24 on the linear contact model of Example 2.10, with
-$(C,\alpha,h)=(\mathbb{R}^{3}(q,p,z),\,\mathrm{d}z-p\,\mathrm{d}q,\,h=z-p),\qquad I_{1}\mathrel{\vbox{\hbox{\scriptsize.}\hbox{\scriptsize.}}}=e^{-z}.$
+$(C,\alpha,h)=(\mathbb{R}^{3}(q,p,z),\,\mathrm{d}z-p\,\mathrm{d}q,\,h=z-p),\qquad I_{1}\mathrel{\vbox{\hbox{\scriptsize.}\hbox{\scriptsize.}}}=z.$
 This system arises by restriction from the exact symplectic homogeneous system
@@
 with Liouville-transverse hypersurface $C=\{r=1\}$. The Jacobi subalgebra is
-$\mathcal{A}_{\mathrm{cl}}=\mathrm{Jacobi\text{-}alg}\langle I_{0}=h,\,I_{1}=e^{-z}\rangle\subset\mathscr{C}^{\infty}(C).$
+$\mathcal{A}_{\mathrm{cl}}=\mathrm{Jacobi\text{-}alg}\langle I_{0}=h,\,I_{1}=z\rangle\subset\mathscr{C}^{\infty}(C).$
 Assume that $M$ is quantized by semiclassical Weyl quantization on $\mathbb{R}^{4}$, so that
@@
 for all $f\in\mathcal{A}_{\mathrm{cl}}$.
-As a consequence, the Lindblad operators belong to $\mathcal{A}_{\hbar}$ and hence
+As a consequence, the Lindblad operators belong to $\mathcal{A}_{\hbar}$, and hence
 $L_{k,\hbar},A]=[L_{k,\hbar}^{\dagger},A]=0\quad\forall\,A\in\mathcal{A}_{\hbar}.$
@@
 Fix $\hbar>0$ and consider the commuting family of quantized observables
-$\widehat{I}_{0,\hbar}=\widehat{h}_{\hbar},\qquad\widehat{I}_{1,\hbar}=\operatorname{Op}_{\hbar}^{\mathrm{w}}(I_{1}^{\Sigma}),\qquad I_{1}=e^{-z}.$
+$\widehat{I}_{0,\hbar}=\widehat{h}_{\hbar},\qquad\widehat{I}_{1,\hbar}=\operatorname{Op}_{\hbar}^{\mathrm{w}}(I_{1}^{\Sigma}),\qquad I_{1}=z.$
 Let
@@
 5 An Euler-top-type Poisson–Lie pencil as a bi-Lindblad system
-In this section we illustrate how the abstract framework developed above can be implemented in a concrete low-dimensional model of Euler-top type, inspired by the deformed Euler top of Marrero and collaborators [ballesteros2017poisson, G.I.M+2023].
+In this section we illustrate how the abstract framework developed above can be implemented in a concrete low-dimensional model of Euler-top type, inspired by the deformed Euler top of Marrero and collaborators [3, 33].
 Rather than reproducing every physical feature of the rigid-body Euler top, we focus on the structural ingredients that are essential for our construction: (i) a linear Poisson pencil of Poisson–Lie origin, (ii) a bi-Hamiltonian vector field admitting a Lenard–Magri hierarchy of commuting invariants, (iii) a homogeneous lift yielding a completely integrable contact system with nontrivial dissipated quantities, and (iv) an algebraic setting in which one can build Lindblad generators implementing pure dephasing relative to the quantum integrals.
@@
 $\Phi:(S,\omega)\longrightarrow(\mathfrak{so}(3)^{*},\pi_{0}).$
-Such exact symplectic realizations of linear Poisson manifolds are classical and can be constructed explicitly; see, for instance, [C.F.M2021, Section 6.3].
+Such exact symplectic realizations of linear Poisson manifolds are classical and can be constructed explicitly; see, for instance, [12, Section 6.3].
 In particular, for $\mathfrak{so}(3)^{*}\simeq\mathfrak{su}(2)^{*}$ one may take $S=N\times\mathbb{R}^{3}_{m}$, where $N$ is a $3$-manifold equipped with a coframe $\{\vartheta_{1},\vartheta_{2},\vartheta_{3}\}$ satisfying the Maurer–Cartan relations, and define
@@
 $\Delta=\sum_{i=1}^{3}m_{i}\frac{\partial}{\partial m_{i}}\,$,
-and hence $F_{1}$ and $F_{2}$ are homogeneous of degree $2$. Let us define the functions $I_{a}^{\Sigma}\colon U\to\mathbb{R}$, defined on the open subset $U=S\setminus(N\times\{0\})$, given by
+and hence $F_{1}$ and $F_{2}$ are homogeneous of degree $2$, i.e., $\Delta F_{a}=2F_{a},\,a=0,1$.
+Let us define the functions $I_{a}^{\Sigma}\colon U\to\mathbb{R}$, defined on the open subset $U=S\setminus(N\times\{0\})$, given by
 $I_{0}^{\Sigma}=\sqrt{-2F_{0}}=\sqrt{m_{1}^{2}+m_{2}^{2}+m_{3}^{2}}\,,\quad I_{1}^{\Sigma}=\sqrt{F_{1}}=\sqrt{m_{1}^{2}+m_{2}m_{3}}\,.$
@@
 Semiclassically, the joint spectral parameter $\nu=(\nu_{0},\nu_{1})$ labels, up to the usual Bohr–Sommerfeld quantization conditions (see e.g.
-[GuilleminSternberg1990]), the common level sets $\{I_{0}=\nu_{0},\ I_{1}=\nu_{1}\}$ of the underlying contact integrable system.
+[32]), the common level sets $\{I_{0}=\nu_{0},\ I_{1}=\nu_{1}\}$ of the underlying contact integrable system.
 The direct-integral fibres $\mathcal{H}_{\hbar,\nu}$ may therefore be interpreted as quantum counterparts of the classical invariant leaves (Liouville tori/cylinders after reduction), providing the quantum “fibres” associated with the classical invariant foliation.
@@
 $\frac{\mathrm{d}}{\mathrm{d}t}\rho_{\nu\mu}(t)\Big|_{\rm diss}=-\frac{1}{2}\sum_{k=1}^{N}\big(\ell_{k,\nu}-\ell_{k,\mu}\big)^{2}\,\rho_{\nu\mu}(t),$ (5.12)
-while the diagonal blocks $\rho_{\nu\nu}$ are unaffected by the dissipator. Hence the off-diagonal coherences decay exponentially,
-$\|\rho_{\nu\mu}(t)\|\;\lesssim\;\exp\!\Big(-\tfrac{1}{2}\,t\sum_{k=1}^{N}(\ell_{k,\nu}-\ell_{k,\mu})^{2}\Big)\,\|\rho_{\nu\mu}(0)\|,\qquad\nu\neq\mu,$
+while the diagonal blocks $\rho_{\nu\nu}$ are unaffected by the dissipator.
+Hence the off-diagonal coherences decay exponentially, that is, there exists a constant $C>0$ such that, for all $t\geq 0$ and $\nu\neq\mu$,
+$\|\rho_{\nu\mu}(t)\|\;\leq\;C\,\exp\!\Big(-\tfrac{1}{2}\,t\sum_{k=1}^{N}(\ell_{k,\nu}-\ell_{k,\mu})^{2}\Big)\,\|\rho_{\nu\mu}(0)\|\,$,
 whereas the populations $\operatorname{Tr}(\rho P_{\nu})$ are preserved.
@@
 In the present minimal Euler-top pencil, one may simply take $H_{0}^{\mathrm{eff}}=I_{0}$ and $H_{1}^{\mathrm{eff}}=I_{1}$, up to irrelevant affine reparametrizations.
-Let $\widehat{H}_{0,\hbar}$ and $\widehat{H}_{1,\hbar}$ be corresponding semiclassical quantizations.
+Let $\widehat{H}_{0,\hbar}$ and $\widehat{H}_{1,\hbar}$ be the corresponding semiclassical quantizations.
 We define two Lindblad generators on $\mathcal{B}(\mathcal{H}_{\hbar})$ by
@@
 The associated pointer sectors are therefore in canonical correspondence with the leaves of the classical contact foliation singled out by the integrable structure.
-This provides a natural setting in which decoherence, integrability, and the semiclassical limit are aligned by construction rather than enforced by model-dependent assumptions: decoherence dynamically selects the very algebra that encodes the classical integrable geometry.
+This provides a natural setting in which decoherence, integrability, and the semiclassical limit are aligned by construction rather than enforced by model-dependent assumptions: decoherence dynamically selects the algebra that encodes the classical integrable geometry.
 $\diamond$
@@
 This perspective opens the door to systematic constructions of open quantum models in which dissipation, integrability, and semiclassical consistency are aligned by design, rather than imposed by ad hoc assumptions.
-Conflict of Interest Statement.
+Declarations
+Funding
+L. Colombo acknowledges financial support from Grant PID2022-137909-NB-C22 funded by the Spanish Ministry of Science and Innovation.
+Competing Interests
 The authors declare no known competing financial interests or personal relationships that could have appeared to influence the work reported in this manuscript.
-Data Availability Statement.
+Data Availability Statement
 No datasets were generated or analyzed during the current study. All mathematical derivations are contained within the article.